### v3

# Quantum cohomology of projective bundles

Abstract.
We construct an $I$-function of the projective bundle $\mathbb{P}(V)$ associated with a not necessarily split vector bundle $V\to B$ as a Fourier transform of the $S^{1}$-equivariant $J$-function of the total space of $V$ and show that it lies on the Givental Lagrangian cone of $\mathbb{P}(V)$. Using this result, we show that the quantum cohomology $D$-module of $\mathbb{P}(V)$ splits into a direct sum of the quantum cohomology $D$-modules of the base space $B$. This has applications to the semisimplicity of big quantum cohomology.
2020 Mathematics Subject Classification:
Primary 14N35, Secondary 53D45
1. Introduction
For a rank-$r$ vector bundle $V\to B$ over a smooth projective variety $B$, we can construct the projective bundle $\mathbb{P}(V)\to B$ consisting of lines in the fibres of $V\to B$. The ordinary cohomology group $H^{*}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is generated by the relative hyperplane class $p:=c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))$ as an $H^{*}(B)$-algebra as follows:
$H^{*}(\mathbb{P}(V))\cong H^{*}(B)[p]/(p^{r}+c_{1}(V)p^{r-1}+\cdots+c_{r}(V)).$
This isomorphism is a consequence of the Leray-Hirsch theorem that says that $H^{*}(\mathbb{P}(V))$ is a free $H^{*}(B)$-module generated by $1,p,\dots,p^{r-1}$.
Many researchers have pursued a quantum cohomology counterpart to the Leray-Hirsch theorem for projective bundles or even broader fibre bundles, see e.g. [47, 10, 1, 13, 38, 3, 49, 37, 36, 20, 14, 16, 15]. If we count only ‘vertical’ curves contracted by the projection $\mathbb{P}(V)\to B$, we can easily observe that the corresponding ‘small vertical’ quantum cohomology of $\mathbb{P}(V)$ is given by
$H^{*}(B)[p,q]/(p^{r}+c_{1}(V)p^{r-1}+\cdots+c_{r}(V)-q)$
where $q$ is the Novikov variable associated with vertical curves. It decomposes into a direct sum of $H^{*}(B)$ as a ring for $q\neq 0$. This decomposition must persist under full quantum deformation, parametrized by $q$, the Novikov variable $Q$ of $B$ and the bulk parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$. The question then arises as to whether each summand is isomorphic to the quantum cohomology of $B$. This paper aims to affirm this correspondence, i.e. $QH^{*}(\mathbb{P}(V))$ decomposes into a direct sum of $QH^{*}(B)$. Furthermore, we show that the decomposition can be lifted to quantum cohomology $D$-modules.
1.1. Mirror theorem for non-split projective bundles
When the vector bundle $V\to B$ is split, i.e. a direct sum of line bundles, the Elezi-Brown mirror theorem [13, 3] enables us to determine the quantum cohomology of $\mathbb{P}(V)$. This theorem provides an explicit hypergeometric series, called the $I$-function, that lies in the Givental cone of $\mathbb{P}(V)$. Recall that the Givental cone of a smooth projective variety $X$ is a Lagrangian submanifold in the infinite dimensional symplectic vector space $\mathcal{H}_{X}=H^{*}(X)(\!(z^{-1})\!)[\![Q]\!]$. It encodes all genus-zero Gromov-Witten invariants of $X$. The Elezi-Brown $I$-function for $\mathbb{P}(V)$ reconstructs the Givenal cone and hence determines all the genus-zero Gromov-Witten invariants of $\mathbb{P}(V)$.
Our first main result is a generalization of the Elezi-Brown mirror theorem to the non-split case. It is stated in terms of the $S^{1}$-equivariant Gromov-Witten invariants of the vector bundle $V\to B$.
Theorem 1.1 (Theorem 3.3).
Let $B$ be a smooth projective variety and let $V\to B$ be a vector bundle of rank $r\geq 2$. We assume that the dual bundle $V^{\vee}$ is generated by global sections. Consider the $S^{1}$-action on $V$ scaling fibres and let $J_{V}^{\lambda}(\tau)$ denote the $S^{1}$-equivariant $J$-function of the total space of $V$, where $\lambda$ is the $S^{1}$-equivariant parameter. Define the $H^{*}(\mathbb{P}(V))$-valued function $I_{\mathbb{P}(V)}(\tau,t)$ by
$I_{\mathbb{P}(V)}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}\prod_{\begin{subarray}{c}\text{\rm$\delta$: Chern roots}\\ \text{\rm of $V$}\end{subarray}}(p+\delta+cz)}J_{V}^{p+kz}(\tau).$
Then $zI_{\mathbb{P}(V)}(\tau,t)|_{z\to-z}$ lies in the Givental cone of $\mathbb{P}(V)$. Here $q$ denotes the Novikov variable corresponding to the class of a line in a fibre and $p$ is the first Chern class of the relative $\mathcal{O}(1)$ over $\mathbb{P}(V)$.
The function $I_{\mathbb{P}(V)}$ in this theorem coincides with the Elezi-Brown $I$-function when $V$ is a direct sum of line bundles (see Example 3.4). It arises as the Fourier transform of the equivariant $J$-function $J_{V}^{\lambda}$ with respect to the equivariant parameter $\lambda$. We shall elaborate on this point in §1.3 below.
Remark 1.2.
By tensoring $V$ with a sufficiently negative line bundle, we can always assume that $V^{\vee}$ is generated by global sections, without changing $\mathbb{P}(V)$. This assumption ensures that $J_{V}^{\lambda}$ does not contain negative powers of $\lambda$ and that the substitution $\lambda=p+kz$ in the theorem is well-defined (see §2.3). The conclusion of Theorem 1.1 holds true even if we replace the $J$-function $J_{V}^{\lambda}$ with any slices of the Givental cone of $V$ depending polynomially on $\lambda$; see Theorem 3.3 for the details.
Remark 1.3.
The equivariant $J$-function $J_{V}^{\lambda}$ is defined over the Novikov ring of $B$. We use the splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})$ given by the class $p=c_{1}(\mathcal{O}(1))$ to identify Novikov variables for $B$ with Novikov variables for $\mathbb{P}(V)$, see Notation 3.1.
Remark 1.4.
The quantum Riemann-Roch theorem of Coates-Givental [7] determines the $S^{1}$-equivariant $J$-function $J_{V}^{\lambda}$ from the genus-zero Gromov-Witten invariants of the base $B$ and the Chern classes $c_{i}(V)$. Therefore Theorem 1.1 determines all genus-zero Gromov-Witten invariants of $\mathbb{P}(V)$ from those of the base $B$. This recovers the genus-zero part of the results of Fan-Lee [16] and Fan [15], who showed that all genus Gromov-Witten invariants of $\mathbb{P}(V)$ can be reconstructed from those of the base $B$ and the Chern classes $c_{i}(V)$. On the other hand, in contrast to the split case, it is not easy to find a closed formula for $J_{V}^{\lambda}$ in general. We give a non-split example, the cotangent bundle of $\mathbb{P}^{n}$, in Example 3.5.
Remark 1.5.
For other techniques to compute Gromov-Witten invariants of non-split projective bundles, we refer the reader to the reconstruction algorithm [49] and the quantum splitting principle [36, 14].
Remark 1.6.
In the subsequent paper [35], the second-named author generalized Theorem 1.1 to toric bundles associated with non-split vector bundles, combining the techniques in this paper and virtual localization for an additional torus action.
1.2. Decomposition of quantum $D$-modules
An important property of quantum cohomology is that it forms a $D$-module over the parameter space. Roughly speaking, the quantum $D$-module $\operatorname{QDM}(X)$ of a smooth projective variety $X$ is the module $H^{*}(X)[z][\![Q,\tau]\!]$ equipped with the flat covariant derivatives $\nabla_{\tau^{i}}$, $\nabla_{z\partial_{z}}$, $\nabla_{\xi Q\partial_{Q}}$ with respect to the parameters $\tau$, $z$, $Q$ (see §2.2). Here $\tau\in H^{*}(X)$ represents the bulk parameter and $Q$ denotes the Novikov variable. These covariant derivatives, called the quantum connection, are defined in terms of the quantum multiplication. The quantum $D$-module is also equipped with a $z$-sesquilinear pairing $P$, induced from the Poincaré pairing and compatible with the quantum connection. The $I$- or $J$-functions appearing in the previous section are cohomology-valued solutions of the quantum $D$-module.
The Mirror Theorem 1.1 establishes that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ is a Fourier transform of the $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ (see Theorem 1.11 below). Using this fact together with the quantum Riemann-Roch (QRR) theorem of Coates-Givental [7], which connects the quantum $D$-modules of $V$ and $B$, we derive the following decomposition theorem for $\operatorname{QDM}(\mathbb{P}(V))$.
$Q$$q^{-c_{1}(V)/r}Q$$q$$q^{-1/r^{\prime}}$(($\operatorname{QDM}(\mathbb{P}(V))$$\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)$
Figure 1. Decomposition pictured over the Kähler moduli space: $\operatorname{QDM}(\mathbb{P}(V))$ splits into a direct sum of $\operatorname{QDM}(B)$ near the infinity divisor $(q^{-1/r^{\prime}}=0)$, where $Q$ is the Novikov variable of the base $B$ and $q$ is the Novikov variable of the fibre $\mathbb{P}^{r-1}$.
Theorem 1.7 (Theorem 5.1; see also Figure 1).
Let $V\to B$ be as in Theorem 1.1. We have a map $\varsigma=\bigoplus_{j=0}^{r-1}\varsigma_{j}\colon H^{*}(\mathbb{P}(V))\to\bigoplus_{j=0}^{r-1}H^{*}(B)$ and a decomposition of the quantum $D$-modules
$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{i=0}^{r-1}\varsigma_{i}^{*}\operatorname{QDM}(B)_{\rm loc}$
where $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ and $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ denote the localized quantum $D$-modules of $\mathbb{P}(V)$ and $B$ defined over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$, where $Q$ is the Novikov variable of the base $B$ and $r^{\prime}$ is in (5.1).
Corollary 1.8.
The derivative of the map $\varsigma$ induces an isomorphism of the quantum cohomology rings $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{i=0}^{r-1}(H^{*}(B),\star_{\varsigma_{i}(\hat{\tau})})$. In particular, the quantum cohomology of $\mathbb{P}(V)$ is generically semisimple if and only if the same holds for the base $B$.
We can use Corollary 1.8 to deduce the generic semisimplicity of quantum cohomology for certain Grassmannians $G(k,n)$, $IG(k,2n)$ (see Proposition 5.11). The semisimplicity result seems new for $IG(k,2n)$ with $k\geq 3$ (see [45, 17, 11] for $k=2$).
Remark 1.9.
We do not discuss the convergence of quantum cohomology. In this paper, “generic semisimplicity” means formal semisimplicity in the sense of [9].
The proofs of Theorems 1.1 and 1.7 both rely on the QRR theorem [7]. We embed $V$ to the trivial bundle $B\times\mathbb{C}^{N}$. This induces an embedding $\mathbb{P}(V)\subset B\times\mathbb{P}^{N-1}$, which exhibits $\mathbb{P}(V)$ as the zero-locus of a section of a convex vector bundle $\mathcal{Q}(1)$ over $B\times\mathbb{P}^{N-1}$. Theorem 1.1 follows by applying the QRR theorem to the bundle $\mathcal{Q}(1)$. The Decomposition Theorem 1.7 is established by studying a (formal) stationary phase approximation of a certain Fourier integral (5.4) of $J_{V}^{\lambda}$. The QRR theorem for the bundle $V\to B$ implies that this integral takes values in the Givental cone of $B$, thereby defining a morphism $\operatorname{QDM}(\mathbb{P}(V))\to\operatorname{QDM}(B)$ of $D$-modules.
Remark 1.10.
In the split case, a similar decomposition of the quantum $D$-module (more generally, for toric bundles) follows from Brown’s theorem [3]: this has been discussed by Coates-Givental-Tseng [8] and Koto [34]. These previous works rely on the equivariant localization with respect to the (×$(\mathbb{C}^{\times})^{r-1}$-action (which exists only in the split case) and our argument is substantively different from them.
1.3. Equivariant quantum cohomology and Fourier transformation
We explain the idea behind our construction of the $I$-function. Following the conjecture of Teleman [50] on quantum cohomology of symplectic reductions, we view the quantum $D$-module of $\mathbb{P}(V)$ as a Fourier transform of the equivariant quantum $D$-module of $V$. We refer to [29, 31] for a more general conjecture along this line of thought.
Let $(M,\omega)$ be a sympletic manifold equipped with a Hamiltonian $S^{1}$-action, and let $\mu\colon M\to\mathbb{R}$ be a moment map. We denote by $M/\!\!/_{t}S^{1}$ the symplectic reduction $\mu^{-1}(t)/S^{1}$ at the level $t\in\mathbb{R}$. The equivariant volume of $M$ can be written as a Fourier transform of the volume of $M/\!\!/_{t}S^{1}$:
[…]
where $\lambda$ is the $S^{1}$-equivariant parameter and $\omega_{\rm red}$ is the reduced symplectic form on $M/\!\!/_{t}S^{1}$. This formula has played a fundamental role in the works of Duistermaat-Heckman [12] and Jeffrey-Kirwan [32]. We want to ask if a similar relation continues to hold in equivariant quantum cohomology. Since the $J$-function is a quantum analogue of $e^{\omega}$, we could perhaps hope that the equivariant $J$-function of $M$ is a Fourier transform of the $J$-function of the reduction $M/\!\!/_{t}S^{1}$.
We can make this hope more precise by using shift (or Seidel) operators in equivariant quantum cohomology. The $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(M)$ of $M$ admits a shift operator $\mathbb{S}\colon\operatorname{QDM}_{S^{1}}(M)\to\operatorname{QDM}_{S^{1}}(M)$ of the equivariant parameter $\lambda$ [42]. This lifts the Seidel operator [48] on quantum cohomology and satisfies $\mathbb{S}(f(\lambda)\alpha)=f(\lambda-z)\mathbb{S}(\alpha)$ for any $f(\lambda)\in\mathbb{C}[\lambda]$ and $\alpha\in\operatorname{QDM}_{S^{1}}(M)$. We have the following commutation relation:
$[\lambda,\mathbb{S}]=z\mathbb{S}.$
Now we change the point of view and regard $\operatorname{QDM}_{S^{1}}(M)$ as a $\mathbb{C}[\mathbb{S}]$-module rather than a $\mathbb{C}[\lambda]$-module; the equivariant parameter $\lambda$ then behaves like the differential operator $z\mathbb{S}\frac{\partial}{\partial\mathbb{S}}$. In other words, we consider the Fourier transform of the module $\operatorname{QDM}_{S^{1}}(M)$. Teleman’s conjecture suggests that the Fourier transform of $\operatorname{QDM}_{S^{1}}(M)$ corresponds to $\operatorname{QDM}(M/\!\!/_{t}S^{1})$ in such a way that the $\lambda$-action gives the quantum connection in the direction of $\kappa(\lambda)$ and that the $\mathbb{S}$-action gives the Novikov variable action dual to $\kappa(\lambda)$, where $\kappa\colon H^{*}_{S^{1}}(M)\to H^{*}(M/\!\!/_{t}S^{1})$ denotes the Kirwan map.
According to this picture, solutions of $\operatorname{QDM}_{S^{1}}(M)$ and $\operatorname{QDM}(M/\!\!/_{t}S^{1})$ should be related by a Fourier transformation. Let us restrict ourselves to the case where $M$ is the total space of $V\to B$. The equivariant $J$-function $J_{V}^{\lambda}$ gives a solution of $\operatorname{QDM}_{S^{1}}(V)$ which takes values in the (polynomial) Givental space $\mathcal{H}_{V,\rm pol}=H^{*}_{S^{1}}(V)[z,z^{-1}][\![Q]\!]$. The shift action on the Givental space $\mathcal{H}_{V,\rm pol}$ is given by the operator $\mathcal{S}\colon\mathcal{H}_{V,\rm pol}\to\mathcal{H}_{V,\rm pol}$
(1.1) $\mathcal{S}(f(\lambda))=e_{\lambda}(V)f(\lambda-z)$
where $e_{\lambda}(V)=\sum_{i=0}^{r}\lambda^{i}c_{r-i}(V)$ is the $S^{1}$-equivariant Euler class of $V$ (see [27, Definition 3.13]). We consider the following Fourier transformation from $\mathcal{H}_{V,\rm pol}$ to $\mathcal{H}_{\mathbb{P}(V),\rm pol}$:
(1.2) $J(\lambda)\longmapsto\widehat{J}(q)=\sum_{k\in\mathbb{Z}}\kappa(\mathcal{S}^{-k}J)q^{k}$
where $\kappa\colon H^{*}_{S^{1}}(V)\to H^{*}(\mathbb{P}(V))$ is the Kirwan map, mapping $\lambda$ to $p=c_{1}(\mathcal{O}(1))$, and $\mathcal{H}_{\mathbb{P}(V),\rm pol}=H^{*}(\mathbb{P}(V))[z,z^{-1}][\![q,Q]\!]$ is the polynomial Givental space for $\mathbb{P}(V)$. It is easy to see that this transformation intertwines $\lambda$ with $zq\partial_{q}+p$, and $\mathcal{S}$ with $q$.
(1.3) $(\lambda J)^{^}=\left(zq\partial_{q}+p\right)\widehat{J},\qquad(\mathcal{S}J)^{^}=q\widehat{J}.$
Hence the Fourier transform of $J_{V}^{\lambda}$ should give rise to a solution of $\operatorname{QDM}(\mathbb{P}(V))$. It is calculated as
(1.4) $(J_{V}^{\lambda})^{^}$ $=\sum_{k\in\mathbb{Z}}\kappa\left(\frac{\prod_{c=-\infty}^{0}e_{\lambda+cz}(V)}{\prod_{c=-\infty}^{k}e_{\lambda+cz}(V)}J_{V}^{\lambda+kz}\right)q^{k}=\sum_{k\geq 0}\frac{q^{k}}{\prod_{c=1}^{k}e_{p+cz}(V)}J_{V}^{p+kz}.$
The summand with $k<0$ vanishes because of the relation $\kappa(e_{\lambda}(V))=p^{r}+c_{1}(V)p^{r-1}+\cdots+c_{r}(V)=0$ in $H^{*}(\mathbb{P}(V))$. This gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1 restricted to $t=0$.
We can rephrase the mirror theorem (Theorems 1.1, 3.3) as the Fourier duality between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$, as follows:
Theorem 1.11 (Theorem 4.2, Proposition 4.4).
Let $V\to B$ be as in Theorem 1.1. There exist a mirror map $\hat{\tau}\colon H^{*}_{S^{1}}(V)\to H^{*}(\mathbb{P}(V))$, whose leading term is the Kirwan map, and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\cong\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ such that $\operatorname{FT}$ intertwines
  • the shift operator $\mathbb{S}$ with the Novikov variable $q$ and;
  • the equivariant parameter $\lambda$ with the quantum connection $z(\hat{\tau}^{*}\nabla)_{q\partial_{q}}$.
Moreover, the natural pairings on $\operatorname{QDM}_{S^{1}}(V)$ and $\operatorname{QDM}(\mathbb{P}(V))$ can be related by a Fourier transformation.
1.4. Plan of the paper
In §2, we review Gromov-Witten invariants, Givental formalism and shift operators. We prove a technical Lemma 2.8 on miniveral elements of the Givental cone, which plays a crucial role in the proof of the mirror theorem. In §3, we prove our mirror theorem and give examples. In §4, we reformulate the mirror theorem as a Fourier duality between quantum $D$-modules of $V$ and $\mathbb{P}(V)$. In §5, we prove a decomposition of the quantum $D$-module of $\mathbb{P}(V)$.
Acknowledgements
H.I. thanks Fumihiko Sanda, Yuuji Tanaka and Constantin Teleman for very useful discussions on shift operators and equivariant cohomology. H.I. is supported by JSPS grant 16H06335, 20K03582, 21H04994 and 23H01073. Y.K. is supported by JSPS grant 21J23023.
2. Preliminaries
We introduce notation on Gromov-Witten invariants and quantum cohomology. Then we discuss quantum connection, the Givental Lagrangian cone, quantum Riemann-Roch (QRR) theorem and equivariant shift operators.
2.1. Gromov-Witten invariants and quantum cohomology
A reference on Gromov-Witten invariants is made to [40]. Let $X$ be a smooth projective variety over $\mathbb{C}$. For $n\geq 0$ and $d\in H_{2}(X,\mathbb{Z})$, let $X_{0,n,d}$ denote the moduli stack of genus-zero, $n$-points, degree $d$ stable maps to $X$. This is a proper Deligne-Mumford stack carrying a virtual fundamental class $[X_{0,n,d}]_{\rm vir}\in H_{2D}(X_{0,n,d},\mathbb{Q})$ of degree $D=\dim_{\mathbb{C}}X+c_{1}(TX)\cdot d+n-3$. Let $\operatorname{ev}_{i}\colon X_{0,n,d}\to X$ be the evaluation map at the $i$th marked point. The (genus-zero, descendant) Gromov-Witten invariants are the integrals of the form:
$\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle_{0,n,d}^{X}=\int_{[X_{0,n,d}]_{\rm vir}}\prod_{i=1}^{n}\operatorname{ev}_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}$
where $\alpha_{1},\dots,\alpha_{n}\in H^{*}(X)$, $k_{1},\dots,k_{n}$ are nonnegative integers, and $\psi_{i}$ is the first Chern class of the universal cotangent line bundle at the $i$th marked point.
Let $\operatorname{Eff}(X)\subset H_{2}(X,\mathbb{Z})$ denote the semigroup of effective curve classes. For a module $K$, we write $K[\![Q]\!]$ for the space of formal power series $\sum_{d\in\operatorname{Eff}(X)}k_{d}Q^{d}$ with $k_{d}\in K$. The formal variable $Q$ is called the Novikov variable. The module $K[\![Q]\!]$ has a natural ring structure when $K$ is a ring. The quantum cohomology will be defined over the ring $\mathbb{C}[\![Q]\!]$, which is called the Novikov ring.
Let $(\alpha,\beta)_{X}=\int_{X}\alpha\cup\beta$ denote the Poincaré pairing on $H^{*}(X)$. Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(X)$. We denote by $\tau=\sum_{i=0}^{s}\tau^{i}\phi_{i}$ a general point in $H^{*}(X)$; then $\tau^{0},\dots,\tau^{s}$ form a $\mathbb{C}$-linear coordinate system on $H^{*}(X)$. The quantum product $\star_{\tau}$ at a parameter $\tau\in H^{*}(X)$ is defined by
(2.1) $(\alpha\star_{\tau}\beta,\gamma)_{X}=\sum_{n\geq 0}\sum_{d\in\operatorname{Eff}(X)}\left\langle\alpha,\beta,\gamma,\tau,\cdots,\tau\right\rangle_{0,n+3,d}^{X}\frac{Q^{d}}{n!}$
for $\alpha,\beta,\gamma\in H^{*}(X)$. The product $\star_{\tau}$ introduces an associative and supercommutative ring structure on $H^{*}(X)[\![Q,\tau]\!]=H^{*}(X)[\![Q]\!][\![\tau^{0},\dots,\tau^{s}]\!]$ such that $1\in H^{0}(X)$ is the unity. This is called quantum cohomology. The variables $\tau^{i}$ associated with odd classes $\phi_{i}$ have odd parity, and therefore anti-commute with each other by supercommutativity; more precisely we have $\tau^{i}\tau^{j}=(-1)^{|i||j|}\tau^{j}\tau^{i}$, $\tau^{i}\phi_{j}=(-1)^{|i||j|}\phi_{j}\tau^{i}$ with $|i|=\deg\phi_{i}$. The Novikov variable $Q$ is even and commutes with other elements. We also follow the Koszul sign convention when expanding the correlator $\left\langle\alpha,\beta,\gamma,\tau,\dots,\tau\right\rangle_{0,n+3,d}^{X}$ in power series of $\tau^{i}$. The parameter space $H^{*}(X)$ should be viewed as a formal supermanifold over $\mathbb{C}[\![Q]\!]$; we refer to [40] for the details.
2.2. Quantum connection and fundamental solution
The quantum connection is a meromorphic flat connection $\nabla$ on the vector bundle $H^{*}(X)\times(H^{*}(X)\times\mathbb{C})\to H^{*}(X)\times\mathbb{C}$ defined by the formulae
(2.2) $\begin{split}\nabla_{\tau^{i}}&=\partial_{\tau^{i}}+z^{-1}(\phi_{i}\star_{\tau})\\ \nabla_{z\partial_{z}}&=z\partial_{z}-z^{-1}(E_{X}\star_{\tau})+\mu_{X}\end{split}$
where $(\tau,z)$ denotes a point on the base $H^{*}(X)\times\mathbb{C}$, $E_{X}$ is the Euler vector field (a section of this vector bundle)
(2.3) $E_{X}=c_{1}(TX)+\sum_{i=0}^{s}\left(1-\frac{\deg\phi_{i}}{2}\right)\tau^{i}\phi_{i}$
and $\mu_{X}\in\operatorname{End}(H^{*}(X))$ is the grading operator
(2.4) $\mu_{X}(\phi_{i})=\left(\frac{\deg\phi_{i}}{2}-\frac{\dim_{\mathbb{C}}X}{2}\right)\phi_{i}.$
More precisely, the base space $H^{*}(X)\times\mathbb{C}$ should be considered as a formal supermanifold over $\mathbb{C}[\![Q]\!]$ whose ring of regular functions is $\mathbb{C}[z][\![Q,\tau]\!]$. The flat connection can be also extended in the direction of Novikov variables $Q$: we define for $\xi\in H^{2}(X)$
(2.5) $\nabla_{\xi Q\partial_{Q}}=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\tau})$
where $\xi Q\partial_{Q}$ denotes the derivation of $\mathbb{C}[\![Q]\!]$ given by $(\xi Q\partial_{Q})Q^{d}=(\xi\cdot d)Q^{d}$. In algebraic terms, the quantum connection defines pairwise supercommuting operators
$\nabla_{\tau^{i}},\ \nabla_{z\partial_{z}},\ \nabla_{\xi Q\partial_{Q}}\colon H^{*}(X)[z][\![Q,\tau]\!]\to z^{-1}H^{*}(X)[z][\![Q,\tau]\!].$
The quantum connection admits the fundamental solution $M_{X}(\tau)$ defined as follows (see [19, §1], [44, Proposition 2], [25, Proposition 2.4], [6, Proposition 3.1] for the properties below):
(2.6) $(M_{X}(\tau)\phi_{i},\phi_{j})_{X}=(\phi_{i},\phi_{j})_{X}+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0)\end{subarray}}\left\langle\phi_{i},\tau,\dots,\tau,\frac{\phi_{j}}{z-\psi}\right\rangle_{0,n+2,d}^{X}\frac{Q^{d}}{n!}$
where $\phi_{j}/(z-\psi)$ in the correlator should be expanded as $\sum_{b=0}^{\infty}\phi_{j}\psi^{b}/z^{b+1}$. The solution $M_{X}(\tau)$ belongs to $\operatorname{End}(H^{*}(X))\otimes\mathbb{C}[z^{-1}][\![Q,\tau]\!]$ and satisfies the following differential equations:
(2.7) $\begin{split}M_{X}(\tau)\circ\nabla_{\tau^{i}}&=\partial_{\tau^{i}}\circ M_{X}(\tau),\\ M_{X}(\tau)\circ\nabla_{\xi Q\partial_{Q}}&=\left(\xi Q\partial_{Q}+z^{-1}\xi\right)\circ M_{X}(\tau),\\ M_{X}(\tau)\circ\nabla_{z\partial_{z}}&=\left(z\partial_{z}-z^{-1}c_{1}(TX)+\mu_{X}\right)\circ M_{X}(\tau)\end{split}$
where $\xi$ and $c_{1}(X)$ on the right-hand side act by the cup product. Let $P_{X}$ denote the $z$-sesquilinear pairing on $H^{*}(X)\otimes\mathbb{C}[z,z^{-1}][\![Q,\tau]\!]$ induced by the Poincaré pairing
(2.8) $P_{X}(f,g)=(f(-z),g(z))_{X}.$
Then the pairing $P_{X}$ is flat for $\nabla$ and $M_{X}$ is isometric with respect to $P_{X}$.
(2.9) $\begin{split}dP_{X}(f,g)&=P_{X}(\nabla f,g)+P_{X}(f,\nabla g)\\ P_{X}(f,g)&=P_{X}(M_{X}(\tau)f,M_{X}(\tau)g)\end{split}$
where $f,g\in H^{*}(X)\otimes\mathbb{C}[z][\![Q,\tau]\!]$. The $J$-function of $X$ is defined to be the image of $1$ under $M_{X}(\tau)$:
(2.10) $\begin{split}J_{X}(\tau)&=M_{X}(\tau)1=1+\sum_{(d,n)\neq(0,0)}\sum_{j=0}^{s}\left\langle 1,\tau,\dots,\tau,\frac{\phi^{j}}{z-\psi}\right\rangle_{0,n+2,d}^{X}\phi_{j}\frac{Q^{d}}{n!}\\ &=1+\frac{\tau}{z}+\sum_{(d,n)\neq(0,0),(0,1)}\sum_{k=0}^{\infty}\sum_{j=0}^{s}\frac{\phi^{j}}{z^{k+2}}\left\langle\tau,\dots,\tau,\phi_{j}\psi^{k}\right\rangle_{0,n+1,d}\frac{Q^{d}}{n!}\end{split}$
where $\{\phi^{j}\}$ is the dual basis of $\{\phi_{i}\}$ such that $(\phi_{i},\phi^{j})_{X}=\delta_{i}^{j}$. We used the String Equation [44, §1.2] in the second line. The other columns $M_{X}(\tau)\phi_{i}$ of $M_{X}(\tau)$ are given by the derivatives $z\partial_{\tau^{i}}J_{X}(\tau)$ of the $J$-function due to the differential equation (2.7).
2.3. The vector bundle case
In this paper, we also consider the quantum cohomology of a non-compact space, the total space of a vector bundle $V\to B$ of rank $r$ over a smooth projective base $B$. We introduce the fibrewise scalar $\mathbb{C}^{\times}$-action on $V$. This induces a $\mathbb{C}^{\times}$-action on the moduli space $V_{0,n,d}$, and the virtual class and the $\psi$-classes have canonical equivariant lifts. The $\mathbb{C}^{\times}$-fixed locus of $V_{0,n,d}$ is identified with $B_{0,n,d}$ and hence is proper. Thus we can define the equivariant Gromov-Witten invariants
$\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle_{0,n,d}^{V}=\int_{[V_{0,n,d}]_{\rm vir}}^{\mathbb{C}^{\times}}\prod_{i=1}^{n}\operatorname{ev}_{i}^{*}(\alpha_{i})\psi_{i}^{k}$
with $\alpha_{1},\dots,\alpha_{n}\in H^{*}_{\mathbb{C}^{\times}}(V)=H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ by the virtual localization formula [22], where $\lambda$ is the $\mathbb{C}^{\times}$-equivariant (or $S^{1}$-equivariant) parameter [^1] In this paper, all $S^{1}$-actions are induced from $\mathbb{C}^{\times}$-actions, so we use the terms ‘$S^{1}$-equivariant’ and ‘$\mathbb{C}^{\times}$-equivariant’ interchangeably.. The equivariant Gromov-Witten invariants take values in $\mathbb{C}[\lambda,\lambda^{-1}]$. In §2.5 below, we describe them as the Gromov-Witten invariants of $B$ twisted by $V$ and the inverse equivariant Euler class.
Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(B)$. This induces a $\mathbb{C}[\lambda]$-basis $\{\phi_{i}\}_{i=0}^{s}$ of $H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ and $\mathbb{C}[\lambda]$-valued coordinates $\tau^{0},\dots,\tau^{s}$ on $H_{S^{1}}^{*}(V)$; $(\tau^{0},\dots,\tau^{s})$ parametrizes a point $\tau=\sum_{i=0}^{s}\tau^{i}\phi_{i}\in H^{*}_{S^{1}}(V)$. We define the quantum product $\star_{\tau}$ on $H^{*}_{S^{1}}(V)$ by the same formula (2.1), replacing the Poincaré pairing and the Gromov-Witten invariants with their equivariant counterparts. Note that the equivariant Poincaré pairing of $V$ is given by the localization formula:
(2.11) $(\alpha,\beta)_{V}:=\int_{B}\alpha\cup\beta\cup e_{\lambda}(V)^{-1}\qquad\text{for $\alpha,\beta\in H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$}$
where $e_{\lambda}(V)=\sum_{i=0}^{r}\lambda^{i}c_{r-i}(V)$ is the equivariant Euler class. In general, the quantum product $\alpha\star_{\tau}\beta$ with $\alpha,\beta\in H^{*}_{S^{1}}(V)$ lies only in the localization $H^{*}_{S^{1}}(V)_{\rm loc}:=H^{*}_{S^{1}}(V)\otimes_{\mathbb{C}[\lambda]}\mathbb{C}[\lambda,\lambda^{-1}]$. As noted in [4, §1.4], we can define $\alpha\star_{\tau}\beta$ without inverting $\lambda$ if $V$ is semi-projective, i.e. projective over an affine variety. In this case, the evaluation map $\operatorname{ev}_{i}$ is proper, and therefore the quantum product can be defined as the push-forward along the evaluation map. We show the following.
Lemma 2.1.
Suppose that $V^{\vee}$ is generated by global sections. Then $V$ is semi-projective. In particular, the quantum product $\star_{\tau}$ is defined on $H^{*}_{S^{1}}(V)[\![Q,\tau]\!]$ without localization.
Proof.
By the assumption, we have a surjection $\mathcal{O}^{\oplus N}\to V^{\vee}$ for some $N>0$. Taking the dual, we find that $V$ is a subbundle of the trivial bundle $B\times\mathbb{C}^{N}$. The closed embedding $V\hookrightarrow B\times\mathbb{C}^{N}$ composed with the projection $B\times\mathbb{C}^{N}\to\mathbb{C}^{N}$ gives a projective morphism $V\to\mathbb{C}^{N}$. Thus $V$ is projective over an affine variety. ∎
Hereafter we shall assume that $V^{\vee}$ is generated by global sections. The quantum connection for $V$ is defined similarly to (2.2) and (2.5) except that the $\tau$-parameter space $H^{*}_{S^{1}}(V)$ is now infinite-dimensional over $\mathbb{C}$. Consider the $\mathbb{C}$-basis $\{\phi_{i}\lambda^{k}\}_{0\leq i\leq s,k\geq 0}$ of $H^{*}_{S^{1}}(V)$ and introduce the $\mathbb{C}$-linear coordinates $\{\tau^{i,k}\}$ dual to $\{\phi_{i}\lambda^{k}\}$ so that $\tau^{i}=\sum_{k=0}^{\infty}\tau^{i,k}\lambda^{k}$. The quantum connection is given by
$\nabla_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\phi_{i}\lambda^{k}\star_{\tau})$
$\nabla_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\tau})$
$\nabla_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{V}\star_{\tau})+\mu_{V}$
where $\xi\in H^{2}(V)=H^{2}(B)$ and $E_{V}$, $\mu_{V}$ are given by
$E_{V}$ $=c_{1}^{S^{1}}(TV)+\sum_{i=0}^{s}\sum_{k\geq 0}\left(1-\frac{\deg\phi_{i}}{2}-k\right)\tau^{i,k}\phi_{i}\lambda^{k}$
$\mu_{V}(\phi_{i}\lambda^{k})$ $=\left(\frac{\deg\phi_{i}}{2}+k-\frac{\dim B+r}{2}\right)\phi_{i}\lambda^{k}$
with $c_{1}^{S^{1}}(TV)=c_{1}(TB)+c_{1}(V)+r\lambda$ the equivariant first Chern class. Note that $\mu_{V}$ contains the derivation $\lambda\partial_{\lambda}$. These operators define maps
$\nabla_{\tau^{i,k}},\ \nabla_{\xi Q\partial_{Q}},\ \nabla_{z\partial_{z}}\colon H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]\to z^{-1}H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]$
where $K[\![\boldsymbol{\tau}]\!]$ denotes the space of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}$ with coefficients in $K$ for a module $K$ (we use the boldface symbol $\boldsymbol{\tau}$ to distinguish infinitely many parameters $\{\tau^{i,k}\}$ from finitely many ones $\{\tau^{i}\}$).
We can define the fundamental solution $M_{V}(\tau)$ for $V$ similarly to (2.6) by replacing the Poincaré pairing and the Gromov-Witten invariants with their equivariant counterparts. It satisfies the properties (2.7), (2.9) with $X$ and $c_{1}(TX)$ there replaced with $V$ and $c_{1}^{S^{1}}(TV)$ respectively. Since $V$ is semi-projective and the evaluation maps are proper, $M_{V}(\tau)$ is defined without localization with respect to $\lambda$, i.e.
(2.12) $M_{V}(\tau)\in\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[\lambda,z^{-1}][\![Q,\tau]\!].$
We note that the coefficient of $M_{V}(\tau)$ in front of a fixed monomial $Q^{d}\tau^{I}$ is a polynomial in $z^{-1}$ (instead of a formal power series in it). This follows from the virtual localization formula: the $\psi$-classes $\psi_{i}$ are nilpotent on the fixed locus $B_{0,n,d}$. These properties ensure that the substitution $\lambda\to p+kz$ in $J_{V}(\tau)=M_{V}(\tau)1$ (in Theorem 1.1) is well-defined.
Remark 2.2.
Some of the quantities here (such as $\star_{\tau}$, $M_{V}(\tau)$, $J_{V}(\tau)$) are defined over $\mathbb{C}[\lambda][\![\tau]\!]$ while others (such as $E_{V}$, $\nabla_{z\partial_{z}}$) are defined over $\mathbb{C}[\lambda][\![\boldsymbol{\tau}]\!]$. Note that we have the natural inclusion $\mathbb{C}[\lambda][\![\tau]\!]\subset\mathbb{C}[\lambda][\![\boldsymbol{\tau}]\!]$ sending $\tau^{i}$ to $\sum_{k\geq 0}\tau^{i,k}\lambda^{k}$. This inclusion is, however, not preserved by $\lambda\partial_{\lambda}$ and hence we need to work with $\mathbb{C}[\lambda][\![\boldsymbol{\tau}]\!]$ when considering $\nabla_{z\partial_{z}}$.
2.4. Givental Lagrangian cones
We briefly review Givental’s symplecitc formalism for genus-zero Gromov-Witten theory, as developed in [18, 7]. The Givental symplectic space of a smooth projective variety $X$ is the infinite-dimensional $\mathbb{C}[\![Q]\!]$-vector space
(2.13) $\mathcal{H}_{X}=H^{*}(X)(\!(z^{-1})\!)[\![Q]\!]$
equipped with the symplectic form
$\Omega(f,g)=-\operatorname{Res}_{z=\infty}(f(-z),g(z))_{X}dz.$
The Givental space $\mathcal{H}_{X}$ has a canonical decomposition $\mathcal{H}_{X}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ into maximally isotropic subspaces $\mathcal{H}_{\pm}$, where
$\mathcal{H}_{+}=H^{*}(X)[z][\![Q]\!],\quad\mathcal{H}_{-}=z^{-1}H^{*}(X)[\![z^{-1}]\!][\![Q]\!].$
The symplectic form $\Omega$ identifies $\mathcal{H}_{-}$ with the dual space of $\mathcal{H}_{+}$, hence $\mathcal{H}_{X}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with the cotangent bundle $T^{*}\mathcal{H}_{+}$ of $\mathcal{H}_{+}$. Let $\mathcal{F}_{X}$ denote the genus-zero descendant Gromov-Witten potential:
$\mathcal{F}_{X}(-z+\mathbf{t}(z))=\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1),(0,2)\end{subarray}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi)\right\rangle_{0,n,d}^{X}\frac{Q^{d}}{n!}$
where $\mathbf{t}(z)=\sum_{n\geq 0}t_{n}z^{n}$ with $t_{n}=\sum_{i=0}^{s}t^{i}_{n}\phi_{i}\in H^{*}(X)$. The argument $-z+\mathbf{t}(z)$ represents a point in the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$ and $\mathcal{F}_{X}$ is regarded as a function on the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$. The Givental Lagrangian cone $\mathcal{L}_{X}^{\rm orig}\subset\mathcal{H}_{X}$ is a submanifold-germ defined as the graph of $d\mathcal{F}_{X}$ in $T^{*}\mathcal{H}_{+}\cong\mathcal{H}_{X}$. It consists of
(2.14) $-z+\mathbf{t}(z)+\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\frac{\partial\mathcal{F}_{X}}{\partial t_{k}^{i}}\\ =-z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{X}\frac{Q^{d}}{n!}$
with $\mathbf{t}(z)\in\mathcal{H}_{+}$. More rigorously, for a set $x=(x_{1},x_{2},\dots)$ of formal parameters, we say that an element $f\in\mathcal{H}_{X}[\![x]\!]$ is a $\mathbb{C}[\![Q,x]\!]$-valued point of $\mathcal{L}_{X}^{\rm orig}$ if $f$ is of the form (2.14) for some $\mathbf{t}(z)\in\mathcal{H}_{+}[\![x]\!]$ with $\mathbf{t}(z)|_{Q=x=0}=0$. We refer to [5, Appendix B] for the definition of $\mathcal{L}_{X}$ as an infinite-dimensional formal scheme over $\mathbb{C}[\![Q]\!]$. The nilpotence [^2] This may fail in equivariant theory, and this is a reason why we allow infinitely many negative powers of $z$ in general. However we do not deal with such cases in this paper. of the $\psi$-class implies that $\mathcal{L}_{X}^{\rm orig}$ is contained in the following polynomial form
(2.15) $\mathcal{H}_{X,\rm pol}=H^{*}(X)\otimes\mathbb{C}[z,z^{-1}][\![Q]\!],$
i.e. a $\mathbb{C}[\![Q,x]\!]$-valued point on $\mathcal{L}_{X}^{\rm orig}$ is contained in $\mathcal{H}_{X,\rm pol}[\![x]\!]$.
Example 2.3.
After multiplying by $z$ and flipping the sign of $z$, the $J$-function $J_{X}(\tau)$ in (2.10) lies on the $\mathcal{L}_{X}^{\rm orig}$:
$zJ_{X}(\tau)|_{z\to-z}\in\mathcal{L}_{X}^{\rm orig}.$
It defines a $\mathbb{C}[\![Q,\tau]\!]$-valued point on $\mathcal{L}_{X}^{\rm orig}$.
Notation 2.4.
The $J$-function and the Givental cone have opposite sign conventions for $z$. In the following discussion, to ease the notation, we shall instead use the Givental cone $\mathcal{L}_{X}:=\mathcal{L}_{X}^{\rm orig}|_{z\to-z}$ with the sign of $z$ flipped, so that $zJ_{X}(\tau)\in\mathcal{L}_{X}$.
As the name suggests, $\mathcal{L}_{X}$ is a cone, i.e. the dilation vector field on $\mathcal{H}_{X}$ (the vector field assigning $f$ to any point $f\in\mathcal{H}_{X}$) is tangent to $\mathcal{L}_{X}$. Moreover it satisfies the following “overruled” property [18, Theorem 1]:
the tangent space $T$ of $\mathcal{L}_{X}$ at any point of $\mathcal{L}_{X}$ is tangent to $\mathcal{L}_{X}$ precisely along $zT\subset T$.
Since $\mathcal{L}_{X}$ is a cone, the tangent space $T$ at $f\in\mathcal{L}_{X}$ (regarded as an affine subspace of $\mathcal{H}_{X}$) contains both the origin and $f$. The overruled property says that (1) $T\cap\mathcal{L}_{X}=zT$ and (2) the tangent space of $\mathcal{L}_{X}$ at any point on $zT$ equals $T$. It follows that the Givental cone is a union of semi-infinite subspaces $zT$ over all tangent spaces $T$. Moreover, the tangent spaces to $\mathcal{L}_{X}$ form a finite-dimensional family: every tangent space equals the tangent space $T_{\tau}$ at $zJ_{X}(\tau)$ for some $\tau$, and $T_{\tau}$ is a free $\mathbb{C}[z][\![Q]\!]$-module generated by the derivatives $z\partial_{\tau^{i}}J_{X}(\tau)=M_{X}(\tau)\phi_{i}$. Hence we have
(2.16) $\mathcal{L}_{X}=\bigcup_{\tau}zM_{X}(\tau)\mathcal{H}_{+}.$
A precise meaning of this equality is as follows: any $\mathbb{C}[\![Q,x]\!]$-valued point on $\mathcal{L}_{X}$ can be written as $zM_{X}(\tau)f$ for some $\tau\in H^{*}(X)\otimes\mathbb{C}[\![Q,x]\!]$ and $f\in\mathcal{H}_{+}[\![x]\!]$ such that $\tau|_{Q=x=0}=0$ and $f|_{Q=x=0}=1$.
2.4.1. The vector bundle case
We explain the Givental cone associated with equivariant Gromov-Witten invariants of the vector bundle $V\to B$. We again assume that $V^{\vee}$ is generated by global sections so that $V$ is semi-projective (see Lemma 2.1). The Givental symplectic space $\mathcal{H}_{V}$ for $V$ is defined similarly to (2.13) by replacing $H^{*}(X)$ with $H^{*}_{S^{1}}(V)$.
$\mathcal{H}_{V}=H^{*}_{S^{1}}(V)(\!(z^{-1})\!)[\![Q]\!]=H^{*}(B)\otimes\mathbb{C}[\lambda](\!(z^{-1})\!)[\![Q]\!].$
The genus-zero descendant Gromov-Witten potential $\mathcal{F}_{V}$ for $V$, defined in terms of the equivariant Gromov-Witten invariants of $V$, takes values in $\mathbb{C}[\lambda,\lambda^{-1}][\![Q]\!]$. The properness of the evaluation maps implies that the points (2.14) on the Givental cone with $X=V$ do not contain negative powers of $\lambda$ whenever the input $\mathbf{t}(z)$ lies in $\mathcal{H}_{+}=H^{*}_{S^{1}}(V)[z][\![Q]\!]$. Hence we get the Givental cone $\mathcal{L}_{V}$ for $V$ as a subset of $\mathcal{H}_{V}$. To be precise, with notation as in §2.3, a $\mathbb{C}[\lambda][\![Q,x]\!]$-valued point of $\mathcal{L}_{V}$ is a point of the form
$z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{e_{\lambda}(V)\phi^{i}}{z^{k+1}}\left\langle\mathbf{t}(-\psi),\dots,\mathbf{t}(-\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{V}\frac{Q^{d}}{n!}$
for some $\mathbf{t}(z)\in\mathcal{H}_{+}[\![x]\!]=H^{*}_{S^{1}}(V)[z][\![Q,x]\!]$ with $\mathbf{t}(z)|_{Q=x=0}=0$, where $\{\phi^{i}\}$ is a basis of $H^{*}(B)$ dual to $\{\phi_{i}\}$ in the sense that $(\phi_{i},\phi^{j})_{B}=\delta_{i}^{j}$: hence $(\phi_{i},e_{\lambda}(V)\phi^{j})_{V}=\delta_{i}^{j}$. As discussed at the end of §2.3, the virtual localization formula furthermore shows that the Givental cone $\mathcal{L}_{V}$ is contained in the following polynomial form $\mathcal{H}_{V,\rm pol}\subset\mathcal{H}_{V}$:
(2.17) $\mathcal{H}_{V,\rm pol}=H^{*}_{S^{1}}(V)\otimes_{\mathbb{C}[\lambda]}\mathbb{C}[\lambda,z,z^{-1}][\![Q]\!].$
2.5. Twisted Gromov-Witten invariants and quantum Riemann-Roch theorem
We review the twisted Gromov-Witten invariants and the quantum Riemann-Roch (QRR) theorem due to Coates-Givental [7]. The data of the twist are given by a vector bundle and a characteristic class. Let $\mathbf{c}$ denote the universal multiplicative characteristic class with parameters $\mathbf{s}=(s_{0},s_{1},\dots)$:
$\mathbf{c}(\cdot)=\exp\left(\sum_{k=0}^{\infty}s_{k}\operatorname{ch}_{k}(\cdot)\right).$
For a vector bundle $W\to X$, we introduce the $(W,\mathbf{c})$-twisted Gromov-Witten invariants of $X$ as
$\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle^{X,(W,\mathbf{c})}_{0,n,d}=\int_{[X_{0,n,d}]_{\rm vir}}\left(\prod_{i=1}^{n}\operatorname{ev}_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}\right)\cup\mathbf{c}(W_{0,n,d})$
where $W_{0,n,d}=\mathbb{R}\pi_{*}f^{*}W\in K^{0}(X_{0,n,d})$ is the complex of vector bundles constructed by the following universal family of stable maps:
${C_{0,n,d}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{f}$${X}$${X_{0,n,d}}$
We can similarly define the twisted versions of quantum cohomology, quantum connection, fundamental solution, $J$-function, Givental symplectic space and Givental cone, by using the twisted Gromov-Witten invariants and the twisted Poincaré pairing
(2.18) $(\alpha,\beta)_{X,(W,\mathbf{c})}=\int_{X}\alpha\cup\beta\cup\mathbf{c}(W).$
In this paper, the twisted Gromov-Witten invariants appear mainly in the following two cases.
  • (a)
    $W$ is convex, i.e. we have $H^{1}(C,f^{*}W)=0$ for any genus-zero stable maps $C\to X$ and $\mathbf{c}$ is the equivariant Euler class $e_{\lambda}$. This corresponds to the choice of parameters
    $s_{k}=\begin{cases}\log\lambda&k=0;\\ (-1)^{k-1}(k-1)!\lambda^{-k}&k\geq 1.\end{cases}$
  • (b)
    $W^{\vee}$ is generated by global sections and $\mathbf{c}$ is the inverse equivariant Euler class $e_{\lambda}^{-1}$. This corresponds to the negative of the parameters $s_{k}$ given above.
In the case (a), the $(W,e_{\lambda})$-twisted Gromov-Witten invariants are closely related to the Gromov-Witten invariants of the zero-locus $Z\subset X$ of a regular section of $W$. The convexity assumption implies that $W_{0,n,d}$ is represented by a vector bundle. Hence we can take the non-equivariant limit $\lambda\to 0$ of the $(W,e_{\lambda})$-twisted invariants. The functoriality of virtual classes [33] implies
$\lim_{\lambda\to 0}\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle_{0,n,d}^{X,(W,e_{\lambda})}=\sum_{i_{*}d^{\prime}=d}\left\langle i^{*}(\alpha_{1})\psi^{k_{1}},\dots,i^{*}(\alpha_{n})\psi^{k_{n}}\right\rangle_{0,n,d^{\prime}}^{Z}$
where $i\colon Z\to X$ is the inclusion map. Moreover, the fundamental solution $M_{X,(W,e_{\lambda})}$ in the $(W,e_{\lambda})$-twisted theory satisfies (see [44, §2.1], [26, Proposition 2.4])
$\lim_{\lambda\to 0}i^{*}M_{X,(W,e_{\lambda})}(\tau)\phi=M_{Z}(i^{*}\tau)i^{*}\phi\Bigr{|}_{H_{2}(Z,\mathbb{Z})\to H_{2}(X,\mathbb{Z})}$
for $\tau,\phi\in H^{*}(X)$, where the subscript $H_{2}(Z,\mathbb{Z})\to H_{2}(X,\mathbb{Z})$ means to replace the Novikov variable $Q^{d}$ with $Q^{i_{*}d}$ for $d\in H_{2}(Z,\mathbb{Z})$. This means that points on the $(W,e_{\lambda})$-twisted Givental cone yield, after applying $i^{*}$ and taking the non-equivariant limit, points on the Givental cone of $Z$.
The case (b) corresponds to the equivariant Gromov-Witten invariants of the total space of $W$. The virtual localization formula [22] implies that the equivariant Gromov-Witten invariants of the vector bundle $V\to B$ equal the $(V,e_{\lambda}^{-1})$-twisted Gromov-Witten invariants of $B$:
$\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle^{V}_{0,n,d}=\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle^{B,(V,e_{\lambda}^{-1})}_{0,n,d}.$
We now state the (genus-zero) QRR theorem of Coates-Givental [7]. Let $\mathcal{H}_{X,(W,\mathbf{c})}=H^{*}(X)(\!(z^{-1})\!)[\![Q]\!][\![\mathbf{s}]\!]$ denote the Givental space for the $(W,\mathbf{c})$-twisted theory: it is equipped with the symplectic form induced by the twisted Poincaré pairing (2.18). Define the symplectic operator [^3] In view of our convention from Notation 2.4, we flipped the sign of $z$ from the original definition. $\Delta_{(W,\mathbf{c})}\colon\mathcal{H}_{X}[\![\mathbf{s}]\!]\to\mathcal{H}_{X,(W,\mathbf{c})}$ as
(2.19) $\Delta_{(W,\mathbf{c})}=\exp\left(\sum_{l,m\geq 0}s_{l+m-1}\frac{B_{m}}{m!}\operatorname{ch}_{l}(W)(-z)^{m-1}\right)$
where we set $s_{-1}=0$ and $\{B_{m}\}_{m=0}^{\infty}$ are the Bernoulli numbers defined by $\sum_{m=0}^{\infty}\frac{B_{m}}{m!}x^{m}=x/(e^{x}-1)$; we have $B_{0}=1,B_{1}=-\frac{1}{2},B_{2}=\frac{1}{6}$, etc., and $B_{2k+1}=0$ for all $k\geq 1$. Let $\mathcal{L}_{X,(W,\mathbf{c})}\subset\mathcal{H}_{X,(W,\mathbf{c})}$ denote the Givental cone for the $(W,\mathbf{c})$-twisted theory.
Theorem 2.5 ([7, Corollary 4]).
We have $\mathcal{L}_{X,(W,\mathbf{c})}=\Delta_{(W,\mathbf{c})}\mathcal{L}_{X}$.
Remark 2.6.
In the case of $e_{\lambda}$ or $e_{\lambda}^{-1}$-twist, we have $s_{0}=\pm\log\lambda$ whereas $s_{k}$’s with positive $k$ are proportional to $\lambda^{-k}$. The existence of $\log\lambda$ makes it difficult to treat the $e_{\lambda}^{\pm}$-twisted Givental cone formally in terms of the $\lambda^{-1}$-adic topology. We deal with the $e_{\lambda}^{\pm}$-twist in the following two steps: first we consider the twist by $\tilde{e}_{\lambda}^{\pm}$, where
$\tilde{e}_{\lambda}(W)=\sum_{i\geq 0}\lambda^{-i}c_{i}(W)=\lambda^{-\operatorname{rank}W}e_{\lambda}(W),$
and get the $\tilde{e}_{\lambda}^{\pm}$-twisted Givental cone $\widetilde{\mathcal{L}}$ defined over $\mathbb{C}[\![Q,\lambda^{-1}]\!]$; then points on the $e_{\lambda}^{\pm}$-twisted cone arise from points on $\widetilde{\mathcal{L}}$ by the change of variables $Q^{d}\to Q^{d}\lambda^{\pm c_{1}(W)\cdot d}$ (this follows easily from the definition of points on the cone, together with $\operatorname{rank}W_{0,n,d}=\operatorname{rank}W+c_{1}(W)\cdot d$ and the change of the twisted Poincaré pairings).
2.6. Moving points on the Givental cone via differential operators
We review a well-known technique to construct points on the Givental cone via the action of differential operators, and show that any “miniversal” slices on the cone arise from the $J$-function in this way. This plays a key role in the proof of the main theorems. Let $x=(x_{1},x_{2},\dots)$, $y=(y_{1},y_{2},\dots)$ be two sets of formal parameters. We consider a differential operator $F=F(x,z\partial_{x},z)$ such that
$F\in\mathbb{C}[z][\![x]\!]\langle z\partial_{x_{1}},z\partial_{x_{2}},\dots\rangle[\![y,Q]\!],\qquad F|_{y=Q=0}=0.$
The second condition ensures that $\exp(F/z)$ converges in the $(Q,y)$-adic topology.
Lemma 2.7 ([7, §8], [5, §4.2]).
The action of the differential operator $\exp(F/z)$ preserves $\mathbb{C}[\![Q,x,y]\!]$-valued points on $\mathcal{L}_{X}$.
Proof.
This is well-known, at least in the case where $F$ is a constant differential operator (i.e. do not depend on $x$). We include a brief discussion. The overruled property implies that $z^{-1}Ff$ for a $\mathbb{C}[\![Q,x,y]\!]$-valued point $f$ of $\mathcal{L}_{X}$ lies in the tangent space of $\mathcal{L}_{X}$ at $f$. Thus the assignment $f\mapsto z^{-1}Ff$ defines a vector field on the mapping scheme [^4] In [5, Appendix B], the formal scheme $\mathcal{L}_{X}$ was given as a functor from topological $\mathbb{C}[\![Q]\!]$-algebras to sets. It is represented by a certain formal power series ring over $\mathbb{C}[\![Q]\!]$. The mapping scheme here is given by the functor $R\mapsto\mathcal{L}_{X}(R[\![x,y]\!])$, which is again represented by a formal power series ring over $\mathbb{C}[\![Q]\!]$. $\operatorname{Map}(\operatorname{Spf}(\mathbb{C}[\![Q,x,y]\!]),\mathcal{L}_{X})$ and $\exp(F/z)f$ is the time-one flow of the vector field. ∎
We also prove a sort of converse statement. Let $\tau=\sum_{i=0}^{s}\tau^{i}\phi_{i}$ be a parameter in $H^{*}(X)$ as before. Recall from Example 2.3 that $zJ_{X}(\tau)$ is a $\mathbb{C}[\![Q,\tau]\!]$-valued point on $\mathcal{L}_{X}$. We say that a $\mathbb{C}[\![Q,\tau,y]\!]$-valued point $f$ on the Givental cone $\mathcal{L}_{X}$ is a miniversal slice if it satisfies
$f|_{Q=y=0}=z+\tau+O(z^{-1}).$
By the definition of $\mathcal{L}_{X}$ (see (2.14)), this is equivalent to $f|_{Q=y=0}=ze^{\tau/z}$.
Lemma 2.8.
Any miniversal slice of the Givental cone $\mathcal{L}_{X}$ can be obtained from $zJ_{X}(\tau)$ by applying a differential operator of the form $\exp(F/z)$ for some $F=F(\tau,z\partial_{\tau},z)$ such that $F\in\sum_{i=0}^{s}\mathbb{C}[z][\![Q,\tau,y]\!]z\partial_{\tau^{i}}$ and $F|_{Q=y=0}=0$.
Proof.
Let $f\in\mathcal{L}_{X}(\mathbb{C}[\![Q,\tau,y]\!])$ be a miniversal slice. By (2.16), we have $f=zM_{X}(\sigma)g$ for some $\sigma\in H^{*}(X)[\![Q,\tau,y]\!]$, $g\in H^{*}(X)[z][\![Q,\tau,y]\!]$ with $\sigma|_{Q=\tau=y=0}=0$, $g|_{Q=\tau=y=0}=1$. The miniversality assumption implies that $\sigma|_{Q=y=0}=\tau$ and $g|_{Q=y=0}=1$. We claim that there exists a vector field $x\in\sum_{i=0}^{s}\mathbb{C}[\![Q,\tau,y]\!]\partial_{\tau^{i}}$ such that
$\exp(x)\sigma=\tau,\qquad x|_{Q=y=0}=0.$
This can be shown easily by induction on powers of $Q$ and $y$ (see [23, Theorem 3.17] for a related result). Then we have
$\exp(x)f=\exp(x)\left(zM_{X}(\sigma)g\right)=zM_{X}(\tau)g^{\prime}$
with $g^{\prime}=\exp(x)g$. We have $g^{\prime}|_{Q=y=0}=1$. Next we claim that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that
(2.20) $\exp\left(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}}\right)g^{\prime}=1,\qquad w_{i}|_{Q=y=0}=0.$
We construct $w_{i}$ by induction on powers of $Q$ and $y$. Take an ample class $\omega$ of $X$, and consider the additive valuation $v\colon\mathbb{C}[\![Q,y]\!]\to\mathbb{R}\cup\{\infty\}$ given by $v(Q^{d})=\omega\cdot d$ and $v(y_{i})=i$. Suppose by induction that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that the equation (2.20) holds modulo terms with valuation $>v_{0}$ for some $v_{0}\in v(\mathbb{C}[\![Q,y]\!])$. Let $v_{1}>v_{0}$ be the next possible value of the valuation $v$. Write $\exp(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}})g^{\prime}=1-\sum_{i=0}^{s}\delta_{i}\phi_{i}$ for some $\delta_{i}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ with $v(\delta_{i})\geq v_{1}$. Using the fact that $z\nabla_{\tau^{i}}1=\phi_{i}$, we have that $\exp\left(\sum_{i=0}^{s}(w_{i}+\delta_{i})z\nabla_{\tau^{i}}\right)g^{\prime}$ equals 1 modulo terms with valuation $>v_{1}$. This completes the induction and the proof of the claim. Setting $G=\sum_{i=0}^{s}w_{i}z\partial_{\tau^{i}}$, we have
$\exp(G)\exp(x)f$ $=\exp(G)zM_{X}(\tau)g^{\prime}=zM_{X}(\tau)\exp\left(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}}\right)g^{\prime}$
$=zM_{X}(\tau)1=zJ_{X}(\tau)$
where we used $\partial_{\tau^{i}}\circ M_{X}(\tau)=M_{X}(\tau)\circ\nabla_{\tau^{i}}$. Since $\sum_{i=0}^{s}\mathbb{C}[z][\![Q,\tau,y]\!]\partial_{\tau^{i}}$ is closed under commutator, we can rewrite the left-hand side by the Baker-Campbell-Hausdorff formula and obtain $F\in\sum_{i=0}^{s}\mathbb{C}[z][\![Q,\tau,y]\!]z\partial_{\tau^{i}}$ such that $\exp(-F/z)f=zJ_{X}(\tau)$ and $F|_{Q=y=0}=0$, as claimed. ∎
2.7. Shift operators
Shift operators of equivariant parameters have been introduced by Okounkov-Pandharipande [42] and studied in [41, 2, 27, 39, 21]. They can be viewed as lifts of the Seidel representation [48] on quantum cohomology algebras to quantum $D$-modules. In this section we review basic properties of shift operators in the case of the vector bundle $V\to B$. As before, we assume that $V^{\vee}$ is generated by global sections and thus $V$ is semi-projective by Lemma 2.1.
In this section, we work with the $\mathbb{C}$-basis $\{\phi_{i}\lambda^{k}\}$ and the associated $\mathbb{C}$-linear coordinates $\{\tau^{i,k}\}$ of $H^{*}_{S^{1}}(V)$ as in §2.3. The shift operator shall be defined over the ring $\mathbb{C}[\![\boldsymbol{\tau}]\!]$ of formal power series in $\{\tau^{i,k}\}$. Since $V$ is semiprojective and the fibrewise scalar $S^{1}$-action is seminegative [^5] On the other hand, the inverse $S^{1}$-action is not seminegative: this is the reason why $\mathbb{S}(\tau)^{-1}$ is not defined without localization. (in the sense of [27, Definition 3.3]), the $S^{1}$-action defines the shift operator [27, Definition 3.9]
$\mathbb{S}(\tau)\colon H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]\to H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!].$
This operator satisfies the following properties:
(2.21) $\begin{split}\mathbb{S}(\tau)(f(\lambda,z)x)&=f(\lambda-z,z)\mathbb{S}(\tau)x\\ [\nabla_{\tau^{i,k}},\mathbb{S}(\tau)]&=[\nabla_{\xi Q\partial_{Q}},\mathbb{S}(\tau)]=[\nabla_{z\partial_{z}},\mathbb{S}(\tau)]=0\\ M_{V}(\tau)\circ\mathbb{S}(\tau)&=\mathcal{S}\circ M_{V}(\tau)\end{split}$
where $f(\lambda,z)\in\mathbb{C}[\lambda,z][\![Q,\boldsymbol{\tau}]\!]$, $x\in H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]$, $\xi\in H^{2}(B)$ and $\mathcal{S}=e_{\lambda}(V)e^{-z\partial_{\lambda}}$ is the shift operator (1.1) on the polynomial Givental space $\mathcal{H}_{V,\rm pol}=H^{*}_{S^{1}}(V)\otimes_{\mathbb{C}[\lambda]}\mathbb{C}[\lambda,z^{\pm}][\![Q]\!]$. Here $e^{-z\partial_{\lambda}}$ denotes the operator sending $f(\lambda)$ to $f(\lambda-z)$. The third property of (2.21) is in [27, Theorem 3.14] and the second follows from the third together with the differential equations (2.7) for $M_{V}$ and the commutativity
$[\partial_{\tau^{i,k}},\mathcal{S}]=[\xi Q\partial_{Q}+z^{-1}\xi,\mathcal{S}]=[z\partial_{z}+z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V},\mathcal{S}]=0.$
The shift operator is compatible with the pairing $P_{V}$ (2.8) in the sense that
(2.22) $e^{-z\partial_{\lambda}}P_{V}(\mathbb{S}(\tau)f,g)=P_{V}(f,\mathbb{S}(\tau)g)$
with $f,g\in H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]$. This follows from the fact (2.9) that $M_{V}(\tau)$ is isometric with respect to $P_{V}$, the intertwining property (2.21) between $\mathbb{S}(\tau)$ and $\mathcal{S}$ and the identity
$e^{-z\partial_{\lambda}}P_{V}(\mathcal{S}f,g)=P_{V}(f,\mathcal{S}g)$
that can be verified easily. The operator $\mathbb{S}(\tau)$ is invertible over the localization $\mathbb{C}(\lambda,z)[\![Q,\boldsymbol{\tau}]\!]$ since the same holds for $\mathcal{S}$. The precise image of $\mathbb{S}(\tau)$ is given in the following remark.
Remark 2.9.
Using (2.9) and (2.21), we have
$P_{V}(\phi_{i},\mathbb{S}(\tau)\phi_{j})=P_{V}(M_{V}(\tau)\phi_{i},\mathcal{S}M_{V}(\tau)\phi_{j})=\int_{B}(M_{V}(\tau,-z)\phi_{i})\cup e^{-z\partial_{\lambda}}(M_{V}(\tau,z)\phi_{j})$
where we write $M_{V}(\tau,z)$ for $M_{V}(\tau)$. The left-hand side belongs to $\mathbb{C}[\lambda,\lambda^{-1},z][\![Q,\boldsymbol{\tau}]\!]$ whereas the right-hand side belongs to $\mathbb{C}[\lambda,z,z^{-1}][\![Q,\boldsymbol{\tau}]\!]$ by (2.12). Thus $P_{V}(\phi_{i},\mathbb{S}(\tau)\phi_{j})$ belongs to $\mathbb{C}[\lambda,z][\![Q,\boldsymbol{\tau}]\!]$; they define an invertible matrix because $P_{V}(\phi_{i},\mathbb{S}(\tau)\phi_{j})|_{Q=\tau=0}=\int_{B}\phi_{i}\cup\phi_{j}$. This implies that the image of $\mathbb{S}(\tau)$ is precisely
$\mathbb{S}(\tau)(H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!])=e_{\lambda}(V)\cdot H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!].$
The image of $\mathbb{S}(\tau)$ is a sub $D$-module since $\mathbb{S}(\tau)$ commutes with the connection. This fact is related to the quantum Serre [7, Corollary 9]: the above submodule can be identified with the quantum $D$-module of the $(V^{\vee},e_{\lambda})$-twisted theory [30, Theorems 2.11, 3.14].
3. Proof of the mirror theorem
In this section we prove Theorem 1.1 (and a more general version Theorem 3.3 below). Let $\mathbb{P}(V)\to B$ be the projective bundle associated with a vector bundle $V\to B$ of rank $r\geq 2$. Let $p=c_{1}(\mathcal{O}(1))$ be the first Chern class of the relative $\mathcal{O}(1)$ over $\mathbb{P}(V)$. This class defines the following splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})$:
(3.1) $H_{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}\oplus\operatorname{Ker}p\cong\mathbb{Z}\oplus H_{2}(B,\mathbb{Z}).$
Notation 3.1.
We write $Q$ for the Novikov variable for $B$, and $q$ for the Novikov variable for $\mathbb{P}(V)$ associated with the class of a line in a fibre. Via the above splitting (3.1), we identify $(q,Q)$ with the Novikov variables for $\mathbb{P}(V)$: a curve class $(k,d)\in\mathbb{Z}\oplus H_{2}(B,\mathbb{Z})\cong H_{2}(\mathbb{P}(V),\mathbb{Z})$ corresponds to the monomial $q^{k}Q^{d}$.
As before, we consider the fibrewise scalar $S^{1}$-action on $V$ and denote by $\lambda$ the $S^{1}$-equivariant parameter. Let $\{\phi_{i}\}_{i=0}^{s}$ be a basis of $H^{*}(B)$. Via the canonical isomorphism $H^{*}_{S^{1}}(V)\cong H^{*}(B)\otimes\mathbb{C}[\lambda]$, $\{\phi_{i}\}_{i=0}^{s}$ gives a basis of $H^{*}_{S^{1}}(V)$ over $\mathbb{C}[\lambda]$. Let $(\tau^{0},\dots,\tau^{s})\mapsto\tau=\sum_{i=0}^{s}\tau^{i}\phi_{i}$ be the linear coordinates on $H^{*}_{S^{1}}(V)$ associated with $\{\phi_{i}\}_{i=0}^{s}$.
3.1. Outline
Since $V^{\vee}$ is globally generated, we have a surjection $\mathcal{O}^{\oplus N}\to V^{\vee}$ for some $N$. Dualizing, we get an exact sequence of vector bundles over $B$:
(3.2) $0\to V\to\mathcal{O}^{\oplus N}\to\mathcal{Q}\to 0$
The inclusion $V\hookrightarrow\mathcal{O}^{\oplus N}$ induces an embedding $\mathbb{P}(V)\hookrightarrow B\times\mathbb{P}^{N-1}$. Since $\mathcal{Q}$ is a quotient of the trivial bundle $\mathcal{O}^{\oplus N}$, we have a natural section $s$ of $\mathcal{Q}(1)=\pi_{1}^{*}\mathcal{Q}\otimes\pi_{2}^{*}\mathcal{O}(1)$ over $B\times\mathbb{P}^{N-1}$ given by the composition $\pi_{2}^{*}\mathcal{O}(-1)\to\mathcal{O}_{B\times\mathbb{P}^{N-1}}^{\oplus N}\to\pi_{1}^{*}\mathcal{Q}$, where $\pi_{1}\colon B\times\mathbb{P}^{N-1}\to B$ and $\pi_{2}\colon B\times\mathbb{P}^{N-1}\to\mathbb{P}^{N-1}$ are the projections. The section $s$ is transverse to the zero-section and cuts out $\mathbb{P}(V)$ as the zero-locus. Clearly, $\mathcal{Q}(1)$ is generated by global sections and therefore $\mathcal{Q}(1)$ is convex.
Our starting point is the fact that Theorem 3.3 holds for the trivial bundle $B\times\mathbb{C}^{N}$: the $J$-function of $B\times\mathbb{P}^{N-1}$ at parameter $\tau+tp$ (with $\tau\in H^{*}(B)$) is given by
$J_{B\times\mathbb{P}^{N-1}}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}J_{B}(\tau).$
This can be viewed as a special case of the Elezi-Brown mirror theorem [3]. As discussed in §2.5, since $\mathcal{Q}(1)$ is convex, the Gromov-Witten invariants of $\mathbb{P}(V)$ can be computed as the non-equivariant limits of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Gromov-Witten invariants of $B\times\mathbb{P}^{N-1}$. We shall compute a slice of the Givental cone of the $(\mathcal{Q}(1),e_{\lambda})$-twisted theory and take its non-equivariant limit.
3.2. Informal proof
We first give a slightly informal proof, ignoring the issue on adic convergence. A rigorous proof follows in the next section, but it is only technically complicated and we recommend the reader to begin with the discussion in this section. We introduce a modified QRR operator:
(3.3) $\Delta^{\lambda}_{W}:=e^{\operatorname{rank}(W)(\lambda\log\lambda-\lambda)/z}\Delta_{(W,e_{\lambda}^{-1})}$
where $\Delta_{(W,e_{\lambda}^{-1})}$ is the QRR operator (2.19) of the inverse Euler twist. This arises from the asymptotic expansion (Stirling’s approximation) of the $\Gamma$-function:
$\Delta^{\lambda}_{W}\sim\prod_{\delta}\sqrt{\frac{z}{2\pi}}z^{\frac{\lambda+\delta}{z}}\Gamma\left(1+\frac{\lambda+\delta}{z}\right)$
where $\delta$ ranges over the Chern roots of $W$. By the String Equation, the Givental cone is preserved by the action of $e^{g(z)/z}$, with $g(z)$ a nonnegative power series of $z$ with scalar coefficients. Hence we can use $\Delta_{W}^{\lambda}$ instead of $\Delta_{(W,e_{\lambda}^{-1})}$ in the QRR Theorem 2.5. The identity $\Gamma(1+x)=x\Gamma(x)$ implies that
(3.4) $\Delta_{W}^{\lambda+kz}/\Delta_{W}^{\lambda}=\prod_{c=1}^{k}\prod_{\delta}(\lambda+\delta+cz).$
The exact sequence (3.2) implies $\Delta_{V}^{\lambda}\Delta_{\mathcal{Q}}^{\lambda}=\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$. Since $\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$ preserves the Givental cone (again by the String Equation), we have $\Delta_{\mathcal{Q}}^{\lambda}\mathcal{L}_{B,(V,e_{\lambda}^{-1})}=\mathcal{L}_{B}$ by the QRR Theorem 2.5. Let $J_{V}^{\lambda}(\tau)$ be the equivariant $J$-function of $V$ so that $zJ_{V}^{\lambda}(\tau)\in\mathcal{L}_{B,(V,e_{\lambda}^{-1})}$. Then $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)$ lies in $\mathcal{L}_{B}$. As in Lemma 2.8, we expect to have a differential operator $F(\lambda)=F(\lambda;\tau,z\partial_{\tau},z)$ in $\tau$ such that
(3.5) $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)=e^{F(\lambda)/z}zJ_{B}(\tau).$
This is not justified by the argument in the proof of Lemma 2.8 due to the lack of an appropriate formal topology; $\Delta^{\lambda}_{\mathcal{Q}}$ contains both positive and negative powers of $\lambda$ and also $\log\lambda$. Let us proceed assuming that we have such an $F(\lambda)$.
Consider the differential operator $\exp(F(\lambda+z\partial_{t})/z)$ in $t$ and $\tau$; here we expand $F(\lambda+z\partial_{t})$ in Taylor series of $z\partial_{t}$. By Lemma 2.7, $e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$ lies in $\mathcal{L}_{B\times\mathbb{P}^{N-1}}$. Therefore
$I^{\lambda}(\tau,t):=(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$
lies in the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$ by the QRR Theorem 2.5. We calculate
$I^{\lambda}(\tau,t)$ $=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+p+kz)/z}J_{B}(\tau)$
$=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}}^{\lambda+p})^{-1}\Delta^{\lambda+p+kz}_{\mathcal{Q}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.5)
(3.6) $=\sum_{k\geq 0}e^{tp/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+p+\epsilon+cz)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.4)
where $\epsilon$ ranges over the Chern roots of $\mathcal{Q}$ in the last line. As discussed in §2.5, $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ lies in the Givental cone $\mathcal{L}_{\mathbb{P}(V)}$ of $\mathbb{P}(V)$, where $i\colon\mathbb{P}(V)\hookrightarrow B\times\mathbb{P}^{N-1}$ is the inclusion. This gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1. Note that $i_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\to H_{2}(B\times\mathbb{P}^{N-1},\mathbb{Z})$ is an isomorphism whenever $r=\operatorname{rank}(V)\geq 2$.
Remark 3.2.
Another point that we glossed over in the above discussion is that the modified Riemann-Roch operator (3.3)
(3.7) $\Delta^{\lambda}_{W}=\prod_{\delta}e^{\frac{(\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta)}{z}+\frac{1}{2}\log(\lambda+\delta)+\sum_{n\geq 2}\frac{B_{n}}{n(n-1)}\left(\frac{z}{\lambda+\delta}\right)^{n-1}}$
arising from the Stirling approximation does not converge in an adic topology, because the exponent contains [^6] This is also the case with the original operator $\Delta_{(W,e_{\lambda}^{-1})}$, but this was not a problem because the coefficient of $z^{-1}$ in $\log(\Delta_{(W,e_{\lambda}^{-1})})$ is nilpotent. both positive and negative powers of $z$. This converges in the standard topology to power series in $z$ and $z^{-1}$ (infinite in both directions) with coefficients in analytic functions of $\lambda$ near $\lambda=e$, but we are not sure whether this fact would be useful. The only properties of $\Delta_{W}^{\lambda}$ we use in the rigorous argument below (e.g. (3.4)) can be interpreted in terms of the well-defined logarithm $\log(\Delta^{\lambda}_{W})\in z^{-1}H^{*}(B)\otimes\mathbb{C}[\lambda,\lambda^{-1},\log\lambda][\![z]\!]$ and the convergence issue does not exist.
3.3. Rigorous proof
We explain how to make the above argument rigorous, using the $\lambda^{-1}$-adic topology. We follow the strategy explained in Remark 2.6. Let $\tilde{e}_{\lambda}(\cdot)=\sum_{i\geq 0}\lambda^{-i}c_{i}(\cdot)$ be the modified equivariant Euler class and let $\widetilde{\Delta}_{W}^{\lambda}:=\Delta_{(W,\tilde{e}_{\lambda}^{-1})}$ denote the QRR operator (2.19) of the $(W,\tilde{e}_{\lambda}^{-1})$-twist. We have $\widetilde{\Delta}_{W}^{\lambda}=1+O(\lambda^{-1})$. We also set $\widetilde{J}_{V}^{\lambda}(\tau):=J_{V}^{\lambda}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$; this does not contain positive powers of $\lambda$ and gives the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory. We have
$\Delta_{W}^{\lambda}$ $=e^{(\operatorname{rank}(W)\theta(\lambda)+c_{1}(W)\log\lambda)/z}\widetilde{\Delta}_{W}^{\lambda}\quad\text{with $\theta(\lambda):=\lambda\log\lambda-\lambda+\frac{\log\lambda}{2}z$}$
$J_{V}^{\lambda}(\tau)$ $=e^{(\log\lambda)(c_{1}(V)-z\partial_{c_{1}(V)})/z}\widetilde{J}_{V}^{\lambda}(\tau)$
where $\partial_{c_{1}(V)}$ means the directional derivative along $c_{1}(V)$ in the $\tau$-space. The second equation follows from the Divisor Equation. Combining these equations, we have
(3.8) $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)\quad\text{with $f(\lambda):=\operatorname{rank}(\mathcal{Q})\theta(\lambda)-(\log\lambda)z\partial_{c_{1}(V)}$.}$
Since $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{\Delta}_{V}^{\lambda}=1$ (by the exact sequence (3.2)), $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)$ is a $\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5. This is a miniversal slice, i.e. $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)|_{Q=\lambda^{-1}=0}=z+\tau+O(z^{-1})$. Therefore, we can apply Lemma 2.8 to this slice, and obtain a differential operator $\widetilde{F}(\lambda)\in\sum_{i=0}^{s}\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]z\partial_{\tau^{i}}$ such that $\widetilde{F}(\lambda)|_{Q=\lambda^{-1}=0}=0$ and
(3.9) $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)=e^{\widetilde{F}(\lambda)/z}J_{B}(\tau).$
Mimicking the previous argument in §3.2, we consider applying the differential operator $e^{f(\lambda+z\partial_{t})/z}e^{\widetilde{F}(\lambda+z\partial_{t})/z}$ to $J_{B\times\mathbb{P}^{N-1}}(\tau,t)$; the first factor $e^{f(\lambda+z\partial_{t})/z}$ was added in view of the difference (3.8) between $\Delta^{\lambda}_{\mathcal{Q}}J_{V}^{\lambda}$ and $\widetilde{\Delta}^{\lambda}_{\mathcal{Q}}\widetilde{J}_{V}^{\lambda}$. To work with $\lambda^{-1}$-power series, we subtract from $f(\lambda+z\partial_{t})$ the first few terms of its Taylor expansion in $z\partial_{t}$ and define:
$f_{\rm reg}(\lambda,z\partial_{t}):=f(\lambda+z\partial_{t})-f(\lambda)-\operatorname{rank}(\mathcal{Q})(\log\lambda)z\partial_{t}.$
This belongs to $\mathbb{C}[z\partial_{t},z\partial_{c_{1}(V)}][\![\lambda^{-1}]\!]$ and $f_{\rm reg}(\lambda,z\partial_{t})|_{\lambda^{-1}=0}=0$. Therefore, Lemma 2.7 and the QRR Theorem 2.5 imply that
$\widetilde{I}^{\lambda}(\tau,t):=(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{f_{\rm reg}(\lambda,z\partial_{t})/z}e^{\widetilde{F}(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$
is a $\mathbb{C}[\![Q,\tau,t,\lambda^{-1}]\!]$-valued point of the $(\mathcal{Q}(1),\tilde{e}_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$. By calculations similar to, but more cumbersome than, the previous section §3.2, we arrive at the following result (cf. (3.6)).
$\widetilde{I}^{\lambda}(\tau,t)=\sum_{k\geq 0}e^{pt/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}\left(1+\frac{p+\epsilon+cz}{\lambda}\right)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$
where $\prod_{\epsilon}$ is the product over the Chern roots $\epsilon$ of $\mathcal{Q}$. For this calculation, we used (3.9) and the following identities:
$\widetilde{\Delta}^{\lambda}_{\mathcal{Q}(1)}$ $=e^{(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+p)-\theta(\lambda)-p\log\lambda)+c_{1}(\mathcal{Q})\log(1+\frac{p}{\lambda}))/z}\widetilde{\Delta}^{\lambda+p}_{\mathcal{Q}},$
$\widetilde{\Delta}_{\mathcal{Q}}^{\lambda+kz}/\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}$ $=e^{-\left(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+kz)-\theta(\lambda))+c_{1}(\mathcal{Q})\log(1+\frac{kz}{\lambda})\right)/z}\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+\epsilon+cz),$
$\ J_{V}^{\lambda+x}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ $=e^{-c_{1}(\mathcal{Q})\log(1+\frac{x}{\lambda})/z}e^{-\log(1+\frac{x}{\lambda})\partial_{c_{1}(V)}}\widetilde{J}_{V}^{\lambda+x}(\tau).$
The first identity follows from $\Delta^{\lambda}_{\mathcal{Q}(1)}=\Delta^{\lambda+p}_{\mathcal{Q}}$, the second from (3.4), and the third from the Divisor Equation. Finally we make the change of variables $Q^{d}\to Q^{d}\lambda^{c_{1}(\mathcal{Q})\cdot d}=Q^{d}\lambda^{-c_{1}(V)\cdot d}$, $q^{k}\to q^{k}\lambda^{\operatorname{rank}(\mathcal{Q})k}$ for $\widetilde{I}^{\lambda}$ and obtain a point $I^{\lambda}$ of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone. This gives the same result as (3.6). The non-equivariant limit $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ exists and gives the $I$-function $I_{\mathbb{P}(V)}$ lying on the Givental cone of $\mathbb{P}(V)$. The proof of Theorem 1.1 is now complete.
3.4. A generalization
In this section we observe that the $J$-function $J_{V}^{\lambda}$ in Theorem 1.1 can be replaced with any other slices of $\mathcal{L}_{V}$ depending polynomially on $\lambda$.
Theorem 3.3.
Let $V\to B$ as in Theorem 1.1. Let $x=(x_{1},x_{2},\dots)$ be a set of formal parameters and let $zI_{V}^{\lambda}$ be a $\mathbb{C}[\lambda][\![Q,x]\!]$-valued point of the Givental cone $\mathcal{L}_{V}$. Define the $H^{*}(\mathbb{P}(V))$-valued function $I_{\mathbb{P}(V)}$ by
$I_{\mathbb{P}(V)}=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}\prod_{\begin{subarray}{c}\text{\rm$\delta$: Chern roots}\\ \text{\rm of $V$}\end{subarray}}(p+\delta+cz)}I_{V}^{p+kz}.$
Then $zI_{\mathbb{P}(V)}$ lies in the Givental cone $\mathcal{L}_{\mathbb{P}(V)}$ of $\mathbb{P}(V)$.
Proof.
We can write $I_{V}^{\lambda}=M_{V}(\tau(\lambda))f(\lambda)$ for some $\tau(\lambda)\in H^{*}(B)[\lambda][\![Q,x]\!]$ and $f(\lambda)\in H^{*}(B)[\lambda,z][\![Q,x]\!]$ with $\tau(\lambda)|_{Q=x=0}=0$ and $f(\lambda)|_{Q=x=0}=1$. We may assume that the basis $\{\phi_{i}\}_{i=0}^{s}$ is chosen so that $\phi_{0}=1$ and write $f(\lambda)=\sum_{i=0}^{s}f^{i}(\lambda)\phi_{i}$. We have
$I_{V}^{\lambda}=\sum_{i=0}^{s}M_{V}(\tau(\lambda))f^{i}(\lambda)\phi_{i}=f^{0}(\lambda)J_{V}^{\lambda}(\tau(\lambda))+\sum_{i=1}^{s}f^{i}(\lambda)(z\partial_{\tau^{i}}J_{V}^{\lambda})(\tau(\lambda))$
with $f^{0}(\lambda)|_{Q=x=0}=1$, $f^{i}(\lambda)|_{Q=x=0}=0$ for $i\geq 1$. Writing $\tau(\lambda)=\tau(0)+\delta(\lambda)$, $\delta(\lambda)=\sum_{j=0}^{s}\delta^{j}(\lambda)\phi_{j}$, we have
$I_{V}^{\lambda}=f^{0}(\lambda)\left(e^{\sum_{j=0}^{s}\delta^{j}(\lambda)\partial_{\tau^{j}}}J^{\lambda}_{V}\right)(\tau(0))+\sum_{i=1}^{s}f^{i}(\lambda)\left(z\partial_{\tau^{i}}e^{\sum_{j=0}^{s}\delta^{j}(\lambda)\partial_{\tau^{j}}}J_{V}^{\lambda}\right)(\tau(0))$
If we write $I_{\mathbb{P}(V)}(\tau,t)$ for the $I$-function in Theorem 1.1, the $I$-function in this Theorem 3.3 can be written as
$f^{0}(z\partial_{t})\left(e^{\sum_{j=0}^{s}\delta^{j}(z\partial_{t})\partial_{\tau^{j}}}I_{\mathbb{P}(V)}\right)(\tau(0),t)+\sum_{i=1}^{s}f^{i}(z\partial_{t})\left(z\partial_{\tau^{i}}e^{\sum_{j=0}^{s}\delta^{j}(z\partial_{t})\partial_{\tau^{j}}}I_{\mathbb{P}(V)}\right)(\tau(0),t).$
We have that $g(\tau,t):=e^{\sum_{j=0}^{s}\delta^{j}(z\partial_{t})\partial_{\tau^{j}}}zI_{\mathbb{P}(V)}(\tau,t)$ is a $\mathbb{C}[\![Q,\tau,t,x]\!]$-valued point of the Givental cone $\mathcal{L}_{\mathbb{P}(V)}$ by Lemma 2.7. Writing $T$ for the tangent space of $\mathcal{L}_{\mathbb{P}(V)}$ at $g$, we have from the overruled property that
$h(\tau,t)=(f^{0}(z\partial_{t})-1)g(\tau,t)+\sum_{i=1}^{s}f^{i}(z\partial_{t})z\partial_{\tau^{i}}g(\tau,t)$
gives an element of $zT$ that vanishes at $Q=x=0$; hence $g(\tau,t)+h(\tau,t)$ lies in the cone $\mathcal{L}_{\mathbb{P}(V)}$. By specializing $\tau$ to $\tau(0)$ in $g(\tau,t)+h(\tau,t)$, we obtain the $I$-function multiplied by $z$ as a $\mathbb{C}[\![Q,t,x]\!]$-valued point of $\mathcal{L}_{\mathbb{P}(V)}$. ∎
Example 3.4.
When $V$ is split, the $I$-function $I_{\mathbb{P}(V)}$ coincides with the Elezi-Brown $I$-function. Suppose that $V=L_{1}\oplus\cdots\oplus L_{r}$ with $L_{i}$ a line bundle over $B$. By the assumption that $V^{\vee}$ is generated by global sections, each $L_{i}^{-1}$ is nef. Let
$J_{B}(\tau)=\sum_{d\in\operatorname{Eff}(B)}J_{d}(\tau)Q^{d}$
be the $J$-function of $B$. The hypergeometric modification
$I_{V}^{\lambda}=\sum_{d\in\operatorname{Eff}(B)}\left(\prod_{i=1}^{r}\prod_{c=0}^{-c_{1}(L_{i})\cdot d-1}(\lambda+c_{1}(L_{i})-cz)\right)J_{d}(\tau)Q^{d}$
then lies in $\mathcal{L}_{V}$ by the quantum Lefschetz theorem [7]. Applying Theorem 3.3 to this $I_{V}^{\lambda}$, we recover the Elezi-Brown $I$-function
[…]
Example 3.5.
Let $B=\mathbb{P}^{n}$ with $n\geq 2$ and consider the cotangent bundle $T^{\vee}_{\mathbb{P}^{n}}$. We compute the $I$-function of $\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}})$ by using Theorem 3.3. The Euler sequence
$0\to T^{\vee}_{\mathbb{P}^{n}}(1)\to\mathcal{O}^{\oplus(n+1)}\to\mathcal{O}(1)\to 0,$
implies that $T^{\vee}_{\mathbb{P}^{n}}(1)$ satisfies the assumption in Theorem 1.1. It also implies that $\Delta_{T^{\vee}_{\mathbb{P}^{n}}(1)}^{\lambda}\Delta_{\mathcal{O}(1)}^{\lambda}=\Delta_{\mathcal{O}^{\oplus(n+1)}}^{\lambda}$. Since $\Delta_{\mathcal{O}^{\oplus(n+1)}}^{\lambda}$ preserves the Givental cone $\mathcal{L}_{\mathbb{P}^{n}}$, we have
$\mathcal{L}_{\mathbb{P}^{n},(T^{\vee}_{\mathbb{P}^{n}}(1),e_{\lambda}^{-1})}=\mathcal{L}_{\mathbb{P}^{n},(\mathcal{O}(1),e_{\lambda})}.$
Therefore, by the quantum Lefschetz theorem [7] for the $(\mathcal{O}(1),e_{\lambda})$-twist,
$I_{T^{\vee}_{\mathbb{P}^{n}}(1)}^{\lambda}=\sum_{d=0}^{\infty}e^{h\tau/z}Q^{d}e^{d\tau}\frac{\prod_{c=1}^{d}(h+\lambda+cz)}{\prod_{c=1}^{d}(h+cz)^{n+1}}$
lies in the cone of $T^{\vee}_{\mathbb{P}^{n}}(1)$, where $h\in H^{2}(B,\mathbb{Z})$ is the ample generator. Applying Theorem 3.3 to $I_{T^{\vee}_{\mathbb{P}^{n}}(1)}^{\lambda}$, we obtain the $I$-function for $\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}})=\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}}(1))$.
$I_{\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}})}=\sum_{k=0}^{\infty}\sum_{d=0}^{\infty}e^{(pt+h\tau)/z}q^{k}e^{kt}Q^{d}e^{d\tau}\frac{\prod_{c=1}^{k+d}(p+h+cz)}{\prod_{c=1}^{k}(p+cz)^{n+1}\prod_{c=1}^{d}(h+cz)^{n+1}}.$
We can obtain this $I$-function in a different way. The space $\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}})$ is isomorphic to the 2-step flag variety $\operatorname{Fl}(1,n-1;n)$, which can be embedded to $\mathbb{P}^{n}\times\mathbb{P}^{n}$ as a degree $(1,1)$-hypersurface. The $I$-function $I_{\mathbb{P}(T^{\vee}_{\mathbb{P}^{n}})}$ can be also obtained as a hypergeometric modification of the $J$-function of $\mathbb{P}^{n}\times\mathbb{P}^{n}$.
4. Fourier duality between quanum $D$-modules
In this section, we restate the mirror theorem (Theorem 3.3) as a Fourier duality between the equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ of $V$ and the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{P}(V)$. More precisely, we construct an isomorphism between $\operatorname{QDM}_{S^{1}}(V)$ and $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ that intertwines the shift operator $\mathbb{S}(\tau)$ with the Novikov variable $q$ and the equivariant parameter $\lambda$ with the quantum connection $z(\hat{\tau}^{*}\nabla)_{q\partial_{q}}$. Here $\hat{\tau}\colon H^{*}_{S^{1}}(V)\to H^{*}(\mathbb{P}(V))$ is a ‘mirror map’ that equals the Kirwan map modulo higher order terms in $q,Q$. We continue to assume that the dual bundle $V^{\vee}$ is globally generated.
4.1. Quantum $D$-modules
We fix notation and definitions concering the quantum $D$-modules of $\mathbb{P}(V)$ and $V$.
4.1.1. Quantum $D$-module of $\mathbb{P}(V)$
Let $\hat{\tau}\in H^{*}(\mathbb{P}(V))$ denote the parameter of the quantum cohomology of $\mathbb{P}(V)$. Following Notation 3.1, we write $(q,Q)$ for the Novikov variables of $\mathbb{P}(V)$, which are dual to the decomposition $H^{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}p\oplus H^{2}(B,\mathbb{Z})$. The quantum $D$-module of $\mathbb{P}(V)$ is the module
$\operatorname{QDM}(\mathbb{P}(V))=H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$
equipped with the quantum connection (see §2.2)
$\nabla_{\hat{\tau}^{i}},\ \nabla_{q\partial_{q}},\ \nabla_{\xi Q\partial_{Q}},\nabla_{z\partial_{z}}\colon\operatorname{QDM}(\mathbb{P}(V))\to z^{-1}\operatorname{QDM}(\mathbb{P}(V))$
where $\xi\in H^{2}(B)\subset H^{2}(\mathbb{P}(V))$ and the flat $z$-sesquilinear pairing (2.8)
$P_{\mathbb{P}(V)}\colon\operatorname{QDM}(\mathbb{P}(V))\times\operatorname{QDM}(\mathbb{P}(V))\to\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$
induced by the Poincaré pairing of $\mathbb{P}(V)$. We note that $\nabla_{q\partial_{q}}=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}})$. We also note that we defined $\operatorname{QDM}(\mathbb{P}(V))$ over $\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$ instead of over $\mathbb{C}[z][\![q,Q,\hat{\tau}]\!]$: it is possible because of the homogeneity in Gromov-Witten theory and the fact that both $z$ and $q$ have positive degrees.
4.1.2. Quantum $D$-module of $V$
Let $\tau\in H^{*}_{S^{1}}(V)$ denote the parameter of the equivariant quantum cohomology of $V$. Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(B)$ as before; here we assume that $\phi_{0}=1$. Then $\{\phi_{i}\lambda^{k}\}_{0\leq i\leq s,k\geq 0}$ gives a $\mathbb{C}$-basis of $H^{*}_{S^{1}}(V)$. We write $\tau=\sum_{i=0}^{s}\sum_{k=0}^{\infty}\tau^{i,k}\phi_{i}\lambda^{k}$ with $\tau^{i,k}$ being $\mathbb{C}$-valued coordinates. The equivariant quantum $D$-module of $V$ is the module
$\operatorname{QDM}_{S^{1}}(V)=H^{*}_{S^{1}}(V)\otimes_{\mathbb{C}[\lambda]}\mathbb{C}[\lambda,z][\![Q,\boldsymbol{\tau}]\!]=H^{*}(B)\otimes\mathbb{C}[\lambda,z][\![Q,\boldsymbol{\tau}]\!]$
equipped with the quantum connection (see §2.3)
$\nabla_{\tau^{i,k}},\ \nabla_{\xi Q\partial_{Q}},\ \nabla_{z\partial_{z}}\colon\operatorname{QDM}_{S^{1}}(V)\to z^{-1}\operatorname{QDM}_{S^{1}}(V)$
where $\xi\in H^{2}(B)=H^{2}(V)$, the shift operator $\mathbb{S}(\tau)\colon\operatorname{QDM}_{S^{1}}(V)\to\operatorname{QDM}_{S^{1}}(V)$ (see §2.7) and the flat $z$-sesquilinear pairing (2.8)
$P_{V}\colon\operatorname{QDM}_{S^{1}}(V)\times\operatorname{QDM}_{S^{1}}(V)\to\mathbb{C}(\lambda,z)[\![Q,\boldsymbol{\tau}]\!]$
induced by the Poincaré pairing (2.11) of $V$.
4.2. Fourier transformation for quantum $D$-modules
Recall the Fourier transformation (1.2) on the level of the Givental spaces from the introduction.
$\widehat{\phantom{A}}\colon$ $\mathcal{H}_{V,\rm pol}\to\mathcal{H}_{\mathbb{P}(V),\rm pol}$
$J(\lambda)\longmapsto\widehat{J}(q)=\sum_{k\in\mathbb{Z}}\kappa(\mathcal{S}^{-k}J)q^{k}=\sum_{k=0}^{\infty}\frac{J(p+kz)}{\prod_{c=1}^{k}e_{p+cz}(V)}q^{k},$
where $\mathcal{H}_{V,\rm pol}$, $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ are the polynomial Givental spaces given in (2.17), (2.15). (Note that $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ is defined over the Novikov ring $\mathbb{C}[\![q,Q]\!]$.)
Example 4.1 (cf. (1.4)).
Let $\tau_{\rm ne}=\sum_{i=0}^{s}\tau^{i,0}\phi_{i}\in H^{*}(V)$ denote the non-equivariant class. The Fourier transform of the $J$-function $J_{V}(\tau_{\rm ne}+\tau^{0,1}\lambda\phi_{0})=e^{\lambda\tau^{0,1}/z}J_{V}(\tau_{\rm ne})$ gives the $I$-function in Theorem 1.1 with $t=\tau^{0,1}$.
$J_{V}(\tau_{\rm ne}+\tau^{0,1}\lambda\phi_{0})^{^}=I_{\mathbb{P}(V)}(\tau_{\rm ne},\tau^{0,1}).$
More generally, we obtain $J_{V}(\textstyle\sum_{i,k}\tau^{i,k}\phi_{i}\lambda^{k})^{^}=e^{\sum_{k\geq 1,(i,k)\neq(0,1)}\tau^{i,k}(zq\partial_{q}+p)^{k}\partial_{\tau^{i}}}I_{\mathbb{P}(V)}(\tau_{\rm ne},\tau^{0,1})$ by using (1.3).
We show that the Fourier transformation induces an isomorphism between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$.
Theorem 4.2.
Let $V\to B$ be as in Theorem 1.1. There exist a mirror map $\hat{\tau}=\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[q][\![Q,\boldsymbol{\tau}]\!]$ and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules which satisfy the following properties:
  • (1)
    $\operatorname{FT}$ intertwines $\mathbb{S}(\tau)$ with $q$;
  • (2)
    $\operatorname{FT}$ intertwines $\lambda$ with $z\nabla^{\prime}_{q\partial_{q}}$;
  • (3)
    $\operatorname{FT}$ intertwines $\nabla_{\tau^{i,k}}$ with $\nabla^{\prime}_{\tau^{i,k}}$;
  • (4)
    $\operatorname{FT}$ intertwines $\nabla_{\xi Q\partial_{Q}}$ with $\nabla^{\prime}_{\xi Q\partial_{Q}}$ for $\xi\in H^{2}(B)$;
  • (5)
    $\operatorname{FT}$ intertwines $\nabla_{z\partial_{z}}+\frac{1}{2}$ with $\nabla^{\prime}_{z\partial_{z}}$;
  • (6)
    $\hat{\tau}(\tau)$ is homogeneous of degree $2$ and $\operatorname{FT}$ is homogeneous of degree 00 with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(\mathbb{P}(V))$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2c_{1}(TV)\cdot d=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
  • (7)
    $\hat{\tau}(\tau)|_{q=Q=0}=\kappa(\tau)$ and $\hat{\tau}(\tau)|_{Q=0,\tau^{*,r}=\tau^{*,r+1}=\cdots=0}\equiv\kappa(\tau)$ modulo $(\tau^{*,\leq r-1})^{2}$, where $(\tau^{*,\leq r-1})$ denotes the ideal generated by $\tau^{i,k}$ with $0\leq i\leq s$, $k\leq r-1$;
  • (8)
    $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $10\leq k\leq r-$ 1;
  • (9)
    $(M_{V}(\tau)s)^{^}=M_{\mathbb{P}(V)}(\hat{\tau}(\tau))\operatorname{FT}(s)$, i.e. the following diagram commutes:
    ${\operatorname{QDM}_{S^{1}}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{FT}\quad}$$\scriptstyle{M_{V}(\tau)}$${\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{\mathbb{P}(V)}(\hat{\tau}(\tau))}$${\mathcal{H}_{V,\rm pol}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\phantom{aa}}}$${\mathcal{H}_{\mathbb{P}(V),\rm pol},}$
where we set $\nabla^{\prime}=\hat{\tau}^{*}\nabla$.
Remark 4.3.
The pull-back $\nabla^{\prime}=\hat{\tau}^{*}\nabla$ is defined in the standard way as follows:
$\nabla^{\prime}_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\partial_{\tau^{i,k}}\hat{\tau})\star_{\hat{\tau}(\tau)},$
$\nabla^{\prime}_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}(\tau)}+(q\partial_{q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
$\nabla^{\prime}_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\hat{\tau}(\tau)}+(\xi Q\partial_{Q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
$\nabla^{\prime}_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{\mathbb{P}(V)}\star_{\hat{\tau}(\tau)})+\mu_{\mathbb{P}(V)}.$
Proof of Theorem 4.2.
Consider the $J$-function $J_{V}(\tau)$ with $\tau=\sum_{i,k}\tau^{i,k}\phi_{i}\lambda^{k}$. By Theorem 3.3, the Fourier transform $z\widehat{J}_{V}(\tau)$ lies in the Givental cone for $\mathbb{P}(V)$. Hence there exist unique $\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[\![q,Q,\boldsymbol{\tau}]\!]$ and $\Upsilon(\tau)\in H^{*}(\mathbb{P}(V))[z][\![q,Q,\boldsymbol{\tau}]\!]$ such that
$(M_{V}(\tau)1)^{^}=\widehat{J}_{V}(\tau)=M_{\mathbb{P}(V)}(\hat{\tau}(\tau))\Upsilon(\tau).$
By differentiating this in $\tau^{i,k}$, we obtain (using (2.7))
(4.1) $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}=M_{\mathbb{P}(V)}(\hat{\tau}(\tau))z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau).$
This shows that $M_{\mathbb{P}(V)}(\hat{\tau}(\tau))$ arises from the Birkhoff factorization of the matrix with column vectors $\{(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}:0\leq i\leq s,0\leq k\leq r-1\}$ because $M_{\mathbb{P}(V)}=\operatorname{id}+O(z^{-1})$ and $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ does not contain negative powers of $z$ (see [7]). Note that
(4.2) $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,\big{|}_{Q=q=0}=e^{\kappa(\tau)/z}\phi_{i}p^{k}$
and these vectors with $0\leq i\leq s$, $10\leq k\leq r-$ 1 form a basis of $H^{*}(\mathbb{P}(V))$. The mirror map $\hat{\tau}(\tau)$ is then determined by $M_{\mathbb{P}(V)}(\hat{\tau}(\tau))1=1+z^{-1}\hat{\tau}(\tau)+O(z^{-2})$. The formula (4.2) with $i=k=0$ implies that $\hat{\tau}(\tau)|_{Q=q=0}=\kappa(\tau)$ and $\Upsilon(\tau)|_{Q=q=0}=1$. Similarly, a direct calculation shows that $(M_{V}(\tau)1)^{^}\,|_{Q=0,\tau^{*,r}=\tau^{*,r+1}=\cdots=0}\equiv 1+z^{-1}\kappa(\tau)+O(z^{-2})$ modulo $(\tau^{*,\leq r-1})^{2}$. Part (7) follows from these computations.
$M_{V}(\tau)\phi_{i}\lambda^{k}$ is homogeneous of degree $\deg(\phi_{i}\lambda^{k})$. It is easy to see that its Fourier transform $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,$ is also homogeneous of the same degree. This implies the homogeneity of the Birkhoff factors $M_{\mathbb{P}}(V)(\hat{\tau}(\tau))$, $\Upsilon(\tau)$, $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ and hence of the mirror map $\hat{\tau}(\tau)$. The homogeneity implies that $\hat{\tau}(\tau)$ and $\Upsilon(\tau)$ depend polynomially on $q$ (for each fixed power of $Q$ and $\tau$), i.e. $\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!]$, $\Upsilon(\tau)\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$.
We define a $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-linear map $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ by sending $\phi_{i}\lambda^{k}$ to $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$. Then $\operatorname{FT}$ is homogeneous of degree zero and satisfies Part (9) by (4.1). Parts (1)-(5) follow from Part (9) together with the following relations
$(\mathcal{S}f)^{^}$ $=q\hat{f}$
$(\lambda f)^{^}$ $=(zq\partial_{q}+p)\hat{f}$
$(\partial_{\tau^{i,k}}f)^{^}$ $=\partial_{\tau^{i,k}}\hat{f}$
$((\xi Q\partial_{Q}+z^{-1}\xi)f)^{^}$ $=(\xi Q\partial_{Q}+z^{-1}\xi)\hat{f}$
$((z\partial_{z}-z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V}+\tfrac{1}{2})f)^{^}$ $=(z\partial_{z}-z^{-1}c_{1}(T\mathbb{P}(V))+\mu_{\mathbb{P}(V)})\hat{f}$
and the intertwining properties (2.7) of $M_{V}$ and $M_{\mathbb{P}(V)}$.
Finally we show that $\operatorname{FT}$ is an isomorphism (along with part (8)). It suffices to show that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}\in H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q]$ with $0\leq i\leq s$, $k\geq 0$ form a basis over $\mathbb{C}[z]$. We have
$(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,\big{|}_{Q=\tau=0}=\sum_{l=0}^{\infty}\frac{\phi_{i}(p+lz)^{k}}{\prod_{c=1}^{l}e_{p+cz}(V)}q^{l}=\begin{cases}\phi_{i}p^{k}+O(z^{-1})&k\leq r-1\\ \phi_{i}p^{r}+q\phi_{i}+O(z^{-1})&k=r\end{cases}$
which shows that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $10\leq k\leq r-$ 1 and, since the multiplication by $\lambda$ corresponds to $z\nabla^{\prime}_{q\partial_{q}}$ under $\operatorname{FT}$, that
$z\nabla^{\prime}_{q\partial_{q}}(\phi_{i}p^{k})\big{|}_{Q=\tau=0}=\begin{cases}\phi_{i}p^{k+1}&0\leq k\leq r-2;\\ \phi_{i}p^{r}+q\phi_{i}&k=r-1.\end{cases}$
Hence for $k=rn+m$ with $10\leq m\leq r-$ 1, we have
$\operatorname{FT}(\phi_{i}\lambda^{k})\big{|}_{Q=\tau=0}$ $=(z\nabla^{\prime}_{q\partial_{q}})^{k}\phi_{i}\big{|}_{Q=\tau=0}$
$=\phi_{i}p^{m}q^{n}+(\text{polynomial in $q$ of degree $\leq(n-1)$})+O(z).$
The homogeneity of $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}$ implies that they form a $\mathbb{C}[z]$-basis of $H^{*}(\mathbb{P}(V))\otimes\mathbb{C}[z,q]$. ∎
We can also identify the pairing $P_{\mathbb{P}(V)}$ with a Fourier transform of $P_{V}$. Define the pairing $\widehat{P}_{V}$ on $\operatorname{QDM}_{S^{1}}(V)$ as follows (see Remark 4.6 for another definition):
(4.3) $\widehat{P}_{V}(x,y):=\sum_{k,l\in\mathbb{Z}}q^{k+l}\operatorname{Res}_{\lambda=0}P_{V}(\mathbb{S}(\tau)^{-k}x,\mathbb{S}(\tau)^{-l}y)d\lambda.$
It is clear from the definition that $\widehat{P}_{V}(\mathbb{S}(\tau)x,y)=\widehat{P}_{V}(x,\mathbb{S}(\tau)y)=q\widehat{P}_{V}(x,y)$.
Proposition 4.4.
The pairing $\widehat{P}_{V}$ is well-defined and takes values in $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$. We have
$\widehat{P}_{V}(x,y)=P_{\mathbb{P}(V)}(\operatorname{FT}(x),\operatorname{FT}(y))$
for $x,y\in\operatorname{QDM}_{S^{1}}(V)$.
Proof.
For $x,y\in\operatorname{QDM}_{S^{1}}(V)$, we set $f=M_{V}(\tau)x$ and $g=M_{V}(\tau)y$. We have
$\widehat{P}_{V}(x,y)$ $=\sum_{k,l\in\mathbb{Z}}q^{k+l}\operatorname{Res}_{\lambda=0}P_{V}(\mathcal{S}^{-k}f,\mathcal{S}^{-l}g)d\lambda$ by (2.9) and (2.21)
$=\sum_{k,l\in\mathbb{Z}}q^{k+l}P_{\mathbb{P}(V)}(\kappa(\mathcal{S}^{-k}f),\kappa(\mathcal{S}^{-l}g))$ by Lemma 4.5 below
$=P_{\mathbb{P}(V)}(\hat{f},\hat{g})$
$=P_{\mathbb{P}(V)}(M_{\mathbb{P}(V)}(\hat{\tau})\operatorname{FT}(x),M_{\mathbb{P}(V)}(\hat{\tau})\operatorname{FT}(y))$ by Theorem 4.2(9)
[…]
Note that the summands are non-zero only when $k,l\geq 0$, and hence the sum is well-defined. It is clear that $P_{\mathbb{P}(V)}(\operatorname{FT}(x),\operatorname{FT}(y))$ lies in $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$. ∎
Lemma 4.5.
Let $\pi\colon\mathbb{P}(V)\to B$ denote the projection. For $f=f(\lambda)\in H^{*}(B)[\![\lambda]\!]$, we have
$\pi_{*}(\kappa(f))=\operatorname{Res}_{\lambda=0}\frac{f(\lambda)}{e_{\lambda}(V)}d\lambda.$
In particular, by integrating it over $B$, we have
$\int_{\mathbb{P}(V)}\kappa(f)=\operatorname{Res}_{\lambda=0}\left(\int^{S^{1}}_{V}f\right)d\lambda.$
Proof.
The map $f\mapsto\pi_{*}(\kappa(f))$ is uniquely characterized by the following properties:
  • (1)
    it is a homomorphism of $H^{*}(B)$-modules;
  • (2)
    it vanishes on $e_{\lambda}(V)H^{*}(B)[\![\lambda]\!]$;
  • (3)
    $\pi_{*}(\kappa(\lambda^{i}))=0$ for $20\leq i\leq r-$ 2 and $\pi_{*}(\kappa(\lambda^{r-1}))=1$.
It is easy to check that the right-hand side satisfies the same properties. ∎
Remark 4.6.
For $x,y\in\operatorname{QDM}_{S^{1}}(V)$, we have
$P_{V}(\mathbb{S}(\tau)^{-k}x,\mathbb{S}(\tau)^{-l}y)=e^{-kz\partial_{\lambda}}P_{V}(x,\mathbb{S}(\tau)^{-k-l}y)=e^{-kz\partial_{\lambda}}P_{V}(f,\mathcal{S}^{-k-l}g)$
with $f=M_{V}(\tau)x$, $g=M_{V}(\tau)y$ (where we used (2.22)) and $P_{V}(f,\mathcal{S}^{-d}g)=P_{V}(f,e_{\lambda+z}(V)^{-1}\cdots e_{\lambda+dz}(V)^{-1}e^{dz\partial_{\lambda}}g)$ has poles only at $\lambda=-jz$ with $j\in[0,d]\cap\mathbb{Z}$ as a function of $\lambda$. Therefore the pairing $\widehat{P}_{V}$ can be rewritten as
$\widehat{P}_{V}(x,y)$ $=\sum_{d\in\mathbb{Z}}q^{d}\sum_{k+l=d}\operatorname{Res}_{\lambda=0}e^{-kz\partial_{\lambda}}P_{V}(f,\mathcal{S}^{-d}g)d\lambda$
$=\sum_{d\in\mathbb{Z}}q^{d}\sum_{0\leq k\leq d}\operatorname{Res}_{\lambda=-kz}P_{V}(f,\mathcal{S}^{-d}g)d\lambda$
$=-\sum_{d\in\mathbb{Z}}q^{d}\operatorname{Res}_{\lambda=\infty}P_{V}(f,\mathcal{S}^{-d}g)d\lambda$
$=-\sum_{d\in\mathbb{Z}}q^{d}\operatorname{Res}_{\lambda=\infty}P_{V}(x,\mathbb{S}(\tau)^{-d}y)d\lambda.$
Example 4.7.
Consider the vector bundle $V=\mathcal{O}(-1)\oplus\mathcal{O}^{\oplus(r-1)}$ over $\mathbb{P}^{1}$. We set the parameter $\tau$ to be zero and denote by $Q$ the Novikov variable corresponding to the positive generator of $H_{2}(\mathbb{P}^{1},\mathbb{Z})$. In this example, we can consider the equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ over the polynomial ring $\mathbb{C}[z,\lambda,Q]$; $\operatorname{QDM}_{S^{1}}(V)$ is generated by $1$ as a $\mathbb{C}[z]\langle Q,z\nabla_{Q\partial_{Q}},\lambda,\mathbb{S}\rangle$-module and is defined by the relations
$((z\nabla_{Q\partial_{Q}})^{2}+Q(z\nabla_{Q\partial_{Q}}-\lambda))\cdot 1$ $=0,$
$(\mathbb{S}-\lambda^{r-1}(\lambda-z\nabla_{Q\partial_{Q}}))\cdot 1$ $=0.$
The hyperplane class $h\in H^{2}(\mathbb{P}^{1})$ is given by $h=z\nabla_{Q\partial_{Q}}1$. The quantum connection and the shift operator can be written in the $\mathbb{C}[z,\lambda,Q]$-basis $\{1,h\}$ as
$\nabla=d+\frac{1}{z}\begin{pmatrix}0&\lambda Q\\ 1&-Q\end{pmatrix}\frac{dQ}{Q},\qquad\mathbb{S}=\begin{pmatrix}\lambda^{r}&-\lambda^{r}Q\\ -\lambda^{r-1}&\lambda^{r}+Q\lambda^{r-1}\end{pmatrix}\circ e^{-z\partial_{\lambda}}$
where we omitted the connection in the $z$-direction. Note that the quantum connection does not depend on the rank $r$ but the shift operator does. We identify $\mathbb{S}$ with the Novikov variable $q$ for $\mathbb{P}(V)$ associated with the class of a line in the fibre. We can check that $\{1,\lambda,\dots,\lambda^{r-1},h,h\lambda,\dots,h\lambda^{r-1}\}$ gives a basis of $\operatorname{QDM}_{S^{1}}(V)$ over $\mathbb{C}[z,q,Q]=\mathbb{C}[z,\mathbb{S},Q]$. The action of $\lambda$ defines the connection $z\nabla_{q\partial_{q}}$ in the $q$-direction; the Fourier-transformed connection on the $(q,Q)$-space can be written in the above basis as
$zd+\left(\begin{array}[]{ccccc|ccccc}0&0&&&q&0&0&&&qQ\\ 1&0&&&0&0&0&&&0\\ 0&1&&&&&&&&\\ &&\ddots&&&&&&&\\ 0&&&1&0&0&0&&&0\\ \hline\cr 0&0&&&0&0&0&&&q\\ &&&&&1&0&&&0\\ &&&&&0&1&&&\\ 0&0&&&0&&&\ddots&&\\ 0&0&&&1&0&&&1&0\end{array}\right)\frac{dq}{q}+\left(\begin{array}[]{ccccc|ccccc}0&0&&&0&0&0&&&qQ\\ 0&0&&&&Q&0&&&0\\ &&&&&0&Q&&&\\ &&&&&&&\ddots&&\\ 0&&&&0&&&&Q&0\\ \hline\cr 1&0&&&0&-Q&0&&&0\\ 0&1&&&&0&-Q&&&\\ &&\ddots&&&&&\ddots&&\\ &&&&&&&&-Q&0\\ 0&&&&1&0&&&0&0\end{array}\right)\frac{dQ}{Q}.$
The matrices appearing here represent the multiplication by $p$ and $h$ in the quantum cohomology of $\mathbb{P}(\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}^{\oplus(r-1)})$ with respect to the basis $\{1,p,\dots,p^{r-1},h,hp,\dots,hp^{r-1}\}$.
5. Decomposition of the quantum $D$-module of $\mathbb{P}(V)$
In this section we construct a decomposition (Theorem 1.7) of the quantum $D$-module of $\mathbb{P}(V)$ into a direct sum of copies of the quantum $D$-module of $B$.
5.1. Statement
We state a more precise version of Theorem 1.7. Let $\operatorname{QDM}(\mathbb{P}(V))$ be the quantum $D$-module of $\mathbb{P}(V)$ as in §4.1.1. Let $\{\phi_{i}\}_{i=0}^{s}$ be a basis of $H^{*}(B)$ with $\phi_{0}=1$ as before. Then $\{\phi_{i}p^{k}\}_{0\leq i\leq s,0\leq k\leq r-1}$ form a basis of $H^{*}(\mathbb{P}(V))$; we write $\hat{\tau}=\sum_{i=0}^{s}\sum_{k=0}^{r-1}\hat{\tau}^{i,k}\phi_{i}p^{k}$ for the parameter in $H^{*}(\mathbb{P}(V))$. Set
(5.1) $r^{\prime}:=\begin{cases}r&\text{when $r-1$ is even;}\\ 2r&\text{when $r-1$ is odd}.\end{cases}$
We introduce the localized quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ as
$\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}=\operatorname{QDM}(\mathbb{P}(V))\otimes_{\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!].$
It is equipped with the connection $\nabla$ and the pairing $P_{\mathbb{P}(V)}$ as before.
Theorem 5.1.
There exist maps $H^{*}(\mathbb{P}(V))\to H^{*}(B)$, $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ with $\varsigma_{j}(\hat{\tau})\in-c_{1}(V)\log\lambda_{j}+H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$, $j=0,\dots,r-1$ with $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$ and an isomorphism
$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
such that
  • (1)
    $\Phi$ intertwines $\nabla$ with $\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\nabla$;
  • (2)
    $\Phi$ intertwines $P_{\mathbb{P}(V)}$ with $\bigoplus_{j=0}^{r-1}P_{B}$;
  • (3)
    $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ is homogeneous of degree $2$ and $\Phi$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}(\mathbb{P}(V))$, $H^{*}(B)$ and the following degree of the variables: $\deg\hat{\tau}^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
  • (4)
    $\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}-c_{1}(V)\log\lambda_{j}+O(q^{-1/r})$;
  • (5)
    the Jacobian matrix of $\varsigma(\hat{\tau})=\bigoplus_{j=0}^{r-1}\varsigma_{j}(\hat{\tau})$ along $Q=\hat{\tau}=0$ is $(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})|_{Q=\hat{\tau}=0}=\phi_{i}\lambda_{j}^{k}(1+O(q^{-1/r}))$ and is invertible over $\mathbb{C}(\!(q^{-1/r})\!)$;
  • (6)
    $\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}(\phi_{i}+O(q^{-1/r}))$
where $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ is the module
$\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}=H^{*}(B)\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$
equipped with the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection of $B$ and the pairing $P_{B}$ (2.8).
Remark 5.2.
Although the maps $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ contain $\log q$, the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection is well-defined by the Divisor Equation. We can eliminate $\log q$ from the statement of the above theorem by considering a coordinate change for the Novikov variables. Consider the map
$\varsigma^{\rm reg}\colon(\hat{\tau},q,Q)\longmapsto(\varsigma,Q_{B})=\left(\{\varsigma_{j}(\hat{\tau})+c_{1}(V)(\log q)/r\}_{j=0}^{r-1},q^{-c_{1}(V)/r}Q\right)$
where $\varsigma$ is a parameter of $\bigoplus_{j=0}^{r-1}H^{*}(B)$ and $Q_{B}$ is the Novikov variable for $B$. The map $\varsigma^{\rm reg}$ does not contain $\log q$. Then the Divisor Equation shows that $\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\nabla$ is the same as $(\varsigma^{\rm reg})^{*}(\nabla^{\oplus r})$. The identification $Q_{B}^{d}=q^{-c_{1}(V)\cdot d/r}Q^{d}$ of the Novikov variables is compatible with their degrees.
Remark 5.3.
As suggested in [28, 24], we expect that the (formal) decomposition of quantum cohomology $D$-modules occurs over formal power series in $z$. We defined the localized quantum $D$-modules over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$, but more precisely we can define them over the ‘homogeneous’ completion $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ consisting of finite sums of homogeneous elements in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$. We observe that $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ is contained in the ring $\mathbb{C}[q^{-1/r^{\prime}},q^{1/r^{\prime}}][\![z]\!]$ of formal power series in $z$. If we have convergence of quantum cohomology with respect to the parameters $q$, $Q$, $\hat{\tau}$, we expect that the decomposition $\Phi$ cannot be analytified in $z$ but should be defined over $\mathcal{O}[\![z]\!]$, with $\mathcal{O}$ the sheaf of analytic functions in $q,Q,\hat{\tau}$. As the picture in [28, 24] suggests, this formal (in $z$) decomposition should be lifted to an analytic decomposition over an angular sector on the $z$-plane, and we expect that the analytic decomposition can be related (at certain values of the parameters) to the semi-orthogonal decomposition [43] of derived categories of coherent sheaves
$D^{b}(\mathbb{P}(V))=\langle D^{b}(B)_{0},\dots,D^{b}(B)_{r-1}\rangle$
via the $\widehat{\Gamma}$-integral structure, where $D^{b}(B)_{i}=D^{b}(B)\otimes\mathcal{O}(i)$.
5.2. Idea of the proof
We first observe that the modified QRR operator $\Delta_{V}^{\lambda}$ (3.3) gives rise to a solution of the difference-differential module $\operatorname{QDM}_{S^{1}}(V)$. The equation (3.4) implies that $\Delta_{V}^{\lambda}$ solves the following difference equation:
(5.2) $e^{-z\partial_{\lambda}}\circ(\Delta_{V}^{\lambda})^{-1}=(\Delta_{V}^{\lambda})^{-1}\circ\mathcal{S}.$
The operator $\Delta_{V}^{\lambda}$ clearly commutes with $\partial_{\tau^{i,k}}$ and $\xi Q\partial_{Q}+z^{-1}\xi$ (with $\xi\in H^{2}(B)$). Therefore the map assigning to a section $s\in\operatorname{QDM}_{S^{1}}(V)$ the $H^{*}(B)$-valued function $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)s$ intertwines the actions of $\mathbb{S}(\tau)$, $\nabla_{\tau^{i,k}}$, $\nabla_{\xi Q\partial_{Q}}$ with the operators $e^{-z\partial_{\lambda}}$, $\partial_{\tau^{i,k}}$, $\xi Q\partial_{Q}+z^{-1}\xi$ respectively (thanks to the intertwining properties (2.7) of $M_{V}(\tau)$) and can be regarded as a solution for $\operatorname{QDM}_{S^{1}}(V)$. Therefore its Fourier transform
[…]
should give a solution for $\operatorname{QDM}(\mathbb{P}(V))$ through the Fourier isomorphism $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\cong\operatorname{QDM}_{S^{1}}(V)$ in Theorem 4.2. Despite the convergence issue of $\Delta_{V}^{\lambda}$ raised in Remark 3.2, we can formally make sense of this Fourier transformation using the stationary phase approximation (i.e. the method of steepest descent). The QRR theorem tells us that this lies in the Givental cone of $B$ and induces projections $\operatorname{QDM}(\mathbb{P}(V))\to\varsigma_{j}^{*}\operatorname{QDM}(B)$, $j=0,\dots,r-1$, where the $r$ choices result from the choice of critical points we use for the stationary phase method.
Remark 5.4.
The QRR operator $\Delta_{V}^{\lambda}$ induces an isomorphism between the quantum $D$-modules of $V$ and $B$ after localizing $\lambda$. The isomorphism is given by a gauge transformation $R(\tau)\in\lambda^{-r/2}\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[z](\!(\lambda^{-1})\!)[\![Q,\tau]\!]$ from $\operatorname{QDM}_{S^{1}}(V)$ to $\sigma^{*}\operatorname{QDM}(B)$ satisfying
$M_{B}(\sigma(\tau))R(\tau)=(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$
and a coordinate change $\tau\mapsto\sigma=\sigma(\tau)\in r(\lambda-\lambda\log\lambda)-c_{1}(V)\log\lambda+H^{*}(B)[\lambda,\lambda^{-1}][\![Q,\tau]\!]$ of the form $\sigma(\tau)=\sigma_{\rm pert}+\tau+O(Q)$ with
$\sigma_{\rm pert}=\sum_{\delta:\text{Chern roots of $V$}}\left[(\lambda+\delta)-(\lambda+\delta)\log(\lambda+\delta)\right].$
The shift operator acts on $\operatorname{QDM}(B)$ by the trivial one $e^{-z\partial_{\lambda}}$, but it acts highly non-trivially on the pull-back by the $\lambda$-dependent coordinate change $\tau\mapsto\sigma(\tau)$. From this point of view, the solution $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$ considered above corresponds to the standard solution $M_{B}(\sigma(\tau))$ for $\sigma^{*}\operatorname{QDM}(B)$.
5.3. Fourier transformation via stationary phase method
In this section we define the stationary phase asymptotics of the Fourier integral
[…]
for $J(\lambda)\in H^{*}_{S^{1}}(V)[z,z^{-1}]=H^{*}(B)[\lambda,z,z^{-1}]$. The integral can be written as
[…]
where $\varphi(\lambda)=r(\lambda\log\lambda-\lambda)-\lambda\log q$ and $K(\lambda)=(\widetilde{\Delta}_{V}^{\lambda})^{-1}J(\lambda)$ (see §3.3 for $\widetilde{\Delta}_{V}^{\lambda}$). We note that $K(\lambda)$ lies in $H^{*}(B)[z,z^{-1}](\!(\lambda^{-1})\!)$. We expand the integrand at the critical point $\lambda_{0}:=q^{1/r}$ of the phase function $\varphi(\lambda)$. We have
$\varphi(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})=-r\lambda_{0}+\frac{r}{2}s^{2}+\varphi_{\geq 3}(s,\lambda_{0})\quad\text{with $\varphi_{\geq 3}(s,\lambda_{0})=r\sum_{k=3}^{\infty}\frac{k-1}{k!}\frac{s^{k}}{\sqrt{\lambda_{0}}^{k-2}}$}$
and hence the Fourier integral can be rewritten in the form
[…]
with $L(s,\lambda_{0})\in H^{*}(B)[z,z^{-1},s](\!(\lambda_{0}^{-1/2})\!)$ given by
$L(s,\lambda_{0})$ $=e^{-(\frac{r}{2}-1)u}e^{-uc_{1}(V)/z}e^{-\varphi_{\geq 3}(s,\lambda_{0})/z}K(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})\qquad\text{with $u=s/\sqrt{\lambda_{0}}$.}$
We expand $L(s,\lambda_{0})$ in power series in $s$ and perform the Gaussian integral (5.3) termwise over $s\in\mathbb{R}$ (assuming $z>0$) and arrive at the formal stationary phase asymptotics
[…]
with $\mathscr{F}_{0}(J)\in\lambda_{0}^{-(r-1)/2}H^{*}(B)[z,z^{-1}](\!(\lambda_{0}^{-1})\!)[\log\lambda_{0}]$ given by
$\mathscr{F}_{0}(J)=\frac{1}{\sqrt{r}}\lambda_{0}^{-(r-1)/2}e^{-c_{1}(V)\log\lambda_{0}/z}\left[e^{z(\partial_{s})^{2}/(2r)}L(s,\lambda_{0})\right]_{s=0}.$
Note that half-integer powers of $\lambda_{0}$ in $L(s,\lambda_{0})$ do not contribute. The monodromy transformation on the $q$-plane with respect to the $(2\pi j)$ rotation maps $\lambda_{0}$ to $\lambda_{j}:=e^{2\pi\sqrt{-1}j/r}q^{1/r}$ and $\mathscr{F}_{0}(J)$ to $\mathscr{F}_{j}(J):=\mathscr{F}_{0}(J)|_{\lambda_{0}\to\lambda_{j}}$. We have defined the Fourier transformations $\mathscr{F}_{j}\colon\mathcal{H}_{V,\rm pol}\to\mathcal{H}_{B,\mathbb{C}(\!(q^{-1/r^{\prime}})\!)}$, $j=0,\dots,r-1$ taking values in
$\mathcal{H}_{B,\mathbb{C}(\!(q^{-1/r^{\prime}})\!)}:=H^{*}(B)\otimes\mathbb{C}[z,z^{-1}](\!(q^{-1/r^{\prime}})\!)[\![Q]\!][\log q]$
where $r^{\prime}$ is as in (5.1).
Proposition 5.5.
The map $\mathscr{F}_{j}\colon\mathcal{H}_{V,\rm pol}\to\mathcal{H}_{B,\mathbb{C}(\!(q^{-1/r^{\prime}})\!)}$ satisfies the following properties:
  • (1)
    $\mathscr{F}_{j}(\lambda J)=(\lambda_{j}+zq\partial_{q})\mathscr{F}_{j}(J)$;
  • (2)
    $\mathscr{F}_{j}(\mathcal{S}J)=q\mathscr{F}_{j}(J)$;
  • (3)
    $\mathscr{F}_{j}((z\partial_{z}-z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V}+\tfrac{1}{2})J)=(z\partial_{z}-z^{-1}(c_{1}(TB)+r\lambda_{j})+\mu_{B})\mathscr{F}_{j}(J)$.
Proof.
Parts (1), (2) follow by taking the stationary phase asymptotics of the standard properties of the Fourier transformation:
[…]
where we shift $\lambda$ as $\lambda\mapsto\lambda-z$ and used (5.2) in the last line. We note that the shift $\lambda\mapsto\lambda-z$ translates into the change of variables
$s\mapsto s+\sqrt{\lambda_{j}}\log\left(1-z\lambda_{j}^{-1}e^{-s/\sqrt{\lambda_{j}}}\right)$
in terms of the coordinate $s$ around the critical point (given by $\lambda=\lambda_{j}e^{s/\sqrt{\lambda_{j}}}$) and this makes sense in the ring $\mathbb{C}[z,z^{-1},s](\!(\lambda_{j}^{-1/2})\!)$. To show part (3), observe that the map $J\mapsto e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J)$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$ and $\deg q=2r$, $\deg z=2$. This means
$\mathscr{F}_{j}((z\partial_{z}+\mu_{V}+\tfrac{1}{2})J)=(z\partial_{z}+rq\partial_{q}+\mu_{B}+z^{-1}c_{1}(V))\mathscr{F}_{j}(J).$
Part (3) follows from this together with part (1), the fact that $\mathscr{F}_{j}$ is an $H^{*}(B)[z,z^{-1}]$-module homomorphism and $c_{1}^{S^{1}}(TV)=c_{1}(TB)+c_{1}(V)+r\lambda$. ∎
5.4. Fourier transform of the $J$-function
In this section we show that the Fourier transform $z\mathscr{F}_{j}(J_{V}(\tau))$ of the $J$-function of $V$ lies in the Givental cone of $B$. Roughly speaking, the argument goes as follows. The QRR theorem says that $z(\Delta_{V}^{\lambda})^{-1}J_{V}(\tau)$ “lies” in $\mathcal{L}_{B}$ modulo the convergence issue. The formal asymptotic expansion of the integral
[…]
is given by applying the differential operator $\sqrt{\frac{2\pi z}{\varphi^{\prime\prime}(\lambda_{j})}}e^{z(\partial_{\lambda})^{2}/(2\varphi^{\prime\prime}(\lambda_{j}))}$ to the integrand multiplied by $e^{\varphi^{\prime\prime}(\lambda_{j})(\lambda-\lambda_{j})^{2}/(2z)}$ and evaluating it at the critical point $\lambda=\lambda_{j}$. Hence it should lie in $\mathcal{L}_{B}$ by a property of the cone (see Lemma 2.7) up to powers of $z$. A more careful analysis shows the following:
Proposition 5.6.
Write $\tau(\lambda)=\sum_{i,k}\tau^{i,k}\lambda^{k}\phi_{i}$ for the parameter in $H^{*}_{S^{1}}(V)$. There exist
$\sigma_{j}$ $\in-c_{1}(V)\log\lambda_{j}+\tau(\lambda_{j})+\mathfrak{m}H^{*}(B)[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
$\upsilon_{j}$ $\in\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}(1+\mathfrak{m}^{\prime}H^{*}(B)[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!])$
with $\lambda_{j}=q^{1/r}e^{2\pi\sqrt{-1}j/r}$ such that $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$, where
  • $q^{\bullet/r}\boldsymbol{\tau}$ stands for the infinite set $\{q^{k/r}\tau^{i,k}\}$ of variables;
  • $K[\![q^{-c_{1}(V)/r}Q]\!]=\left\{\sum_{d\in\operatorname{Eff}(B)}c_{d}q^{-c_{1}(V)\cdot d/r}Q^{d}:c_{d}\in K\right\}$ for a module $K$; and
  • $\mathfrak{m}$ ((((resp. $\mathfrak{m}^{\prime}$)))) is the closed ideal of $\mathbb{C}[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ ((((resp. of $\mathbb{C}[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$)))) generated by $q^{-1/r}$ and $q^{-c_{1}(V)\cdot d/r}Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$.
Moreover, $\sigma_{j}+c_{1}(V)\log\lambda_{j}$ and $\upsilon_{j}$ are homogeneous of degree $2$ and $-(r-1)$ respectively with respect to the usual grading on $H^{*}(B)$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$ and $\deg z=2$.
Proof.
We start with the $J$-function $J_{V}(\tau)$ with parameter $\tau=\sum_{i}\tau^{i}\phi_{i}$. By Remark 2.6, $\widetilde{J}:=J_{V}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ is the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory. Thus $I:=z(\widetilde{\Delta}^{\lambda}_{V})^{-1}\widetilde{J}$ is a $\mathbb{C}[\![\lambda^{-1},Q,\tau]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5. By substituting $\sum_{k\geq 0}\tau^{i,k}\lambda^{k}$ for $\tau^{i}$, we regard it as a $\mathbb{C}[\![\lambda^{-1},Q,\lambda^{\bullet}\boldsymbol{\tau}]\!]$-valued point on $\mathcal{L}_{B}$, where $\lambda^{\bullet}\boldsymbol{\tau}$ denotes the set $\{\lambda^{k}\tau^{i,k}\}$ of variables. By the definition of $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))$ in the previous section, we have
$ze^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))\Bigr{|}_{Q\to\lambda_{j}^{c_{1}(V)}Q}\\ =\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\left.e^{z(\partial_{s})^{2}/(2r)}\left[e^{-(\frac{r}{2}-1)u}e^{-uc_{1}(V)/z}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}I\biggr{|}_{\scriptsize\!\!\!\begin{array}[]{l}Q\to Qe^{-uc_{1}(V)}\\ \lambda\to\lambda_{j}e^{u}\end{array}}\right]\right|_{s=0}$
where we set $u=s/\sqrt{\lambda_{j}}$ as before and $Q\to\lambda^{c_{1}(V)}Q$ means $Q^{d}\to\lambda^{c_{1}(V)\cdot d}Q^{d}$. The quantity in the big parenthesis []delimited-[][$\cdots$][ ⋯ ] belongs to $H^{*}(B)[z,z^{-1},s][\![\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and defines a $\mathbb{C}[\![s,\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$-valued point of $\mathcal{L}_{B}$: the Dilaton and String Equations show that $\mathcal{L}_{B}$ is invariant under $e^{-(\frac{r}{2}-1)u}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}$ and the Divisor Equation shows that $\mathcal{L}_{B}$ is invariant under $f(Q)\to e^{-uc_{1}(V)/z}f(Qe^{-uc_{1}(V)})$. By Lemma 2.7, the action of $e^{z(\partial_{s})^{2}/(2r)}$ preserves points on the cone [^7] Strictly speaking, the operator $(z\partial_{s})^{2}/(2r)$ does not vanish at the origin of the formal parameters as required in Lemma 2.7, but the conclusion holds since the action of $e^{z(\partial_{s})^{2}/(2r)}$ is well-defined.. Therefore we have $\tilde{\sigma}_{j}\in H^{*}(B)[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and $\tilde{\upsilon}_{j}\in H^{*}(B)[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ such that $\tilde{\sigma}_{j}\equiv\tau(\lambda_{j})$, $\tilde{\upsilon}_{j}\equiv 1$ modulo the closed ideal in $\mathbb{C}[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ or in $\mathbb{C}[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ generated by $\lambda_{j}^{-1}$ and $Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$ and
$e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))\Bigr{|}_{Q\to\lambda_{j}^{c_{1}(V)}Q}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}M_{B}(\tilde{\sigma}_{j})\tilde{\upsilon}_{j}.$
As in the proof of Theorem 4.2, $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ can be obtained from the Birkhoff factorization of the matrix formed by the derivatives of the left-hand side with respect to $\tau^{i,0}$. It follows from the homogeneity of $e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))$ that $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ are homogeneous of degree zero after the change $Q\to\lambda_{j}^{-c_{1}(V)}Q$ of variables. By the Divisor Equation again, we obtain $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$ with
$\sigma_{j}=-c_{1}(V)\log\lambda_{j}+\tilde{\sigma}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q},\quad\upsilon_{j}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\tilde{\upsilon}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q}.$
The conclusion follows. ∎
Remark 5.7.
The leading term $-c_{1}(V)\log\lambda_{j}$ of $\sigma_{j}$ accounts for difference of the degree of $Q^{d}$ between $B$ and $V$.
Example 5.8.
Computing the formal stationary phase asymptotics of (5.4) at $Q=0$ over the Artinian local ring $H^{*}(B)$, we see that $r\lambda_{j}+\sigma_{j}|_{Q=0}$ is given by the critical value of (cf. (3.7))
$f(\lambda)=\tau(\lambda)+\lambda\log q-\sum_{\delta:\text{Chern roots of $V$}}((\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta))$
at a unique $H^{*}(B)$-valued critical point $\hat{\lambda}_{j}\in\lambda_{j}(1+\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1},\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!])$, which is a solution $\lambda$ to the equation $q=e_{\lambda}(V)e^{-\tau^{\prime}(\lambda)}$. The critical point $\hat{\lambda}_{j}$ and the critical value $f(\hat{\lambda}_{j})$ are the eigenvalues of the quantum products $(p\star)|_{Q=0}$ and $(E_{\mathbb{P}(V)}\star)|_{Q=0}$ in the $H^{*}(B)$-algebra $QH^{*}(\mathbb{P}(V))|_{Q=0}$ respectively. (It is easy to see that $H^{*}(B)$ acts by cup product on $QH^{*}(\mathbb{P}(V))|_{Q=0}$.)
5.5. Projections induced by the Fourier transformations
Let $\sigma_{j}$ be as in Proposition 5.6. We regard it as a multi-valued map from $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ to $H^{*}(B)$, where $r^{\prime}$ is as in (5.1), and consider the pull-back $\sigma_{j}^{*}\operatorname{QDM}(B)$ of the quantum $D$-module of $B$ by $\sigma_{j}$. We define $\sigma_{j}^{*}\operatorname{QDM}(B)$ to be the module
$\sigma_{j}^{*}\operatorname{QDM}(B)=H^{*}(B)[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
equipped with the pull-back $\sigma_{j}^{*}\nabla$ of the quantum connection of $B$:
$(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\partial_{\tau^{i,k}}\sigma_{j})\star_{\sigma_{j}}$
$(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(q\partial_{q}\sigma_{j})\star_{\sigma_{j}}$
$(\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\sigma_{j}}+(\xi Q\partial_{Q}\sigma_{j})\star_{\sigma_{j}})\qquad(\xi\in H^{2}(B))$
$(\sigma_{j}^{*}\nabla)_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}E_{B}\star_{\sigma_{j}}+\mu_{B}$
and the pairing $P_{B}$ as in (2.8). The Divisor Equation show that these pull-backs are well-defined and define maps
$q^{-k/r}(\sigma_{j}^{*}\nabla)_{\tau^{i,k}},\ (\sigma_{j}^{*}\nabla)_{q\partial_{q}},\ (\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}},\ (\sigma_{j}^{*}\nabla)_{z\partial_{z}}\colon\sigma_{j}^{*}\operatorname{QDM}(B)\to z^{-1}\sigma_{j}^{*}\operatorname{QDM}(B).$
The operator $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ has poles along the infinity disivor $(q^{-1/r^{\prime}}=0)$, but this is an apparent singularity: $\partial_{\tau^{i,k}}$ viewed as a vector field on $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ has poles of the same order along $(q^{-1/r^{\prime}}=0)$.
We introduce a localized version along the infinity disivor $(q^{-1/r^{\prime}}=0)$ as follows:
$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ $:=\sigma_{j}^{*}\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$
$=H^{*}(B)[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!].$
Proposition 5.9.
The Fourier transformation $\mathscr{F}_{j}$ induces a map $\Pi_{j}\colon\operatorname{QDM}_{S^{1}}(V)\to\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules such that
(5.5) $\mathscr{F}_{j}(M_{V}(\tau)s)=M_{B}(\sigma_{j})\Pi_{j}(s)\qquad\text{for all $s\in\operatorname{QDM}_{S^{1}}(V)$.}$
The map $\Pi_{j}$ satisfies the following:
  • (1)
    $\Pi_{j}$ intertwines $\mathbb{S}(\tau)$ with $q$;
  • (2)
    $\Pi_{j}$ intertwines $\lambda$ with $\lambda_{j}+z(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$;
  • (3)
    $\Pi_{j}$ intertwines $\nabla_{\tau^{i,k}}$ with $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$;
  • (4)
    $\Pi_{j}$ intertwines $\nabla_{\xi Q\partial_{Q}}$ with $(\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}}$ for $\xi\in H^{2}(B)$;
  • (5)
    $\Pi_{j}$ intertwines $\nabla_{z\partial_{z}}+\frac{1}{2}$ with $(\sigma_{j}^{*}\nabla)_{z\partial_{z}}-r\lambda_{j}/z$;
  • (6)
    $\Pi_{j}$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(B)$ and the degree of the variables as in Proposition 5.6;
  • (7)
    $\Pi_{j}(\phi_{i}\lambda^{k})\in\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}\left(\phi_{i}+\mathfrak{m}H^{*}(B)[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]\right)$
where $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$ and $\mathfrak{m}$ is as in Proposition 5.6.
Proof.
By differentiating the identity $\mathscr{F}_{j}(M_{V}(\tau)1)=M_{B}(\sigma_{j})\upsilon_{j}$ from Proposition 5.6 by $\tau^{i,k}$ and using the differential equations (2.7) for $M_{V}$ and $M_{B}$, we obtain
$\mathscr{F}_{j}(M_{V}(\tau)\phi_{i}\lambda^{k})=M_{B}(\sigma_{j})z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}.$
Therefore the $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-module map $\Pi_{j}$ sending $\phi_{i}\lambda^{k}$ to $z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}$ satisfies (5.5). Parts (1)-(5) follow immediately from (5.5), Proposition 5.5 and the intertwining properties (2.7), (2.21) of $M_{V},M_{B}$. Part (6) follows from the homogeneity of $\sigma_{j}+c_{1}(V)\log\lambda_{j}$, $\upsilon_{j}$ in Proposition 5.6. Part (7) follows by a direct computation of
$\Pi_{j}(\phi_{i}\lambda^{k})=z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}=z\frac{\partial\upsilon_{j}}{\partial\tau^{i,k}}+\frac{\partial\sigma_{j}}{\partial\tau^{i,k}}\star_{\sigma_{j}}\upsilon_{j}$
using Proposition 5.6. ∎
5.6. End of the proof
We finish the proof of Theorem 5.1. We combine the isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\cong\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ in Theorem 4.2 with the maps $\Pi_{j}$ in the above proposition to obtain a map
$\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}.$
By Theorem 4.2 and Proposition 5.9, this is a homomorphism of $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$-modules homogeneous of degree $-(r-1)$, and intertwines the connection $\hat{\tau}^{*}\nabla$ with
$\bigoplus_{j=0}^{r-1}\left(\sigma_{j}^{*}\nabla+\frac{\lambda_{j}}{z}\frac{dq}{q}-r\lambda_{j}\frac{dz}{z^{2}}\right).$
We note that the pull-back $(\sigma_{j}+r\lambda_{j})^{*}\nabla$ of the quantum connection of $B$ by $\sigma_{j}+r\lambda_{j}$ is well-defined [^8] Although $M_{B}(\sigma_{j}+r\lambda_{j})$ is ill-defined in the $q^{-1/r}$-adic topology. and that
$\sigma_{j}^{*}\nabla+\frac{\lambda_{j}}{z}\frac{dq}{q}-r\lambda_{j}\frac{dz}{z^{2}}=(\sigma_{j}+r\lambda_{j})^{*}\nabla.$
Here we use the fact that the shift $\sigma_{j}\to\sigma_{j}+r\lambda_{j}$ in the $H^{0}(B)$-direction does not change the quantum product but affects the Euler vector field (2.3) and the covariant derivative in the $q$-direction (through $q\partial_{q}(r\lambda_{j})=\lambda_{j}$). Hence we obtain a map
(5.6) $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm loc}$
that is fully compatible with the connection. We now restrict the parameter $\tau$ to lie in $\mathsf{H}:=\bigoplus_{k=0}^{r-1}H^{*}(B)\lambda^{k}\subset H^{*}_{S^{1}}(V)$: the mirror map $\tau\mapsto\hat{\tau}(\tau)$ in Theorem 4.2 restricted to $\mathsf{H}$ gives an isomorphism between $\mathsf{H}$ and $H^{*}(\mathbb{P}(V))$ in the formal neighbourhood of $q=Q=0$ since $\hat{\tau}(\tau)|_{q=Q=0}=\kappa(\tau)$. We compose the map $\sigma_{j}$ with the inverse $\hat{\tau}\mapsto\tau=\tau(\hat{\tau})$ of this isomorphism and define
$\varsigma_{j}(\hat{\tau}):=\sigma_{j}(\tau(\hat{\tau}))+r\lambda_{j}.$
Because of the homogeneity of $\hat{\tau}(\tau)$ from Theorem 4.2(6), the inverse mirror map $\tau(\hat{\tau})$ belongs to $H^{*}(B)[q][\![Q,\hat{\tau}]\!]$. Therefore the composition is well-defined; $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ belongs to $H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$ and is homogeneous of degree $2$. Moreover we have
(5.7) $\varsigma_{j}(\hat{\tau})\Bigr{|}_{Q=0}\in r\lambda_{j}-c_{1}(V)\log\lambda_{j}+\alpha+\sum_{k=0}^{r-1}\sum_{i=0}^{s}\hat{\tau}^{i,k}\phi_{i}\lambda_{j}^{k}(1+\beta_{k})+(\hat{\tau})^{2}H^{*}(B)$
for some $\alpha,\beta_{k}\in\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1}]\!]$, where $(\hat{\tau})\subset\mathbb{C}(\!(q^{-1/r})\!)[\![\hat{\tau}]\!]$ denote the ideal generated by $\hat{\tau}^{i,k}$ with $0\leq i\leq s$, $10\leq k\leq r-$ 1. This follows by Theorem 4.2(7) and Proposition 5.6. The above map induces a $\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$-module map
$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
compatible with the connection and homogeneous of degree $-(r-1)$. Combining Theorem 4.2(8) and Proposition 5.9(7), we have
$\Phi_{j}(\phi_{i}p^{k})\Bigr{|}_{Q=\hat{\tau}=0}\in\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}\left(\phi_{i}+q^{-1/r}H^{*}(B)[z][\![q^{-1/r}]\!]\right)$
for $10\leq k\leq r-$ 1 and therefore $\Phi$ induces an isomorphism over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$. Inverting $\Phi|_{Q=\hat{\tau}=0}$, we obtain
(5.8) $\Phi^{-1}(\phi_{i}e_{k})\Bigr{|}_{Q=\hat{\tau}=0}\in\frac{1}{\sqrt{r}}\sum_{j=0}^{r-1}\lambda_{k}^{(r-1)/2-j}p^{j}\left(\phi_{i}+q^{-1/r}H^{*}(B)[z][\![q^{-1/r}]\!]\right)$
where $e_{k}$ is the vector $(0,\dots,0,1,0,\dots,0)$ (with $1$ in the $k$th position) in $\bigoplus_{j=0}^{r-1}H^{*}(B)$.
It remains to show that $\Phi$ preserves the pairing. The asymptotics (5.8) shows that
$P_{\mathbb{P}(V)}$ $(\Phi^{-1}(\phi_{i}e_{k}),\Phi^{-1}(\phi_{h}e_{l}))\Bigr{|}_{Q=\hat{\tau}=0}$
$=\frac{1}{r}\sum_{j+j^{\prime}\geq r-1}\lambda_{k}^{(r-1)/2-j}\lambda_{l}^{(r-1)/2-j^{\prime}}\left(\int_{\mathbb{P}(V)}\phi_{i}\phi_{h}p^{j+j^{\prime}}+O(q^{-1/r})\right)$
$=P_{B}(\phi_{i}e_{k},\phi_{h}e_{l})+O(q^{-1/r}).$
Writing in the $\mathbb{C}$-bases $\{\phi_{i}p^{k}\}$, $\{\phi_{i}e_{k}\}$ of $H^{*}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}H^{*}(B)$, we regard $P_{1}=P_{\mathbb{P}(V)}$, $P_{2}=\bigoplus_{j=0}^{r-1}P_{B}$ and $\Phi$ as square matrices of size $r\dim H^{*}(B)$. Here $P_{1}$ and $P_{2}$ are constant matrices and $\Phi$ has entries in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$. The above computation shows that
$P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=I+O(q^{-1/r})$
where $I$ is the identity matrix. Both of the pairings $P_{2}$ and $\Phi^{-T}P_{1}\Phi^{-1}$ are compatible with the connection $\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\nabla$ (see (2.9)) and hence $P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}$ is flat with respect to the connection induced on the endomorphism bundle. By (5.7) we have
$(\varsigma_{j}^{*}\nabla)_{q\partial_{q}}\Bigr{|}_{Q=\hat{\tau}=0}=q\partial_{q}+\frac{1}{z}\left(\lambda_{j}+\frac{c_{1}(V)}{r}+\alpha^{\prime}\right)$
where $c_{1}(V)$ and $\alpha^{\prime}=q\partial_{q}\alpha\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$ acts by cup product. Writing $x=q^{-1/r}$, we find that $v(x):=P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=v_{0}+v_{1}x+v_{2}x^{2}+\cdots$ satisfies a differential equation of the form
$(x\partial_{x}+x^{-1}S+(N_{0}+N_{1}x+\cdots))v(x)=0$
where $S,N_{i}$ are mutually commuting endomorphisms on $\operatorname{End}(\bigoplus_{j=0}^{r-1}H^{*}(B)))$ such that $S$ is semisimple and $N_{0}$ is nilpotent. We obtain a sequence of equations
$Sv_{n+1}+(n+N_{0})v_{n}+N_{1}v_{n-1}+N_{2}v_{n-2}+\cdots=0.$
From $Sv_{0}=N_{i}v_{0}=0$ and (*$*_{0}$) we have $Sv_{1}=0$. Then (*$*_{1}$) implies $S^{2}v_{2}=0$. Since $S$ is semisimple, $Sv_{2}=0$ and (*$*_{1}$) implies $(1+N_{0})v_{1}=0$; the nilpotence of $N_{0}$ implies $v_{1}=0$. In this way we can show inductively that $v_{1}=v_{2}=\cdots=0$. Therefore $P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=I$; in other words,
$P_{1}^{-1}\Phi^{T}P_{2}\Phi\Bigr{|}_{Q=\hat{\tau}=0}=I.$
We now use the fact that $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ is flat with respect to the connection on the endomorphism bundle of $\operatorname{QDM}(\mathbb{P}(V))$. It satisfies a differential equation in $(Q,\hat{\tau})$ with at most logarithmic singularities at $Q=0$ and nilpotent residues. Expanding $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ in power series in $Q$ and $\hat{\tau}$ and using a similar argument as before, we can easily show that $P_{1}^{-1}\Phi^{T}P_{2}\Phi=I$ identically. Thus $\Phi$ preserves the pairing.
Remark 5.10.
The map (5.6) glues the connections $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$; $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ is defined around $q=Q=\boldsymbol{\tau}=0$ and $(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$ is defined around $q^{-1/r^{\prime}}=q^{-c_{1}(V)/r}Q=q^{\bullet/r}\boldsymbol{\tau}=0$. They glue to a connection over a global “Kähler moduli space” obtained by adding the infinity divisor $(q=\infty)$ to $\operatorname{Spf}(\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!])$. See Figure 1 for the global Kähler moduli space.
5.7. Semisimplicity
Since the decomposition $\Phi$ in Theorem 5.1 intertwines $\nabla_{\hat{\tau}^{i,k}}$ with $\bigoplus_{j=0}^{r-1}(\varsigma_{j}^{*}\nabla)_{\hat{\tau}^{i,k}}$, it follows that
$(\phi_{i}p^{k})\star_{\hat{\tau}}=\Phi^{-1}\left(\textstyle\bigoplus_{j=0}^{r-1}(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})\star_{\varsigma_{j}(\hat{\tau})}\right)\Phi+z\Phi^{-1}\partial_{\hat{\tau}^{i,k}}\Phi.$
Setting $z=0$, this shows that the map sending $\phi_{i}p^{k}$ to $\bigoplus_{j=0}^{r-1}\partial_{\hat{\tau}^{i,k}}\varsigma_{j}$ gives a ring isomorphism $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{j=0}^{r-1}(H^{*}(B),\star_{\varsigma_{j}(\hat{\tau})})$ over $\mathbb{C}(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$. This proves Corollary 1.8.
Corollary 1.8 can be applied to show the semisimplicity of big quantum cohomology of certain partial flag varieties. Let $Fl(k_{1},\dots,k_{l};n)$ denote the partial flag variety consisting of flags $V_{1}\subset V_{2}\subset\cdots\subset V_{l}\subset\mathbb{C}^{n}$ with $\dim V_{i}=k_{i}$. Fix a non-degenerate skew-symmetric 2-form $\omega$ on $\mathbb{C}^{2n}$. Let $IFl(k_{1},\dots,k_{l};2n)$ denote the isotropic partial flag variety consisting of flags $V_{1}\subset V_{2}\subset\cdots\subset V_{l}\subset\mathbb{C}^{2n}$ such that $\dim V_{i}=k_{i}$ and that $V_{l}$ is isotropic with respect to $\omega$. Here we necessarily have $k_{l}\leq n$. The varieties $G(k,n)=Fl(k;n)$ or $IG(k,2n)=IFl(k;2n)$ are also known as ordinary or isotropic Grassmannians. We have the following result.
Proposition 5.11.
The partial flag varieties $Fl(k_{1},\dots,k_{l};n)$ and $IFl(k_{1},\dots,k_{l};2n)$ have generically semisimple big quantum cohomology.
Proof.
In view of Corollary 1.8, it suffices to show that these spaces are connected to a point by a sequence of projective bundles. For example, $IG(3,6)=IFl(3;6)$ fits into the sequence:
$IFl(3;6)\xleftarrow{\mathbb{P}^{2}}IFl(2,3;6)\xrightarrow{\mathbb{P}^{1}}IFl(2;6)\xleftarrow{\mathbb{P}^{1}}IFl(1,2;6)\xrightarrow{\mathbb{P}^{3}}IFl(1;6)=\mathbb{P}^{5}\to\operatorname{pt}.$
The general case is similar. ∎
Remark 5.12.
It is expected that rational homogeneous varieties $G/P$ have semisimple big quantum cohomology. We refer the reader to [46] and references therein for recent developments on this question.
$Q=\hat{\tau}=0$, $q>0$small deformation of $(Q,\hat{\tau})$
Figure 2. Decomposition of eigenvalues of the Euler multiplication of $\mathbb{P}(V)$ into $r$ groups ($r=6$ in this picture).
Remark 5.13.
We can describe a decomposition of the eigenvalues of the quantum multiplication by the Euler vector field of $\mathbb{P}(V)$ along $Q=\hat{\tau}=0$. Because $z\nabla_{z\partial_{z}}=E_{\mathbb{P}(V)}\star_{\hat{\tau}}+O(z)$ for the quantum connection $\nabla$ of $\mathbb{P}(V)$, Theorem 5.1 implies that the eigenvalues of $E_{\mathbb{P}(V)}\star_{\hat{\tau}}$ are the union of the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$, $j=0,\dots,r-1$. When $Q=\hat{\tau}=0$, the quantum product $\star_{\varsigma_{j}(\hat{\tau})}$ for $B$ equals the cup product and the eigenvalues of $E_{B}\star_{\varsigma_{j}(\hat{\tau})}$ are all equal to the $H^{0}$ component $r\lambda_{j}=re^{2\pi\sqrt{-1}j/r}q^{1/r}$ of $E_{B}$ along $Q=\hat{\tau}=0$ (see Theorem 5.1(4)). See Figure 2 for a typical decomposition of eigenvalues. Similar pictures have appeared in [28, Figure 16], [34, Figure 1].

### v4

@@
 1. Introduction
-For a rank-$r$ vector bundle $V\to B$ over a smooth projective variety $B$, we can construct the projective bundle $\mathbb{P}(V)\to B$ consisting of lines in the fibres of $V\to B$.
+For a rank-$r$ vector bundle $V\to B$ over a smooth projective variety $B$, we can construct the projective bundle $\pi\colon\mathbb{P}(V)\to B$ consisting of lines in the fibres of $V\to B$.
 The ordinary cohomology group $H^{*}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is generated by the relative hyperplane class $p:=c_{1}(\mathcal{O}_{\mathbb{P}(V)}(1))$ as an $H^{*}(B)$-algebra as follows:
@@
 $H^{*}(B)[p,q]/(p^{r}+c_{1}(V)p^{r-1}+\cdots+c_{r}(V)-q)$
-where $q$ is the Novikov variable associated with vertical curves.
-It decomposes into a direct sum of $H^{*}(B)$ as a ring for $q\neq 0$.
-This decomposition must persist under full quantum deformation, parametrized by $q$, the Novikov variable $Q$ of $B$ and the bulk parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$.
+where $q$ is the Novikov variable associated with vertical lines.
+For $q\neq 0$, this ring decomposes into a direct sum of $r$ copies of $H^{*}(B)$.
+This decomposition must persist under the full quantum deformation, parametrized by the full Novikov variables of $\mathbb{P}(V)$ and the bulk-deformation parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$.
 The question then arises as to whether each summand is isomorphic to the quantum cohomology of $B$.
 This paper aims to affirm this correspondence, i.e.
-$QH^{*}(\mathbb{P}(V))$ decomposes into a direct sum of $QH^{*}(B)$.
+$QH^{*}(\mathbb{P}(V))$ decomposes into a direct sum of $r$ copies of $QH^{*}(B)$.
 Furthermore, we show that the decomposition can be lifted to quantum cohomology $D$-modules.
@@
 Remark 1.3.
-The equivariant $J$-function $J_{V}^{\lambda}$ is defined over the Novikov ring of $B$.
-We use the splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})$ given by the class $p=c_{1}(\mathcal{O}(1))$ to identify Novikov variables for $B$ with Novikov variables for $\mathbb{P}(V)$, see Notation 3.1.
+The $J$-function $J_{V}^{\lambda}$ is defined over the Novikov ring of $V$.
+In Theorem 1.1, we use the splitting of $\pi_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})=H_{2}(V,\mathbb{Z})$ given by the class $p=c_{1}(\mathcal{O}(1))$ to identify Novikov variables for $V$ with those for $\mathbb{P}(V)$, see Notation 3.1.
 Remark 1.4.
@@
 Therefore Theorem 1.1 determines all genus-zero Gromov-Witten invariants of $\mathbb{P}(V)$ from those of the base $B$.
-This recovers the genus-zero part of the results of Fan-Lee [16] and Fan [15], who showed that all genus Gromov-Witten invariants of $\mathbb{P}(V)$ can be reconstructed from those of the base $B$ and the Chern classes $c_{i}(V)$.
+This recovers the genus-zero part of the results of Fan-Lee [16] and Fan [15], who showed that all genus [^1] The method in this paper is not applicable to higher-genus reconstruction, as it relies on the quantum Lefschetz principle [34] for convex bundles.
+Gromov-Witten invariants of $\mathbb{P}(V)$ can be reconstructed from those of the base $B$ and the Chern classes $c_{i}(V)$.
 On the other hand, in contrast to the split case, it is not easy to find a closed formula for $J_{V}^{\lambda}$ in general.
@@
 Remark 1.5.
-For other techniques to compute Gromov-Witten invariants of non-split projective bundles, we refer the reader to the reconstruction algorithm [49] and the quantum splitting principle [36, 14].
+Several techniques exist for computing the Gromov-Witten invariants of non-split projective bundles.
+Strangeway [52] provided a reconstruction algorithm for certain projective bundles over $\mathbb{P}^{n}$.
+The quantum splitting principle of Lee-Lin-Qu-Wang [39] proposes a reduction to the split case by blowing up the base.
+Fan [14, Theorem 0.1, Question 4] explored a relationship between the $(V,e_{\lambda}^{-1})$-twisted Gromov-Witten invariants of $B$ and the $(\mathcal{O}(-1),e_{\lambda}^{-1})$-twisted (or untwisted) invariants of $\mathbb{P}(V)$, under the opposite assumption that $V$ is globally generated.
+Recently, Hinault-Yu-Zhang-Zhang [23] (see also earlier announcements [35, 36] by Katzarkov-Kontsevich-Pantev-Yu) proposed a reconstruction method based on the decomposition theorem (Theorem 1.7 below).
+The initial condition required for their reconstruction and the explicit algorithm based on Birkhoff factorization are provided in §5.8.
 Remark 1.6.
@@
 These covariant derivatives, called the quantum connection, are defined in terms of the quantum multiplication.
-The quantum $D$-module is also equipped with a $z$-sesquilinear pairing $P$, induced from the Poincaré pairing and compatible with the quantum connection.
+The quantum $D$-module is further equipped with a $z$-sesquilinear pairing $P$, induced from the Poincaré pairing and compatible with the quantum connection.
 The $I$- or $J$-functions appearing in the previous section are cohomology-valued solutions of the quantum $D$-module.
-The Mirror Theorem 1.1 establishes that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ is a Fourier transform of the $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ (see Theorem 1.11 below).
-Using this fact together with the quantum Riemann-Roch (QRR) theorem of Coates-Givental [7], which connects the quantum $D$-modules of $V$ and $B$, we derive the following decomposition theorem for $\operatorname{QDM}(\mathbb{P}(V))$.
-$Q$$q^{-c_{1}(V)/r}Q$$q$$q^{-1/r^{\prime}}$(($\operatorname{QDM}(\mathbb{P}(V))$$\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)$
+The Mirror Theorem 1.1 establishes that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ is the Fourier transform of the $S^{1}$-equivariant quantum $D$-module $\operatorname{QDM}_{S^{1}}(V)$ (see Theorem 1.12 below).
+By combining this result with the quantum Riemann-Roch (QRR) theorem of Coates-Givental [7], which connects the Gromov-Witten invariants of $V$ and $B$, we prove that $\operatorname{QDM}(\mathbb{P}(V))$ decomposes into a direct sum of $r$ copies of $\operatorname{QDM}(B)$ after an appropriate localization.
+To compare the quantum $D$-modules of $\mathbb{P}(V)$ and $B$, we embed the Novikov ring of $B$ into a localization of the Novikov ring of $\mathbb{P}(V)$.
+Let $(q,Q)$ denote the Novikov variable of $\mathbb{P}(V)$, where $q$ corresponds to the class of a vertical line and $Q$ is the Novikov variable of the base $B$, identified via the splitting in Notation 3.1.
+Introducing an additional copy $Q_{B}$ of the Novikov variable for $B$, we define the following embedding [^2] See Remark 5.2 for the geometric interpretation of this embedding.
+of Novikov rings:
+(1.1) $\mathbb{C}[\![Q_{B}]\!]\hookrightarrow\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]\qquad Q_{B}^{d}\mapsto q^{-c_{1}(V)\cdot d/r}Q^{d}$
+where $r^{\prime}$ is $r$ or $2r$ depending on the parity of $r$ (see (5.1)).
+Let $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ denote the base change of $\operatorname{QDM}(\mathbb{P}(V))$ to the localized base $\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$, and let $\operatorname{QDM}(B)_{\rm ext,loc}$ denote the base change of $\operatorname{QDM}(B)$ via the embedding (1.1).
+[…]
 Figure 1.
-Decomposition pictured over the Kähler moduli space: $\operatorname{QDM}(\mathbb{P}(V))$ splits into a direct sum of $\operatorname{QDM}(B)$ near the infinity divisor $(q^{-1/r^{\prime}}=0)$, where $Q$ is the Novikov variable of the base $B$ and $q$ is the Novikov variable of the fibre $\mathbb{P}^{r-1}$.
-Theorem 1.7 (Theorem 5.1; see also Figure 1).
+Decomposition pictured over the global Kähler moduli space: $\operatorname{QDM}(\mathbb{P}(V))$ splits into a direct sum of $\operatorname{QDM}(B)$ near the infinity divisor $(q^{-1/r^{\prime}}=0)$.
+Theorem 1.7 (Theorem 5.1).
 Let $V\to B$ be as in Theorem 1.1.
-We have a map $\varsigma=\bigoplus_{j=0}^{r-1}\varsigma_{j}\colon H^{*}(\mathbb{P}(V))\to\bigoplus_{j=0}^{r-1}H^{*}(B)$ and a decomposition of the quantum $D$-modules
-$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{i=0}^{r-1}\varsigma_{i}^{*}\operatorname{QDM}(B)_{\rm loc}$
-where $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ and $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ denote the localized quantum $D$-modules of $\mathbb{P}(V)$ and $B$ defined over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$, where $Q$ is the Novikov variable of the base $B$ and $r^{\prime}$ is in (5.1).
-Corollary 1.8.
-The derivative of the map $\varsigma$ induces an isomorphism of the quantum cohomology rings $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{i=0}^{r-1}(H^{*}(B),\star_{\varsigma_{i}(\hat{\tau})})$.
-In particular, the quantum cohomology of $\mathbb{P}(V)$ is generically semisimple if and only if the same holds for the base $B$.
+We have a formal invertible map $\varsigma=\bigoplus_{j=0}^{r-1}\varsigma_{j}\colon H^{*}(\mathbb{P}(V))\to\bigoplus_{j=0}^{r-1}H^{*}(B)$ over $\mathbb{C}(\!(q^{-1/r})\!)[\![Q]\!]$ and a decomposition of the quantum $D$-modules over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$:
+$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{i=0}^{r-1}\varsigma_{i}^{*}\operatorname{QDM}(B)_{\rm ext,loc}.$
+The pairs $(q,Q)$ and $(q^{-1/r^{\prime}},Q_{B})$ of Novikov variables, related by $Q_{B}=q^{-c_{1}(V)/r}Q$, may be viewed as two coordinate charts on the global Kähler moduli space illustrated in Figure 1.
+The isomorphism $\Phi$ glues the quantum $D$-modules $\operatorname{QDM}(\mathbb{P}(V))$ and $\operatorname{QDM}(B)_{\rm ext}^{\boxplus r}=\operatorname{QDM}(B)[\![q^{-1/r^{\prime}}]\!]^{\boxplus r}$ over the intersection of their domains of definition — the localized base $\operatorname{Spf}\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$ — to form a global $D$-module.
+Corollary 1.8 (see §5.7 for the proof).
+[…]
 We can use Corollary 1.8 to deduce the generic semisimplicity of quantum cohomology for certain Grassmannians $G(k,n)$, $IG(k,2n)$ (see Proposition 5.10).
@@
 In the split case, a similar decomposition of the quantum $D$-module (more generally, for toric bundles) follows from Brown’s theorem [3]: this has been discussed by Coates-Givental-Tseng [8] and Koto [38].
-These previous works rely on the equivariant localization with respect to the (×$(\mathbb{C}^{\times})^{r-1}$-action (which exists only in the split case) and our argument is substantively different from them.
+These previous works rely on the equivariant localization with respect to the $(\mathbb{C}^{\times})^{r-1}$-action (which exists only in the split case) and our argument is substantively different from them.
+Remark 1.11 (see [28, 35, 30]).
+Theorem 1.7 provides a decomposition of quantum $D$-modules at the formal level.
+Given that quantum cohomology is expected to have convergent structure constants, we anticipate that the mirror map $\varsigma_{j}$ and the isomorphism $\Phi$ in Theorem 1.7 can be analytified with respect to the Novikov variables $q,Q$ and the bulk parameter $\hat{\tau}\in H^{*}(\mathbb{P}(V))$.
+However, we should not expect $\Phi$ to be analytic in $z$; rather, its matrix entries should belong to $\mathcal{O}^{\rm an}[\![z]\!]$, where $\mathcal{O}^{\rm an}$ denotes the sheaf of analytic functions in $q,Q,\hat{\tau}$.
+This formal (in $z$) decomposition should lift to an analytic decomposition over an angular sector in the $z$-plane.
+Furthermore, we expect this analytic decomposition to relate — at specific parameter values — to the semi-orthogonal decomposition [46] of the derived category of coherent sheaves:
+$D^{b}(\mathbb{P}(V))=\langle D^{b}(B)_{0},\dots,D^{b}(B)_{r-1}\rangle,\qquad D^{b}(B)_{i}=D^{b}(B)\otimes\mathcal{O}(i)$
+via the $\widehat{\Gamma}-$integral structure.
 1.3. Equivariant quantum cohomology and Fourier transformation
@@
 The equivariant volume of $M$ can be written as a Fourier transform of the volume of $M/\!\!/_{t}S^{1}$:
-[…]
+$\int_{M}e^{\omega-\lambda\mu}=\int_{\mathbb{R}}e^{-\lambda t}dt\int_{M/\!\!/_{t}S^{1}}e^{\omega_{\rm red}}$
 where $\lambda$ is the $S^{1}$-equivariant parameter and $\omega_{\rm red}$ is the reduced symplectic form on $M/\!\!/_{t}S^{1}$.
@@
 We have the following commutation relation:
-$[\lambda,\mathbb{S}]=z\mathbb{S}.$
+$\lambda,\mathbb{S}]=z\mathbb{S}.$
 Now we change the point of view and regard $\operatorname{QDM}_{S^{1}}(M)$ as a $\mathbb{C}[\mathbb{S}]$-module rather than a $\mathbb{C}[\lambda]$-module; the equivariant parameter $\lambda$ then behaves like the differential operator $z\mathbb{S}\frac{\partial}{\partial\mathbb{S}}$.
@@
 This gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1 restricted to $t=0$.
-We can rephrase the mirror theorem (Theorems 1.1, 3.3) as the Fourier duality between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$, as follows:
+We can rephrase the mirror theorem (Theorems 1.1, 3.4) as the Fourier duality between the quantum $D$-modules of $V$ and $\mathbb{P}(V)$.
+In this paper, we treat $\operatorname{QDM}_{S^{1}}(V)$ as a $D$-module over the infinite-dimensional base $H^{*}_{S^{1}}(V)$ (see §2.3 and §4.1.2 for further details).
 Theorem 1.12 (Theorem 4.2, Proposition 4.3).
@@
 thanks Fumihiko Sanda, Yuuji Tanaka and Constantin Teleman for very useful discussions on shift operators and equivariant cohomology.
+We thank anonymous referees for many valuable comments.
 H.I.
@@
 For $n\geq 0$ and $d\in H_{2}(X,\mathbb{Z})$, let $X_{0,n,d}$ denote the moduli stack of genus-zero, $n$-points, degree $d$ stable maps to $X$.
-This is a proper Deligne-Mumford stack carrying a virtual fundamental class $[X_{0,n,d}]_{\rm vir}\in H_{2D}(X_{0,n,d},\mathbb{Q})$ of degree $D=\dim_{\mathbb{C}}X+c_{1}(TX)\cdot d+n-3$.
+This is a proper Deligne-Mumford stack carrying a virtual fundamental class $X_{0,n,d}]_{\rm vir}\in H_{2D}(X_{0,n,d},\mathbb{Q})$ of degree $D=\dim_{\mathbb{C}}X+c_{1}(TX)\cdot d+n-3$.
 Let $\operatorname{ev}_{i}\colon X_{0,n,d}\to X$ be the evaluation map at the $i$th marked point.
@@
 Let $\operatorname{Eff}(X)\subset H_{2}(X,\mathbb{Z})$ denote the semigroup of effective curve classes.
-For a module $K$, we write $K[\![Q]\!]$ for the space of formal power series $\sum_{d\in\operatorname{Eff}(X)}k_{d}Q^{d}$ with $k_{d}\in K$.
+For a module $K$, we write K[[Q]]K[\![Q]\!] for the space of formal power series $\sum_{d\in\operatorname{Eff}(X)}k_{d}Q^{d}$ with $k_{d}\in K$.
 The formal variable $Q$ is called the Novikov variable.
-The module $K[\![Q]\!]$ has a natural ring structure when $K$ is a ring.
+The module K[[Q]]K[\![Q]\!] has a natural ring structure when $K$ is a ring.
 The quantum cohomology will be defined over the ring $\mathbb{C}[\![Q]\!]$, which is called the Novikov ring.
@@
 for $\alpha,\beta,\gamma\in H^{*}(X)$.
+We call $\tau$ the bulk(-deformation) parameter.
 The product $\star_{\tau}$ introduces an associative and supercommutative ring structure on $H^{*}(X)[\![Q,\tau]\!]=H^{*}(X)[\![Q]\!][\![\tau^{0},\dots,\tau^{s}]\!]$ such that $1\in H^{0}(X)$ is the unity.
@@
 $\nabla_{\tau^{i}},\ \nabla_{z\partial_{z}},\ \nabla_{\xi Q\partial_{Q}}\colon H^{*}(X)[z][\![Q,\tau]\!]\to z^{-1}H^{*}(X)[z][\![Q,\tau]\!].$
-The quantum connection admits the fundamental solution $M_{X}(\tau)$ defined as follows (see [19, §1], [44, Proposition 2], [25, Proposition 2.4], [6, Proposition 3.1] for the properties below):
+The quantum connection admits the fundamental solution $M_{X}(\tau)$ defined as follows (see [19, §1], [47, Proposition 2], [25, Proposition 2.4], [6, Proposition 3.1] for the properties below):
 (2.6) $(M_{X}(\tau)\phi_{i},\phi_{j})_{X}=(\phi_{i},\phi_{j})_{X}+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0)\end{subarray}}\left\langle\phi_{i},\tau,\dots,\tau,\frac{\phi_{j}}{z-\psi}\right\rangle_{0,n+2,d}^{X}\frac{Q^{d}}{n!}$
@@
 where $\{\phi^{j}\}$ is the dual basis of $\{\phi_{i}\}$ such that $(\phi_{i},\phi^{j})_{X}=\delta_{i}^{j}$.
-We used the String Equation [44, §1.2] in the second line.
+We used the String Equation [47, §1.2] in the second line.
 The other columns $M_{X}(\tau)\phi_{i}$ of $M_{X}(\tau)$ are given by the derivatives $z\partial_{\tau^{i}}J_{X}(\tau)$ of the $J$-function due to the differential equation (2.7).
@@
 $\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle_{0,n,d}^{V}=\int_{[V_{0,n,d}]_{\rm vir}}^{\mathbb{C}^{\times}}\prod_{i=1}^{n}\operatorname{ev}_{i}^{*}(\alpha_{i})\psi_{i}^{k}$
-with $\alpha_{1},\dots,\alpha_{n}\in H^{*}_{\mathbb{C}^{\times}}(V)=H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ by the virtual localization formula [22], where $\lambda$ is the $\mathbb{C}^{\times}$-equivariant (or $S^{1}$-equivariant) parameter [^1] In this paper, all $S^{1}$-actions are induced from $\mathbb{C}^{\times}$-actions, so we use the terms ‘$S^{1}$-equivariant’ and ‘$\mathbb{C}^{\times}$-equivariant’ interchangeably..
+with $\alpha_{1},\dots,\alpha_{n}\in H^{*}_{\mathbb{C}^{\times}}(V)=H^{*}_{S^{1}}(V)=H^{*}(B)\otimes\mathbb{C}[\lambda]$ by the virtual localization formula [22], where $\lambda$ is the $\mathbb{C}^{\times}$-equivariant (or $S^{1}$-equivariant) parameter [^3] In this paper, all $S^{1}$-actions are induced from $\mathbb{C}^{\times}$-actions, so we use the terms $S^{1}$-equivariant’ and $\mathbb{C}^{\times}$-equivariant’ interchangeably..
 The equivariant Gromov-Witten invariants take values in $\mathbb{C}[\lambda,\lambda^{-1}]$.
@@
 $\nabla_{\tau^{i,k}},\ \nabla_{\xi Q\partial_{Q}},\ \nabla_{z\partial_{z}}\colon H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]\to z^{-1}H^{*}_{S^{1}}(V)[z][\![Q,\boldsymbol{\tau}]\!]$
-where $K[\![\boldsymbol{\tau}]\!]$ denotes the space of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}$ with coefficients in $K$ for a module $K$ (we use the boldface symbol $\boldsymbol{\tau}$ to distinguish infinitely many parameters $\{\tau^{i,k}\}$ from finitely many ones $\{\tau^{i}\}$).
+where $K[\![\boldsymbol{\tau}]\!]$ denotes the space of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$ with coefficients in $K$ for a module $K$ (we use the boldface symbol $\boldsymbol{\tau}$ to distinguish infinitely many parameters $\{\tau^{i,k}\}$ from finitely many ones $\{\tau^{i}\}$).
 We can define the fundamental solution $M_{V}(\tau)$ for $V$ similarly to (2.6) by replacing the Poincaré pairing and the Gromov-Witten invariants with their equivariant counterparts.
@@
 where $\mathbf{t}(z)=\sum_{n\geq 0}t_{n}z^{n}$ with $t_{n}=\sum_{i=0}^{s}t^{i}_{n}\phi_{i}\in H^{*}(X)$.
-The argument $-z+\mathbf{t}(z)$ represents a point in the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$ and $\mathcal{F}_{X}$ is regarded as a function on the formal neighbourhood of $-z$ in $\mathcal{H}_{+}$.
+The argument $z+\mathbf{t}(z)$ represents a point in the formal neighbourhood of z-z in $\mathcal{H}_{+}$ and $\mathcal{F}_{X}$ is regarded as a function on the formal neighbourhood of z-z in $\mathcal{H}_{+}$.
 The Givental Lagrangian cone $\mathcal{L}_{X}^{\rm orig}\subset\mathcal{H}_{X}$ is a submanifold-germ defined as the graph of $d\mathcal{F}_{X}$ in $T^{*}\mathcal{H}_{+}\cong\mathcal{H}_{X}$.
 It consists of
-(2.14) $-z+\mathbf{t}(z)+\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\frac{\partial\mathcal{F}_{X}}{\partial t_{k}^{i}}\\ =-z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{X}\frac{Q^{d}}{n!}$
+(2.14) $z+\mathbf{t}(z)+\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\frac{\partial\mathcal{F}_{X}}{\partial t_{k}^{i}}\\ =-z+\mathbf{t}(z)+\sum_{\begin{subarray}{c}d\in\operatorname{Eff}(X),n\geq 0\\ (d,n)\neq(0,0),(0,1)\end{subarray}}\sum_{k\geq 0}\sum_{i=0}^{s}\frac{\phi^{i}}{(-z)^{k+1}}\left\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi),\phi_{i}\psi^{k}\right\rangle_{0,n+1,d}^{X}\frac{Q^{d}}{n!}$
 with $\mathbf{t}(z)\in\mathcal{H}_{+}$.
@@
   • (a)
-    $W$ is convex, i.e. we have $H^{1}(C,f^{*}W)=0$ for any genus-zero stable maps $C\to X$ and $\mathbf{c}$ is the equivariant Euler class $e_{\lambda}$. This corresponds to the choice of parameters
+    $W$ is convex, i.e. we have $H^{1}(C,f^{*}W)=0$ for any genus-zero stable map $C\to X$ and $\mathbf{c}$ is the equivariant Euler class $e_{\lambda}$. This corresponds to the choice of parameters
     $s_{k}=\begin{cases}\log\lambda&k=0;\\ (-1)^{k-1}(k-1)!\lambda^{-k}&k\geq 1.\end{cases}$
@@
 $\lim_{\lambda\to 0}\left\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\right\rangle_{0,n,d}^{X,(W,e_{\lambda})}=\sum_{i_{*}d^{\prime}=d}\left\langle i^{*}(\alpha_{1})\psi^{k_{1}},\dots,i^{*}(\alpha_{n})\psi^{k_{n}}\right\rangle_{0,n,d^{\prime}}^{Z}$
-where $i\colon Z\to X$ is the inclusion map. Moreover, the fundamental solution $M_{X,(W,e_{\lambda})}$ in the $(W,e_{\lambda})$-twisted theory satisfies (see [44, §2.1], [26, Proposition 2.4])
+where $i\colon Z\to X$ is the inclusion map. Moreover, the fundamental solution $M_{X,(W,e_{\lambda})}$ in the $(W,e_{\lambda})$-twisted theory satisfies (see [47, §2.1], [26, Proposition 2.4])
 $\lim_{\lambda\to 0}i^{*}M_{X,(W,e_{\lambda})}(\tau)\phi=M_{Z}(i^{*}\tau)i^{*}\phi\Bigr|_{H_{2}(Z,\mathbb{Z})\to H_{2}(X,\mathbb{Z})}$
@@
 $\exp(x)\sigma=\tau,\qquad x|_{Q=y=0}=0.$
-This can be shown easily by induction on powers of $Q$ and $y$ (see [23, Theorem 3.17] for a related result). Then we have
+This can be shown easily by induction on powers of $Q$ and $y$ (see [24, Theorem 3.17] for a related result). Then we have
 $\exp(x)f=\exp(x)\left(zM_{X}(\sigma)g\right)=zM_{X}(\tau)g^{\prime}$
@@
 Take an ample class $\omega$ of $X$, and consider the additive valuation $v\colon\mathbb{C}[\![Q,y]\!]\to\mathbb{R}\cup\{\infty\}$ given by $v(Q^{d})=\omega\cdot d$ and $v(y_{i})=i$.
-Suppose by induction that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that the equation (2.20) holds modulo terms with valuation $>v_{0}$ for some $v_{0}\in v(\mathbb{C}[\![Q,y]\!])$.
+Suppose by induction that there exist $w_{0},\dots,w_{s}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ such that the equation (2.20) holds modulo terms with valuation $v_{0}$ for some $v_{0}\in v(\mathbb{C}[\![Q,y]\!])$.
 Let $v_{1}>v_{0}$ be the next possible value of the valuation $v$.
 Write $\exp(\sum_{i=0}^{s}w_{i}z\nabla_{\tau^{i}})g^{\prime}=1-\sum_{i=0}^{s}\delta_{i}\phi_{i}$ for some $\delta_{i}\in\mathbb{C}[z][\![Q,\tau,y]\!]$ with $v(\delta_{i})\geq v_{1}$.
-Using the fact that $z\nabla_{\tau^{i}}1=\phi_{i}$, we have that $\exp\left(\sum_{i=0}^{s}(w_{i}+\delta_{i})z\nabla_{\tau^{i}}\right)g^{\prime}$ equals 1 modulo terms with valuation $>v_{1}$.
+Using the fact that $z\nabla_{\tau^{i}}1=\phi_{i}$, we have that $\exp\left(\sum_{i=0}^{s}(w_{i}+\delta_{i})z\nabla_{\tau^{i}}\right)g^{\prime}$ equals 1 modulo terms with valuation $v_{1}$.
 This completes the induction and the proof of the claim.
@@
 In this section, we work with the $\mathbb{C}$-basis $\{\phi_{i}\lambda^{k}\}$ and the associated $\mathbb{C}$-linear coordinates $\{\tau^{i,k}\}$ of $H^{*}_{S^{1}}(V)$ as in §2.3.
-The shift operator shall be defined over the ring $\mathbb{C}[\![\boldsymbol{\tau}]\!]$ of formal power series in $\{\tau^{i,k}\}$.
-Since $V$ is semiprojective and the fibrewise scalar $S^{1}$-action is seminegative [^5] On the other hand, the inverse $S^{1}$-action is not seminegative: this is the reason why $\mathbb{S}(\tau)^{-1}$ is not defined without localization.
+The shift operator shall be defined over the ring $\mathbb{C}[\![\boldsymbol{\tau}]\!]$ of formal power series in $\boldsymbol{\tau}=\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$.
+Since $V$ is semi-projective and the fibrewise scalar $S^{1}$-action is seminegative [^7] On the other hand, the inverse $S^{1}$-action is not seminegative: this is the reason why $\mathbb{S}(\tau)^{-1}$ is not defined without localization.
 (in the sense of [27, Definition 3.3]), the $S^{1}$-action defines the shift operator [27, Definition 3.9]
@@
 The third property of (2.21) is in [27, Theorem 3.14] and the second follows from the third together with the differential equations (2.7) for $M_{V}$ and the commutativity
-$[\partial_{\tau^{i,k}},\mathcal{S}]=[\xi Q\partial_{Q}+z^{-1}\xi,\mathcal{S}]=[z\partial_{z}+z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V},\mathcal{S}]=0.$
+$\partial_{\tau^{i,k}},\mathcal{S}]=[\xi Q\partial_{Q}+z^{-1}\xi,\mathcal{S}]=[z\partial_{z}+z^{-1}c_{1}^{S^{1}}(TV)+\mu_{V},\mathcal{S}]=0.$
 The shift operator is compatible with the pairing $P_{V}$ (2.8) in the sense that
@@
 The image of $\mathbb{S}(\tau)$ is a sub $D$-module since $\mathbb{S}(\tau)$ commutes with the connection.
-This fact is related to the quantum Serre [7, Corollary 9]: the above submodule can be identified with the quantum $D$-module of the $(V^{\vee},e_{\lambda})$-twisted theory [30, Theorems 2.11, 3.14].
+This fact is related to the quantum Serre [7, Corollary 9]: the above submodule can be identified with the quantum $D$-module of the $(V^{\vee},e_{\lambda})$-twisted theory [31, Theorems 2.11, 3.14].
 3. Proof of the mirror theorem
 In this section we prove Theorem 1.1 (and a more general version Theorem 3.4 below).
-Let $\mathbb{P}(V)\to B$ be the projective bundle associated with a vector bundle $V\to B$ of rank $r\geq 2$.
+Let $\pi\colon\mathbb{P}(V)\to B$ be the projective bundle associated with a vector bundle $V\to B$ of rank $r\geq 2$ such that $V^{\vee}$ is globally generated.
 Let $p=c_{1}(\mathcal{O}(1))$ be the first Chern class of the relative $\mathcal{O}(1)$ over $\mathbb{P}(V)$.
-This class defines the following splitting of $H_{2}(\mathbb{P}(V),\mathbb{Z})$:
+This class defines a splitting of the natural sequence $0\to\mathbb{Z}\to H_{2}(\mathbb{P}(V),\mathbb{Z})\xrightarrow{\pi_{*}}H_{2}(B,\mathbb{Z})\to 0$ as follows:
 (3.1) $H_{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}\oplus\operatorname{Ker}p\cong\mathbb{Z}\oplus H_{2}(B,\mathbb{Z}).$
 Notation 3.1.
-We write $Q$ for the Novikov variable for $B$, and $q$ for the Novikov variable for $\mathbb{P}(V)$ associated with the class of a line in a fibre.
-Via the above splitting (3.1), we identify $(q,Q)$ with the Novikov variables for $\mathbb{P}(V)$: a curve class $(k,d)\in\mathbb{Z}\oplus H_{2}(B,\mathbb{Z})\cong H_{2}(\mathbb{P}(V),\mathbb{Z})$ corresponds to the monomial $q^{k}Q^{d}$.
+We write $Q$ for the Novikov variable of the base $B$, and $q$ for the Novikov variable of $\mathbb{P}(V)$ associated with the class of a line in a fibre.
+Via the above splitting (3.1), we identify the pair $(q,Q)$ with the Novikov variables for $\mathbb{P}(V)$; more precisely, a curve class $d\in H_{2}(\mathbb{P}(V),\mathbb{Z})$ corresponds to the monomial $q^{p\cdot d}Q^{\pi_{*}d}$.
+The assumption that $V^{\vee}$ is globally generated implies that the class $p$ is nef; consequently the exponent of $q$ in this monomial is nonnegative whenever $d$ is an effective class.
 As before, we consider the fibrewise scalar $S^{1}$-action on $V$ and denote by $\lambda$ the $S^{1}$-equivariant parameter.
@@
 Since $V^{\vee}$ is globally generated, we have a surjection $\mathcal{O}^{\oplus N}\to V^{\vee}$ for some $N$. Dualizing, we get an exact sequence of vector bundles over $B$:
-(3.2) $0\to V\to\mathcal{O}^{\oplus N}\to\mathcal{Q}\to 0$
+(3.2) $0\to V\to\mathcal{O}^{\oplus N}\to\mathcal{Q}\to 0.$
 The inclusion $V\hookrightarrow\mathcal{O}^{\oplus N}$ induces an embedding $\mathbb{P}(V)\hookrightarrow B\times\mathbb{P}^{N-1}$.
@@
 Clearly, $\mathcal{Q}(1)$ is generated by global sections and therefore $\mathcal{Q}(1)$ is convex.
-Our starting point is the fact that Theorem 3.3 holds for the trivial bundle $B\times\mathbb{C}^{N}$: the $J$-function of $B\times\mathbb{P}^{N-1}$ at parameter $\tau+tp$ (with $\tau\in H^{*}(B)$) is given by
-$J_{B\times\mathbb{P}^{N-1}}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}J_{B}(\tau).$
+Our starting point is the fact that Theorem 3.4 holds for the trivial bundle $B\times\mathbb{C}^{N}\to B$: the $J$-function of $B\times\mathbb{P}^{N-1}$ at parameter $\tau+tp$ (with $\tau\in H^{*}(B)$) is given by
+(3.3) $J_{B\times\mathbb{P}^{N-1}}(\tau,t)=\sum_{k=0}^{\infty}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}J_{B}(\tau).$
 This can be viewed as a special case of the Elezi-Brown mirror theorem [3].
@@
 We shall compute a slice of the Givental cone of the $(\mathcal{Q}(1),e_{\lambda})$-twisted theory and take its non-equivariant limit.
-3.2. Informal proof
-We first give a slightly informal proof, ignoring the issue on adic convergence.
-A rigorous proof follows in the next section, but it is only technically complicated and we recommend the reader to begin with the discussion in this section.
-We introduce a modified QRR operator:
-(3.3) $\Delta^{\lambda}_{W}:=e^{\operatorname{rank}(W)(\lambda\log\lambda-\lambda)/z}\Delta_{(W,e_{\lambda}^{-1})}$
-where $\Delta_{(W,e_{\lambda}^{-1})}$ is the QRR operator (2.19) of the inverse Euler twist. This arises from the asymptotic expansion (Stirling’s approximation) of the $\Gamma$-function:
-$\Delta^{\lambda}_{W}\sim\prod_{\delta}\sqrt{\frac{z}{2\pi}}z^{\frac{\lambda+\delta}{z}}\Gamma\left(1+\frac{\lambda+\delta}{z}\right)$
-where $\delta$ ranges over the Chern roots of $W$.
-By the String Equation, the Givental cone is preserved by the action of $e^{g(z)/z}$, with $g(z)$ a nonnegative power series of $z$ with scalar coefficients.
-Hence we can use $\Delta_{W}^{\lambda}$ instead of $\Delta_{(W,e_{\lambda}^{-1})}$ in the QRR Theorem 2.5.
+3.2. Proof of Theorem 1.1
+We provide a rigorous treatment based on the $\lambda^{-1}$-adic topology, following the method of Remark 2.6.
+Throughout this section, $\tau$ is assumed to be a non-equivariant class in $H^{*}(V)=H^{*}(B)$ (i.e.
+independent of $\lambda$).
+A generalization to equivariant classes will be discussed in Theorem 3.4.
+We introduce a modified Quantum Riemann-Roch (QRR) operator:
+(3.4) $\begin{split}\Delta^{\lambda}_{W}&:=e^{\operatorname{rank}(W)(\lambda\log\lambda-\lambda)/z}\Delta_{(W,e_{\lambda}^{-1})}\\ &=\prod_{\delta}e^{\frac{(\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta)}{z}+\frac{1}{2}\log(\lambda+\delta)+\sum_{n\geq 2}\frac{B_{n}}{n(n-1)}\left(\frac{z}{\lambda+\delta}\right)^{n-1}}\end{split}$
+where $\Delta_{(W,e_{\lambda}^{-1})}$ is the QRR operator (2.19) of the inverse Euler twist, and $\delta$ ranges over the Chern roots of $W$.
+This arises from the asymptotic expansion (Stirling’s approximation) of the $\Gamma$-function:
+$\Delta^{\lambda}_{W}\sim\prod_{\delta}\sqrt{\frac{z}{2\pi}}z^{\frac{\lambda+\delta}{z}}\Gamma\left(1+\frac{\lambda+\delta}{z}\right).$
 The identity $\Gamma(1+x)=x\Gamma(x)$ implies that
 (3.5) $\Delta_{W}^{\lambda+kz}/\Delta_{W}^{\lambda}=\prod_{c=1}^{k}\prod_{\delta}(\lambda+\delta+cz).$
-The exact sequence (3.2) implies $\Delta_{V}^{\lambda}\Delta_{\mathcal{Q}}^{\lambda}=\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$.
-Since $\Delta_{\mathcal{O}^{\oplus N}}^{\lambda}$ preserves the Givental cone (again by the String Equation), we have $\Delta_{\mathcal{Q}}^{\lambda}\mathcal{L}_{B,(V,e_{\lambda}^{-1})}=\mathcal{L}_{B}$ by the QRR Theorem 2.5.
-Let $J_{V}^{\lambda}(\tau)$ be the equivariant $J$-function of $V$ so that $zJ_{V}^{\lambda}(\tau)\in\mathcal{L}_{B,(V,e_{\lambda}^{-1})}$.
-Then $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)$ lies in $\mathcal{L}_{B}$.
-As in Lemma 2.8, we expect to have a differential operator $F(\lambda)=F(\lambda;\tau,z\partial_{\tau},z)$ in $\tau$ such that
-(3.5) $\Delta_{\mathcal{Q}}^{\lambda}zJ_{V}^{\lambda}(\tau)=e^{F(\lambda)/z}zJ_{B}(\tau).$
-This is not justified by the argument in the proof of Lemma 2.8 due to the lack of an appropriate formal topology; $\Delta^{\lambda}_{\mathcal{Q}}$ contains both positive and negative powers of $\lambda$ and also $\log\lambda$.
-Let us proceed assuming that we have such an $F(\lambda)$.
-Consider the differential operator $\exp(F(\lambda+z\partial_{t})/z)$ in $t$ and $\tau$; here we expand $F(\lambda+z\partial_{t})$ in Taylor series of $z\partial_{t}$.
-By Lemma 2.7, $e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$ lies in $\mathcal{L}_{B\times\mathbb{P}^{N-1}}$.
-Therefore
-$I^{\lambda}(\tau,t):=(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$
-lies in the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$ by the QRR Theorem 2.5. We calculate
-$I^{\lambda}(\tau,t)$ $=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+p+kz)/z}J_{B}(\tau)$
-$=\sum_{k\geq 0}\frac{e^{tp/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}(\Delta_{\mathcal{Q}}^{\lambda+p})^{-1}\Delta^{\lambda+p+kz}_{\mathcal{Q}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.5)
-(3.6) $=\sum_{k\geq 0}e^{tp/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+p+\epsilon+cz)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)$ by (3.4)
-where $\epsilon$ ranges over the Chern roots of $\mathcal{Q}$ in the last line.
-As discussed in §2.5, $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ lies in the Givental cone $\mathcal{L}_{\mathbb{P}(V)}$ of $\mathbb{P}(V)$, where $i\colon\mathbb{P}(V)\hookrightarrow B\times\mathbb{P}^{N-1}$ is the inclusion.
-This gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1.
-Note that $i_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\to H_{2}(B\times\mathbb{P}^{N-1},\mathbb{Z})$ is an isomorphism whenever $r=\operatorname{rank}(V)\geq 2$.
-Remark 3.2.
-Another point that we glossed over in the above discussion is that the modified Riemann-Roch operator (3.3)
-(3.7) $\Delta^{\lambda}_{W}=\prod_{\delta}e^{\frac{(\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta)}{z}+\frac{1}{2}\log(\lambda+\delta)+\sum_{n\geq 2}\frac{B_{n}}{n(n-1)}\left(\frac{z}{\lambda+\delta}\right)^{n-1}}$
-arising from the Stirling approximation does not converge in an adic topology, because the exponent contains [^6] This is also the case with the original operator $\Delta_{(W,e_{\lambda}^{-1})}$, but this was not a problem because the coefficient of $z^{-1}$ in $\log(\Delta_{(W,e_{\lambda}^{-1})})$ is nilpotent.
-both positive and negative powers of $z$.
-This converges in the standard topology to power series in $z$ and $z^{-1}$ (infinite in both directions) with coefficients in analytic functions of $\lambda$ near $\lambda=e$, but we are not sure whether this fact would be useful.
-The only properties of $\Delta_{W}^{\lambda}$ we use in the rigorous argument below (e.g.
-(3.4)) can be interpreted in terms of the well-defined logarithm $\log(\Delta^{\lambda}_{W})\in z^{-1}H^{*}(B)\otimes\mathbb{C}[\lambda,\lambda^{-1},\log\lambda][\![z]\!]$ and the convergence issue does not exist.
-3.3. Rigorous proof
-We explain how to make the above argument rigorous, using the $\lambda^{-1}$-adic topology.
-We follow the strategy explained in Remark 2.6.
-Let $\tilde{e}_{\lambda}(\cdot)=\sum_{i\geq 0}\lambda^{-i}c_{i}(\cdot)$ be the modified equivariant Euler class and let $\widetilde{\Delta}_{W}^{\lambda}:=\Delta_{(W,\tilde{e}_{\lambda}^{-1})}$ denote the QRR operator (2.19) of the $(W,\tilde{e}_{\lambda}^{-1})$-twist.
+Ignoring issues of adic convergence (see Remark 3.3), we may regard the modified operator $\Delta_{W}^{\lambda}$ as playing the same role as the original $\Delta_{(W,e_{\lambda}^{-1})}$, because $e^{c/z}$ preserves the Givental cone for any constant $c$.
+Let $\tilde{e}_{\lambda}(\cdot)=\sum_{i\geq 0}\lambda^{-i}c_{i}(\cdot)$ be the variant of $e_{\lambda}$ introduced in Remark 2.6 and let $\widetilde{\Delta}_{W}^{\lambda}:=\Delta_{(W,\tilde{e}_{\lambda}^{-1})}$ denote the QRR operator (2.19) associated with the $(W,\tilde{e}_{\lambda}^{-1})$-twist.
 We have $\widetilde{\Delta}_{W}^{\lambda}=1+O(\lambda^{-1})$.
-We also set $\widetilde{J}_{V}^{\lambda}(\tau):=J_{V}^{\lambda}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$; this does not contain positive powers of $\lambda$ and gives the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory.
+We also set $\widetilde{J}_{V}^{\lambda}(\tau):=J_{V}^{\lambda}(\tau)|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$; this does not contain positive powers of $\lambda$ and gives the $J$-function of the $(V,\tilde{e}_{\lambda}^{-1})$-twisted theory by Remark 2.6.
 We have
-$\Delta_{W}^{\lambda}$ $=e^{(\operatorname{rank}(W)\theta(\lambda)+c_{1}(W)\log\lambda)/z}\widetilde{\Delta}_{W}^{\lambda}\quad\text{with $\theta(\lambda):=\lambda\log\lambda-\lambda+\frac{\log\lambda}{2}z$}$
-$J_{V}^{\lambda}(\tau)$ $=e^{(\log\lambda)(c_{1}(V)-z\partial_{c_{1}(V)})/z}\widetilde{J}_{V}^{\lambda}(\tau)$
-where $\partial_{c_{1}(V)}$ means the directional derivative along $c_{1}(V)$ in the $\tau$-space. The second equation follows from the Divisor Equation. Combining these equations, we have
-(3.8) $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)\quad\text{with $f(\lambda):=\operatorname{rank}(\mathcal{Q})\theta(\lambda)-(\log\lambda)z\partial_{c_{1}(V)}$.}$
-Since $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{\Delta}_{V}^{\lambda}=1$ (by the exact sequence (3.2)), $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)$ is a $\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5.
+$\Delta_{W}^{\lambda}=e^{(\operatorname{rank}(W)\theta(\lambda)+c_{1}(W)\log\lambda)/z}\widetilde{\Delta}_{W}^{\lambda}\quad\text{with $\theta(\lambda):=\lambda\log\lambda-\lambda+\frac{z}{2}\log\lambda$}.$
+Introduce a differential operator $f(\lambda)$ in $\tau$ by
+$f(\lambda):=\operatorname{rank}(\mathcal{Q})\theta(\lambda)-(\log\lambda)z\partial_{c_{1}(V)}$
+where $\partial_{c_{1}(V)}$ denotes the directional derivative along $c_{1}(V)$ in the $\tau$-space.
+We consider the differential operator $f(\lambda+z\partial_{t})$ in both $\tau$ and $t$, defined via the Taylor expansion in $z\partial_{t}$, and define its regularization $f_{\rm reg}(\lambda,z\partial_{t})$ by subtracting terms involving $\log\lambda$:
+$f_{\rm reg}(\lambda,z\partial_{t})$ $:=f(\lambda+z\partial_{t})-f(\lambda)-\operatorname{rank}(\mathcal{Q})(\log\lambda)z\partial_{t}$
+$=\operatorname{rank}(\mathcal{Q})(\theta(\lambda+z\partial_{t})-\theta(\lambda)-(\log\lambda)z\partial_{t})-\log(1+\lambda^{-1}z\partial_{t})z\partial_{c_{1}(V)}.$
+This belongs to $\mathbb{C}[z\partial_{t},z\partial_{c_{1}(V)}][\![\lambda^{-1}]\!]$ and satisfies $f_{\rm reg}(\lambda,z\partial_{t})|_{\lambda^{-1}=0}=0$.
+The exact sequence (3.2) shows $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{\Delta}_{V}^{\lambda}=1$.
+Thus $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}z\widetilde{J}_{V}^{\lambda}(\tau)$ is a $\mathbb{C}[\![Q,\tau,\lambda^{-1}]\!]$-valued point of $\mathcal{L}_{B}$ by the QRR Theorem 2.5.
 This is a miniversal slice, i.e.
@@
 (3.6) $\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)=e^{\widetilde{F}(\lambda)/z}J_{B}(\tau).$
-Mimicking the previous argument in §3.2, we consider applying the differential operator $e^{f(\lambda+z\partial_{t})/z}e^{\widetilde{F}(\lambda+z\partial_{t})/z}$ to $J_{B\times\mathbb{P}^{N-1}}(\tau,t)$; the first factor $e^{f(\lambda+z\partial_{t})/z}$ was added in view of the difference (3.8) between $\Delta^{\lambda}_{\mathcal{Q}}J_{V}^{\lambda}$ and $\widetilde{\Delta}^{\lambda}_{\mathcal{Q}}\widetilde{J}_{V}^{\lambda}$.
-To work with $\lambda^{-1}$-power series, we subtract from $f(\lambda+z\partial_{t})$ the first few terms of its Taylor expansion in $z\partial_{t}$ and define:
-$f_{\rm reg}(\lambda,z\partial_{t}):=f(\lambda+z\partial_{t})-f(\lambda)-\operatorname{rank}(\mathcal{Q})(\log\lambda)z\partial_{t}.$
-This belongs to $\mathbb{C}[z\partial_{t},z\partial_{c_{1}(V)}][\![\lambda^{-1}]\!]$ and $f_{\rm reg}(\lambda,z\partial_{t})|_{\lambda^{-1}=0}=0$.
-Therefore, Lemma 2.7 and the QRR Theorem 2.5 imply that
+Lemma 2.7 and the QRR Theorem 2.5 imply that
 $\widetilde{I}^{\lambda}(\tau,t):=(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{f_{\rm reg}(\lambda,z\partial_{t})/z}e^{\widetilde{F}(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$
 is a $\mathbb{C}[\![Q,\tau,t,\lambda^{-1}]\!]$-valued point of the $(\mathcal{Q}(1),\tilde{e}_{\lambda})$-twisted Givental cone of $B\times\mathbb{P}^{N-1}$.
-By calculations similar to, but more cumbersome than, the previous section §3.2, we arrive at the following result (cf.
-(3.6)).
-$\widetilde{I}^{\lambda}(\tau,t)=\sum_{k\geq 0}e^{pt/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}\left(1+\frac{p+\epsilon+cz}{\lambda}\right)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$
-where $\prod_{\epsilon}$ is the product over the Chern roots $\epsilon$ of $\mathcal{Q}$. For this calculation, we used (3.9) and the following identities:
-$\widetilde{\Delta}^{\lambda}_{\mathcal{Q}(1)}$ $=e^{(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+p)-\theta(\lambda)-p\log\lambda)+c_{1}(\mathcal{Q})\log(1+\frac{p}{\lambda}))/z}\widetilde{\Delta}^{\lambda+p}_{\mathcal{Q}},$
-$\widetilde{\Delta}_{\mathcal{Q}}^{\lambda+kz}/\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}$ $=e^{-\left(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+kz)-\theta(\lambda))+c_{1}(\mathcal{Q})\log(1+\frac{kz}{\lambda})\right)/z}\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+\epsilon+cz),$
-$\ J_{V}^{\lambda+x}(\tau)\Bigr{|}_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ $=e^{-c_{1}(\mathcal{Q})\log(1+\frac{x}{\lambda})/z}e^{-\log(1+\frac{x}{\lambda})\partial_{c_{1}(V)}}\widetilde{J}_{V}^{\lambda+x}(\tau).$
-The first identity follows from $\Delta^{\lambda}_{\mathcal{Q}(1)}=\Delta^{\lambda+p}_{\mathcal{Q}}$, the second from (3.4), and the third from the Divisor Equation.
-Finally we make the change of variables $Q^{d}\to Q^{d}\lambda^{c_{1}(\mathcal{Q})\cdot d}=Q^{d}\lambda^{-c_{1}(V)\cdot d}$, $q^{k}\to q^{k}\lambda^{\operatorname{rank}(\mathcal{Q})k}$ for $\widetilde{I}^{\lambda}$ and obtain a point $I^{\lambda}$ of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone.
-This gives the same result as (3.6).
-The non-equivariant limit $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ exists and gives the $I$-function $I_{\mathbb{P}(V)}$ lying on the Givental cone of $\mathbb{P}(V)$.
+We will calculate this element $\widetilde{I}^{\lambda}$.
+First observe the following identities:
+(3.7) $\begin{split}&\widetilde{\Delta}^{\lambda}_{\mathcal{Q}(1)}=e^{(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+p)-\theta(\lambda)-p\log\lambda)+c_{1}(\mathcal{Q})\log(1+\frac{p}{\lambda}))/z}\widetilde{\Delta}^{\lambda+p}_{\mathcal{Q}},\\ &\widetilde{\Delta}_{\mathcal{Q}}^{\lambda+kz}/\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}=e^{-\left(\operatorname{rank}(\mathcal{Q})(\theta(\lambda+kz)-\theta(\lambda))+c_{1}(\mathcal{Q})\log(1+\frac{kz}{\lambda})\right)/z}\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+\epsilon+cz),\\ \ &J_{V}^{\lambda+x}(\tau)\Bigr|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}=e^{-c_{1}(\mathcal{Q})\log(1+\frac{x}{\lambda})/z}e^{-\log(1+\frac{x}{\lambda})\partial_{c_{1}(V)}}\widetilde{J}_{V}^{\lambda+x}(\tau)\end{split}$
+where $\prod_{\epsilon}$ is the product over the Chern roots of $\mathcal{Q}$.
+The first identity follows from $\Delta_{\mathcal{Q}(1)}^{\lambda}=\Delta_{\mathcal{Q}}^{\lambda+p}$, the second from (3.5), and the thrid from the Divisor Equation.
+We have
+$\widetilde{I}^{\lambda}(\tau,t)$ $=(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}\sum_{k\geq 0}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}e^{f_{\rm reg}(\lambda,p+kz)/z}e^{\widetilde{F}(\lambda+p+kz)/z}J_{B}(\tau)$ by (3.3)
+$=\sum_{k\geq 0}\frac{e^{pt/z}q^{k}e^{kt}}{\prod_{c=1}^{k}(p+cz)^{N}}e^{f_{\rm reg}(\lambda,p+kz)/z}(\widetilde{\Delta}_{\mathcal{Q}(1)}^{\lambda})^{-1}\widetilde{\Delta}^{\lambda+p+kz}_{\mathcal{Q}}\widetilde{J}^{\lambda+p+kz}_{V}(\tau)$ by (3.6)
+$=\sum_{k\geq 0}e^{pt/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}\left(1+\frac{p+\epsilon+cz}{\lambda}\right)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau)\Bigr|_{Q^{d}\to Q^{d}\lambda^{c_{1}(V)\cdot d}}$ by (3.7).
+By the change of variables $Q^{d}\to Q^{d}\lambda^{c_{1}(\mathcal{Q})\cdot d}=Q^{d}\lambda^{-c_{1}(V)\cdot d}$, $q^{k}\to q^{k}\lambda^{\operatorname{rank}(\mathcal{Q})k}$ for $\widetilde{I}^{\lambda}$, we obtain a point $I^{\lambda}$ of the $(\mathcal{Q}(1),e_{\lambda})$-twisted Givental cone by Remark 2.6.
+It is given by
+$I^{\lambda}(\tau,t)=\sum_{k\geq 0}e^{tp/z}q^{k}e^{kt}\frac{\prod_{c=1}^{k}\prod_{\epsilon}(\lambda+p+\epsilon+cz)}{\prod_{c=1}^{k}(p+cz)^{N}}J_{V}^{\lambda+p+kz}(\tau).$
+The non-equivariant limit $\lim_{\lambda\to 0}i^{*}I^{\lambda}$ exists and gives the $I$-function $I_{\mathbb{P}(V)}$ in Theorem 1.1.
+As discussed in §2.5, $I_{\mathbb{P}(V)}$ lies on the Givental cone of $\mathbb{P}(V)$ .
 The proof of Theorem 1.1 is now complete.
+Remark 3.2.
+The differential operator $f(\lambda)$ appearing in the proof above is chosen to satisfy the identity $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}\widetilde{\Delta}_{\mathcal{Q}}^{\lambda}\widetilde{J}_{V}^{\lambda}(\tau)$, which leads to
+$\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{f(\lambda)/z}e^{\widetilde{F}(\lambda)/z}J_{B}(\tau).$
+Conversely, if one chooses a differential operator $F(\lambda)=F(\lambda;\tau,z\partial_{\tau},z)$ in $\tau$ such that $\Delta_{\mathcal{Q}}^{\lambda}J_{V}^{\lambda}(\tau)=e^{F(\lambda)/z}J_{B}(\tau)$, then a straightforward calculation using (3.5) shows that the element $(\Delta_{\mathcal{Q}(1)}^{\lambda})^{-1}e^{F(\lambda+z\partial_{t})/z}J_{B\times\mathbb{P}^{N-1}}(\tau,t)$ yields the point $I^{\lambda}(\tau,t)$ defined above.
+This point is expected to lie on the $(\mathcal{Q}(1),e_{\lambda})$-twisted cone of $B\times\mathbb{P}^{N-1}$, and the preceding construction makes this heuristic rigorous via the $\lambda^{-1}$-adic topology.
+Remark 3.3.
+We note that the modified QRR operator (3.4) arising from the Stirling approximation does not converge in an adic topology, because the exponent contains [^8] This also holds for the original operator $\Delta_{(W,e_{\lambda}^{-1})}$; however this is not an issue because the coefficient of $z^{-1}$ in $\log(\Delta_{(W,e_{\lambda}^{-1})})$ is nilpotent.
+both positive and negative powers of $z$.
+This converges in the standard topology to power series in $z$ and $z^{-1}$ (infinite in both directions) with coefficients in analytic functions of $\lambda$ near $\lambda=e$, but we are not sure whether this fact would be useful.
+The properties of $\Delta_{W}^{\lambda}$ we used in the proof (e.g.
+(3.5)) can be interpreted in terms of the well-defined logarithm $\log(\Delta^{\lambda}_{W})\in z^{-1}H^{*}(B)\otimes\mathbb{C}[\lambda,\lambda^{-1},\log\lambda][\![z]\!]$ and the convergence issue does not exist.
 3.3. A generalization
@@
 4.1. Quantum $D$-modules
-We fix notation and definitions concering the quantum $D$-modules of $\mathbb{P}(V)$ and $V$.
+We fix notation and definitions concerning the quantum $D$-modules of $\mathbb{P}(V)$ and $V$.
 4.1.1. Quantum $D$-module of $\mathbb{P}(V)$
-Let $\hat{\tau}\in H^{*}(\mathbb{P}(V))$ denote the parameter of the quantum cohomology of $\mathbb{P}(V)$.
+Let $\hat{\tau}\in H^{*}(\mathbb{P}(V))$ denote the bulk-deformation parameter of the quantum cohomology of $\mathbb{P}(V)$.
 Following Notation 3.1, we write $(q,Q)$ for the Novikov variables of $\mathbb{P}(V)$, which are dual to the decomposition $H^{2}(\mathbb{P}(V),\mathbb{Z})=\mathbb{Z}p\oplus H^{2}(B,\mathbb{Z})$.
@@
 $P_{\mathbb{P}(V)}\colon\operatorname{QDM}(\mathbb{P}(V))\times\operatorname{QDM}(\mathbb{P}(V))\to\mathbb{C}[z,q][\![Q,\hat{\tau}$]\!]
-induced by the Poincaré pairing of $\mathbb{P}(V)$.
-We note that $\nabla_{q\partial_{q}}=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}})$.
-We also note that we defined $\operatorname{QDM}(\mathbb{P}(V))$ over $\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]$ instead of over $\mathbb{C}[z][\![q,Q,\hat{\tau}]\!]$: it is possible because of the homogeneity in Gromov-Witten theory and the fact that both $z$ and $q$ have positive degrees.
+[…]
 4.1.2. Quantum $D$-module of $V$
-Let $\tau\in H^{*}_{S^{1}}(V)$ denote the parameter of the equivariant quantum cohomology of $V$.
+Let $\tau\in H^{*}_{S^{1}}(V)$ denote the bulk-deformation parameter of the equivariant quantum cohomology of $V$.
 Let $\{\phi_{i}\}_{i=0}^{s}$ be a homogeneous basis of $H^{*}(B)$ as before; here we assume that $\phi_{0}=1$.
 Then $\{\phi_{i}\lambda^{k}\}_{0\leq i\leq s,k\geq 0}$ gives a $\mathbb{C}$-basis of $H^{*}_{S^{1}}(V)$.
-We write $\tau=\sum_{i=0}^{s}\sum_{k=0}^{\infty}\tau^{i,k}\phi_{i}\lambda^{k}$ with $\tau^{i,k}$ being $\mathbb{C}$-valued coordinates.
+We write $\tau=\sum_{i=0}^{s}\sum_{k=0}^{\infty}\tau^{i,k}\phi_{i}\lambda^{k}$, where the $\tau^{i,k}$ are $\mathbb{C}$-valued coordinates, and we use $\boldsymbol{\tau}$ to denote the infinite set $\{\tau^{i,k}\}_{0\leq i\leq s,k\geq 0}$ of coordinates.
 The equivariant quantum $D$-module of $V$ is the module
@@
 4.2. Fourier transformation for quantum $D$-modules
-Recall the Fourier transformation (1.2) on the level of the Givental spaces from the introduction.
+Recall the Fourier transformation (1.3) on the level of the Givental spaces from the introduction:
 $\widehat{\phantom{A}}\colon$ $\mathcal{H}_{V,\rm pol}\to\mathcal{H}_{\mathbb{P}(V),\rm pol}$
@@
 where $\mathcal{H}_{V,\rm pol}$, $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ are the polynomial Givental spaces given in (2.17), (2.15).
-(Note that $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ is defined over the Novikov ring $\mathbb{C}[\![q,Q]\!]$.)
+(We define $\mathcal{H}_{\mathbb{P}(V),\rm pol}$ over the completed Novikov ring $\mathbb{C}[\![q,Q]\!]$.)
 Example 4.1 (cf. (1.5)).
@@
 Let $V\to B$ be as in Theorem 1.1.
-There exist a mirror map $\hat{\tau}=\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[q][\![Q,\boldsymbol{\tau}]\!]$ and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules which satisfy the following properties:
+There exist a mirror map $\hat{\tau}=\hat{\tau}(\tau)\in H^{*}(\mathbb{P}(V))[q][\![Q,\boldsymbol{\tau}]\!]$ and an isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\to\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules, such that by defining the pull-back connection $\nabla^{\prime}=\hat{\tau}^{*}\nabla$ via the standard formulae
+$\nabla^{\prime}_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\partial_{\tau^{i,k}}\hat{\tau})\star_{\hat{\tau}(\tau)},$
+$\nabla^{\prime}_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}(\tau)}+(q\partial_{q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
+$\nabla^{\prime}_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\hat{\tau}(\tau)}+(\xi Q\partial_{Q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
+$\nabla^{\prime}_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{\mathbb{P}(V)}\star_{\hat{\tau}(\tau)})+\mu_{\mathbb{P}(V)},$
+the following properties hold:
   • (1)
@@
   • (6)
-    $\hat{\tau}(\tau)$ is homogeneous of degree $2$ and $\operatorname{FT}$ is homogeneous of degree 00 with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(\mathbb{P}(V))$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2c_{1}(TV)\cdot d=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
+    $\hat{\tau}(\tau)$ is homogeneous of degree $2$ and $\operatorname{FT}$ is homogeneous of degree 0 with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(\mathbb{P}(V))$ and the following degrees of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2c_{1}(TV)\cdot d=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
   • (7)
@@
   • (8)
-    $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $10\leq k\leq r-$ 1;
+    $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $0\leq k\leq r-1$;
   • (9)
     $(M_{V}(\tau)s)^{^}=M_{\mathbb{P}(V)}(\hat{\tau}(\tau))\operatorname{FT}(s)$, i.e. the following diagram commutes:
-    ${\operatorname{QDM}_{S^{1}}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{FT}\quad}$$\scriptstyle{M_{V}(\tau)}$${\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{\mathbb{P}(V)}(\hat{\tau}(\tau))}$${\mathcal{H}_{V,\rm pol}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{\phantom{aa}}}$${\mathcal{H}_{\mathbb{P}(V),\rm pol},}$
-where we set $\nabla^{\prime}=\hat{\tau}^{*}\nabla$.
-Remark 4.3.
-The pull-back $\nabla^{\prime}=\hat{\tau}^{*}\nabla$ is defined in the standard way as follows:
-$\nabla^{\prime}_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\partial_{\tau^{i,k}}\hat{\tau})\star_{\hat{\tau}(\tau)},$
-$\nabla^{\prime}_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(p\star_{\hat{\tau}(\tau)}+(q\partial_{q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
-$\nabla^{\prime}_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\hat{\tau}(\tau)}+(\xi Q\partial_{Q}\hat{\tau})\star_{\hat{\tau}(\tau)}),$
-$\nabla^{\prime}_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{\mathbb{P}(V)}\star_{\hat{\tau}(\tau)})+\mu_{\mathbb{P}(V)}.$
+    ${\operatorname{QDM}_{S^{1}}(V)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{FT}\quad}$$\scriptstyle{M_{V}(\tau)}$${\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M_{\mathbb{P}(V)}(\hat{\tau}(\tau))}$${\mathcal{H}_{V,\rm pol}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}^\scriptstyle{\widehat{\phantom{aa}}}$${\mathcal{H}_{\mathbb{P}(V),\rm pol}.}$
 Proof of Theorem 4.2.
@@
 (4.2) $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,\big|_{Q=q=0}=e^{\kappa(\tau)/z}\phi_{i}p^{k}$
-and these vectors with $0\leq i\leq s$, $10\leq k\leq r-$ 1 form a basis of $H^{*}(\mathbb{P}(V))$.
+and these vectors with $0\leq i\leq s$, $0\leq k\leq r-1$ form a basis of $H^{*}(\mathbb{P}(V))$.
 The mirror map $\hat{\tau}(\tau)$ is then determined by $M_{\mathbb{P}(V)}(\hat{\tau}(\tau))1=1+z^{-1}\hat{\tau}(\tau)+O(z^{-2})$.
@@
 It is easy to see that its Fourier transform $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,$ is also homogeneous of the same degree.
-This implies the homogeneity of the Birkhoff factors $M_{\mathbb{P}}(V)(\hat{\tau}(\tau))$, $\Upsilon(\tau)$, $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ and hence of the mirror map $\hat{\tau}(\tau)$.
+This implies the homogeneity of the Birkhoff factors $M_{\mathbb{P}}(V)(\hat{\tau}(\tau))$, $z\nabla^{\prime}_{\tau^{i,k}}\Upsilon(\tau)$ and hence of $\Upsilon(\tau)=\nabla^{\prime}_{\tau^{0,0}}\Upsilon(\tau)$ and the mirror map $\hat{\tau}(\tau)$.
 The homogeneity implies that $\hat{\tau}(\tau)$ and $\Upsilon(\tau)$ depend polynomially on $q$ (for each fixed power of $Q$ and $\tau$), i.e.
@@
 $(M_{V}(\tau)\phi_{i}\lambda^{k})^{^}\,\big|_{Q=\tau=0}=\sum_{l=0}^{\infty}\frac{\phi_{i}(p+lz)^{k}}{\prod_{c=1}^{l}e_{p+cz}(V)}q^{l}=\begin{cases}\phi_{i}p^{k}+O(z^{-1})&k\leq r-1\\ \phi_{i}p^{r}+q\phi_{i}+O(z^{-1})&k=r\end{cases}$
-which shows that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $10\leq k\leq r-$ 1 and, since the multiplication by $\lambda$ corresponds to $z\nabla^{\prime}_{q\partial_{q}}$ under $\operatorname{FT}$, that
+which shows that $\operatorname{FT}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=\phi_{i}p^{k}$ for $0\leq k\leq r-1$ and, since the multiplication by $\lambda$ corresponds to $z\nabla^{\prime}_{q\partial_{q}}$ under $\operatorname{FT}$, that
 $z\nabla^{\prime}_{q\partial_{q}}(\phi_{i}p^{k})\big|_{Q=\tau=0}=\begin{cases}\phi_{i}p^{k+1}&0\leq k\leq r-2;\\ \phi_{i}p^{r}+q\phi_{i}&k=r-1.\end{cases}$
-Hence for $k=rn+m$ with $10\leq m\leq r-$ 1, we have
+Hence for $k=rn+m$ with $0\leq m\leq r-1$, we have
 $\operatorname{FT}(\phi_{i}\lambda^{k})\big|_{Q=\tau=0}$ $=(z\nabla^{\prime}_{q\partial_{q}})^{k}\phi_{i}\big|_{Q=\tau=0}$
@@
 $=\sum_{k,l\in\mathbb{Z}}q^{k+l}P_{\mathbb{P}(V)}(\kappa(\mathcal{S}^{-k}f),\kappa(\mathcal{S}^{-l}g))$ by Lemma 4.4 below
-$=P_{\mathbb{P}(V)}(\hat{f},\hat{g})$
+[…]
 $=P_{\mathbb{P}(V)}(M_{\mathbb{P}(V)}(\hat{\tau})\operatorname{FT}(x),M_{\mathbb{P}(V)}(\hat{\tau})\operatorname{FT}(y))$ by Theorem 4.2(9)
@@
   • (3)
-    $\pi_{*}(\kappa(\lambda^{i}))=0$ for $20\leq i\leq r-$ 2 and $\pi_{*}(\kappa(\lambda^{r-1}))=1$.
+    $\pi_{*}(\kappa(\lambda^{i}))=0$ for $0\leq i\leq r-2$ and $\pi_{*}(\kappa(\lambda^{r-1}))=1$.
 It is easy to check that the right-hand side satisfies the same properties. ∎
@@
 5.1. Statement
-We state a more precise version of Theorem 1.7.
-Let $\operatorname{QDM}(\mathbb{P}(V))$ be the quantum $D$-module of $\mathbb{P}(V)$ as in §4.1.1.
+We now present a more precise version of Theorem 1.7.
 Let $\{\phi_{i}\}_{i=0}^{s}$ be a basis of $H^{*}(B)$ with $\phi_{0}=1$ as before.
-Then $\{\phi_{i}p^{k}\}_{0\leq i\leq s,0\leq k\leq r-1}$ form a basis of $H^{*}(\mathbb{P}(V))$; we write $\hat{\tau}=\sum_{i=0}^{s}\sum_{k=0}^{r-1}\hat{\tau}^{i,k}\phi_{i}p^{k}$ for the parameter in $H^{*}(\mathbb{P}(V))$.
+Then $\{\phi_{i}p^{k}\}_{0\leq i\leq s,0\leq k\leq r-1}$ form a basis of $H^{*}(\mathbb{P}(V))$.
+We denote by $\sigma=\sum_{i=0}^{s}\sigma^{i}\phi_{i}\in H^{*}(B)$ the bulk parameter for $B$ and by $\hat{\tau}=\sum_{i=0}^{s}\sum_{k=0}^{r-1}\hat{\tau}^{i,k}\phi_{i}p^{k}\in H^{*}(\mathbb{P}(V))$ the bulk parameter for $\mathbb{P}(V)$.
 Set
 (5.1) $r^{\prime}:=\begin{cases}r&\text{when $r-1$ is even;}\\ 2r&\text{when $r-1$ is odd}.\end{cases}$
-We introduce the localized quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ as
-$\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}=\operatorname{QDM}(\mathbb{P}(V))\otimes_{\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!].$
-It is equipped with the connection $\nabla$ and the pairing $P_{\mathbb{P}(V)}$ as before.
+As in Notation 3.1, we use the pair $(q,Q)$ to denote the Novikov variable of $\mathbb{P}(V)$.
+Recall from §4.1.1 that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is defined over the ring $\mathbb{C}[z,q][\![Q]\!]$.
+Its localized version is defined as
+$\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}:=\operatorname{QDM}(\mathbb{P}(V))\otimes_{\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!],
+equipped with the induced connection $\nabla$ and the pairing $P_{\mathbb{P}(V)}$.
+In this section, we introduce an additional copy $Q_{B}$ of the Novikov variable for $B$ and define the quantum $D$-module of $B$ to be the module
+$\operatorname{QDM}(B)=H^{*}(B)[z][\![Q_{B},\sigma]\!]$
+equipped with the quantum connection $\nabla_{\sigma^{i}},\nabla_{\xi Q_{B}\partial_{Q_{B}}},\nabla_{z\partial_{z}}$ (with $\xi\in H^{2}(B)$) and the $z$-sesquilinear pairing $P_{B}$ (see §2.2).
+We relate the Novikov variable $Q_{B}$ of $B$ to the variables $(q,Q)$ of $\mathbb{P}(V)$ via the following formula:
+(5.2) $Q_{B}^{d}=q^{-c_{1}(V)\cdot d/r}Q^{d}$
+for $d\in H_{2}(B,\mathbb{Z})$, which is consistent with the degrees $\deg Q_{B}^{d}=2c_{1}(TB)\cdot d$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$.
+The formula (5.2) defines an embedding of the coefficient ring $\mathbb{C}[z][\![Q_{B}]\!]$ for $\operatorname{QDM}(B)$ into the ring $\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q]\!]\subset\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$.
+We introduce the following base changes of $\operatorname{QDM}(B)$:
+(5.3) $\begin{split}\operatorname{QDM}(B)_{\rm ext}&:=\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![Q_{B},\sigma]\!]}\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,\sigma]\!],\\ \operatorname{QDM}(B)_{\rm ext,loc}&:=\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![Q_{B},\sigma]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\sigma]\!].\end{split}$
+These extended modules are equipped with the quantum connection
+$\nabla_{\sigma^{i}}$ $=\nabla_{\sigma^{i}}\otimes 1+1\otimes\partial_{\sigma^{i}}$ $=\partial_{\sigma^{i}}+z^{-1}(\phi_{i}\star_{\sigma})$
+$\nabla_{q\partial_{q}}$ $=\nabla_{-(c_{1}(V)/r)Q_{B}\partial_{Q_{B}}}\otimes 1+1\otimes q\partial_{q}$ $=q\partial_{q}-z^{-1}((c_{1}(V)/r)\star_{\sigma})$
+$\nabla_{\xi Q\partial_{Q}}$ $=\nabla_{\xi Q_{B}\partial_{Q_{B}}}\otimes 1+1\otimes\xi Q\partial_{Q}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\sigma})$
+$\nabla_{z\partial_{z}}$ $=\nabla_{z\partial_{z}}\otimes 1+1\otimes z\partial_{z}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\sigma})+\mu_{B}$
+where $\xi\in H^{2}(B)$, and the $z$-sesquilinear pairing $P_{B}$. These structures are induced from those for $\operatorname{QDM}(B)$ via pull-back along the map $(Q,q)\mapsto Q_{B}$ given by (5.2).
 Theorem 5.1.
-There exist maps $H^{*}(\mathbb{P}(V))\to H^{*}(B)$, $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ with $\varsigma_{j}(\hat{\tau})\in-c_{1}(V)\log\lambda_{j}+H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$, $j=0,\dots,r-1$ with $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$ and an isomorphism
-$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
-such that
+[…]
+$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
+where $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ is the module $H^{*}(B)\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!] equipped with the pull-back quantum connection $\varsigma_{j}^{*}\nabla$
+$\varsigma_{j}^{*}\nabla_{\hat{\tau}^{i,k}}$ $=\partial_{\hat{\tau}^{i,k}}+z^{-1}(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})\star_{\varsigma_{j}(\hat{\tau})}$
+[…]
+[…]
+$\varsigma_{j}^{*}\nabla_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\varsigma_{j}(\hat{\tau})})+\mu_{B}$
+and the pairing $P_{B}$ as in (2.8), such that the following holds with $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$:
   • (1)
@@
   • (3)
-    $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ is homogeneous of degree $2$ and $\Phi$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}(\mathbb{P}(V))$, $H^{*}(B)$ and the following degree of the variables: $\deg\hat{\tau}^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
+[…]
   • (4)
-    $\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}-c_{1}(V)\log\lambda_{j}+O(q^{-1/r})$;
+    $\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+O(q^{-1/r})$;
   • (5)
-    the Jacobian matrix of $\varsigma(\hat{\tau})=\bigoplus_{j=0}^{r-1}\varsigma_{j}(\hat{\tau})$ along $Q=\hat{\tau}=0$ is $(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})|_{Q=\hat{\tau}=0}=\phi_{i}\lambda_{j}^{k}(1+O(q^{-1/r}))$ and is invertible over $\mathbb{C}(\!(q^{-1/r})\!)$;
+[…]
   • (6)
-    $\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}(\phi_{i}+O(q^{-1/r}))$
-where $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ is the module
-$\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}=H^{*}(B)\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$
-equipped with the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection of $B$ and the pairing $P_{B}$ (2.8).
+    $\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}(\phi_{i}+O(q^{-1/r}))$.
 Remark 5.2.
-Although the maps $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ contain $\log q$, the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection is well-defined by the Divisor Equation.
-We can eliminate $\log q$ from the statement of the above theorem by considering a coordinate change for the Novikov variables.
-Consider the map
-$\varsigma^{\rm reg}\colon(\hat{\tau},q,Q)\longmapsto(\varsigma,Q_{B})=\left(\{\varsigma_{j}(\hat{\tau})+c_{1}(V)(\log q)/r\}_{j=0}^{r-1},q^{-c_{1}(V)/r}Q\right)$
-where $\varsigma$ is a parameter of $\bigoplus_{j=0}^{r-1}H^{*}(B)$ and $Q_{B}$ is the Novikov variable for $B$.
-The map $\varsigma^{\rm reg}$ does not contain $\log q$.
-Then the Divisor Equation shows that $\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\nabla$ is the same as $(\varsigma^{\rm reg})^{*}(\nabla^{\oplus r})$.
-The identification $Q_{B}^{d}=q^{-c_{1}(V)\cdot d/r}Q^{d}$ of the Novikov variables is compatible with their degrees.
+The identification (5.2) corresponds to a splitting of $\pi_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})$ over $\mathbb{Q}$ defined by the kernel of $c_{1}(T_{\rm vert}\mathbb{P}(V))=rp+\pi^{*}c_{1}(V)$.
+Here $T_{\rm vert}\mathbb{P}(V)=\operatorname{Ker}(\pi_{*}\colon T\mathbb{P}(V)\to\pi^{*}TB)$ denotes the vertical tangent bundle.
+This splitting differs from the one in (3.1) defined by the kernel of $p$; however, it is intrinsic to the geometry of the projective bundle $\mathbb{P}(V)\to B$ and is independent of the choice of the vector bundle $V$ (up to tensoring with a line bundle).
 Remark 5.3.
-As suggested in [28, 24], we expect that the (formal) decomposition of quantum cohomology $D$-modules occurs over formal power series in $z$.
-We defined the localized quantum $D$-modules over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$, but more precisely we can define them over the ‘homogeneous’ completion $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ consisting of finite sums of homogeneous elements in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$.
-We observe that $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ is contained in the ring $\mathbb{C}[q^{-1/r^{\prime}},q^{1/r^{\prime}}][\![z]\!]$ of formal power series in $z$.
-If we have convergence of quantum cohomology with respect to the parameters $q$, $Q$, $\hat{\tau}$, we expect that the decomposition $\Phi$ cannot be analytified in $z$ but should be defined over $\mathcal{O}[\![z]\!]$, with $\mathcal{O}$ the sheaf of analytic functions in $q,Q,\hat{\tau}$.
-As the picture in [28, 24] suggests, this formal (in $z$) decomposition should be lifted to an analytic decomposition over an angular sector on the $z$-plane, and we expect that the analytic decomposition can be related (at certain values of the parameters) to the semi-orthogonal decomposition [43] of derived categories of coherent sheaves
-$D^{b}(\mathbb{P}(V))=\langle D^{b}(B)_{0},\dots,D^{b}(B)_{r-1}\rangle$
-via the $\widehat{\Gamma}$-integral structure, where $D^{b}(B)_{i}=D^{b}(B)\otimes\mathcal{O}(i)$.
+We considered the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ and its decomposition over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$.
+More precisely, these can be defined over the ‘homogeneous’ completion $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$, which consists of finite sums of homogeneous elements in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$.
+Because $z$ and $q$ both have positive degrees, $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ is contained in the ring $\mathbb{C}[q^{-1/r^{\prime}},q^{1/r^{\prime}}][\![z]\!]$ of formal power series in $z$.
+This is relevant to Remark 1.11.
 5.2. Idea of the proof
@@
 Therefore its Fourier transform
-[…]
+$s\longmapsto\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)s\,d\lambda$
 should give a solution for $\operatorname{QDM}(\mathbb{P}(V))$ through the Fourier isomorphism $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\cong\operatorname{QDM}_{S^{1}}(V)$ in Theorem 4.2.
@@
 the method of steepest descent).
-The QRR theorem tells us that this lies in the Givental cone of $B$ and induces projections $\operatorname{QDM}(\mathbb{P}(V))\to\varsigma_{j}^{*}\operatorname{QDM}(B)$, $j=0,\dots,r-1$, where the $r$ choices result from the choice of critical points we use for the stationary phase method.
+The QRR theorem tells us that this lies in the Givental cone of $B$ and induces projections $\operatorname{QDM}(\mathbb{P}(V))\to\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$, $j=0,\dots,r-1$, where the $r$ choices result from the choice of critical points we use for the stationary phase method.
 Remark 5.4.
 The QRR operator $\Delta_{V}^{\lambda}$ induces an isomorphism between the quantum $D$-modules of $V$ and $B$ after localizing $\lambda$.
-The isomorphism is given by a gauge transformation $R(\tau)\in\lambda^{-r/2}\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[z](\!(\lambda^{-1})\!)[\![Q,\tau]\!]$ from $\operatorname{QDM}_{S^{1}}(V)$ to $\sigma^{*}\operatorname{QDM}(B)$ satisfying
+The isomorphism is given by a gauge transformation $R(\tau)\in\lambda^{-r/2}\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[z](\!(\lambda^{-1})\!)[\![Q,\tau]\!]$ from $\operatorname{QDM}_{S^{1}}(V)$ to $\sigma^{*}(\operatorname{QDM}(B)|_{Q_{B}\to Q})$ satisfying
 $M_{B}(\sigma(\tau))R(\tau)=(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$
@@
 The shift operator acts on $\operatorname{QDM}(B)$ by the trivial one $e^{-z\partial_{\lambda}}$, but it acts highly non-trivially on the pull-back by the $\lambda$-dependent coordinate change $\tau\mapsto\sigma(\tau)$.
-From this point of view, the solution $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$ considered above corresponds to the standard solution $M_{B}(\sigma(\tau))$ for $\sigma^{*}\operatorname{QDM}(B)$.
+From this point of view, the solution $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$ considered above corresponds to the standard solution $M_{B}(\sigma(\tau))$ for $\sigma^{*}(\operatorname{QDM}(B)|_{Q_{B}\to Q})$.
 5.3. Fourier transformation via stationary phase method
 In this section we define the stationary phase asymptotics of the Fourier integral
-[…]
+$\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$
 for $J(\lambda)\in H^{*}_{S^{1}}(V)[z,z^{-1}]=H^{*}(B)[\lambda,z,z^{-1}]$. The integral can be written as
-[…]
+$\int e^{-\varphi(\lambda)/z}\lambda^{-c_{1}(V)/z}\lambda^{-r/2}K(\lambda)d\lambda$
 where $\varphi(\lambda)=r(\lambda\log\lambda-\lambda)-\lambda\log q$ and $K(\lambda)=(\widetilde{\Delta}_{V}^{\lambda})^{-1}J(\lambda)$ (see §3.2 for $\widetilde{\Delta}_{V}^{\lambda}$).
@@
 We have
-$\varphi(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})=-r\lambda_{0}+\frac{r}{2}s^{2}+\varphi_{\geq 3}(s,\lambda_{0})\quad\text{with $\varphi_{\geq 3}(s,\lambda_{0})=r\sum_{k=3}^{\infty}\frac{k-1}{k!}\frac{s^{k}}{\sqrt{\lambda_{0}}^{k-2}}$}$
+$\varphi(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})=-r\lambda_{0}+\frac{r}{2}s^{2}+\varphi_{\geq 3}(s,\lambda_{0})\quad\text{with $\varphi_{\geq 3}(s,\lambda_{0})=r\sum_{k=3}^{\infty}\frac{k-1}{k!}\frac{s^{k}}{(\sqrt{\lambda_{0}})^{k-2}}$}$
 and hence the Fourier integral can be rewritten in the form
-[…]
+(5.5) $e^{r\lambda_{0}/z}\lambda_{0}^{-c_{1}(V)/z}\lambda_{0}^{-(r-1)/2}\int e^{-rs^{2}/(2z)}L(s,\lambda_{0})ds$
 with $L(s,\lambda_{0})\in H^{*}(B)[z,z^{-1},s](\!(\lambda_{0}^{-1/2})\!)$ given by
 $L(s,\lambda_{0})$ $=e^{-(\frac{r}{2}-1)u}e^{-uc_{1}(V)/z}e^{-\varphi_{\geq 3}(s,\lambda_{0})/z}K(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})\qquad\text{with $u=s/\sqrt{\lambda_{0}}$.}$
-We expand $L(s,\lambda_{0})$ in power series in $s$ and perform the Gaussian integral (5.3) termwise over $s\in\mathbb{R}$ (assuming $z>0$) and arrive at the formal stationary phase asymptotics
-[…]
+By expanding $L(s,\lambda_{0})$ as a power series in $s$ and performing the Gaussian integral (5.5) termwise over $s\in\mathbb{R}$ (assuming $z>0$), we obtain the formal stationary phase asymptotics
+$\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda\sim\sqrt{2\pi z}\cdot e^{r\lambda_{0}/z}\mathscr{F}_{0}(J)$
 with $\mathscr{F}_{0}(J)\in\lambda_{0}^{-(r-1)/2}H^{*}(B)[z,z^{-1}](\!(\lambda_{0}^{-1})\!)[\log\lambda_{0}]$ given by
@@
 Parts (1), (2) follow by taking the stationary phase asymptotics of the standard properties of the Fourier transformation:
-[…]
-[…]
-[…]
+$zq\partial_{q}\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$ $=\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}\lambda J(\lambda)d\lambda$
+$q\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$ $=\int e^{(\lambda+z)\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$
+$=\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}\mathcal{S}J(\lambda)d\lambda$
 where we shift $\lambda$ as $\lambda\mapsto\lambda-z$ and used (5.4) in the last line. We note that the shift $\lambda\mapsto\lambda-z$ translates into the change of variables
@@
 in terms of the coordinate $s$ around the critical point (given by $\lambda=\lambda_{j}e^{s/\sqrt{\lambda_{j}}}$) and this makes sense in the ring $\mathbb{C}[z,z^{-1},s](\!(\lambda_{j}^{-1/2})\!)$.
-To show part (3), observe that the map $J\mapsto e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J)$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$ and $\deg q=2r$, $\deg z=2$.
+To show part (3), observe that the map $J\mapsto e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J)$ is homogeneous of degree (r1)-(r-1) with respect to the usual grading on $H^{*}_{S^{1}}(V)$ and $\deg q=2r$, $\deg z=2$.
 This means
@@
 The formal asymptotic expansion of the integral
-[…]
+(5.6) $\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J_{V}(\tau)d\lambda$
 is given by applying the differential operator $\sqrt{\frac{2\pi z}{\varphi^{\prime\prime}(\lambda_{j})}}e^{z(\partial_{\lambda})^{2}/(2\varphi^{\prime\prime}(\lambda_{j}))}$ to the integrand multiplied by $e^{\varphi^{\prime\prime}(\lambda_{j})(\lambda-\lambda_{j})^{2}/(2z)}$ and evaluating it at the critical point $\lambda=\lambda_{j}$.
@@
 Proposition 5.6.
-Write $\tau(\lambda)=\sum_{i,k}\tau^{i,k}\lambda^{k}\phi_{i}$ for the parameter in $H^{*}_{S^{1}}(V)$. There exist
-$\sigma_{j}$ $\in-c_{1}(V)\log\lambda_{j}+\tau(\lambda_{j})+\mathfrak{m}H^{*}(B)[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
+Write $\tau(\lambda)=\sum_{i,k}\tau^{i,k}\lambda^{k}\phi_{i}$ for the parameter in $H^{*}_{S^{1}}(V)$.
+Let $M_{B}(\sigma)=M_{B}(\sigma;Q_{B})$ denote the fundamental solution (2.6) of the quantum connection of $B$.
+There exist
+$\sigma_{j}$ $\in-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\tau(\lambda_{j})+\mathfrak{m}H^{*}(B)[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
 $\upsilon_{j}$ $\in\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}(1+\mathfrak{m}^{\prime}H^{*}(B)[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!])$
-with $\lambda_{j}=q^{1/r}e^{2\pi\sqrt{-1}j/r}$ such that $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$, where
+with $\lambda_{j}=q^{1/r}e^{2\pi\sqrt{-1}j/r}$ such that $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$, where
   • $q^{\bullet/r}\boldsymbol{\tau}$ stands for the infinite set $\{q^{k/r}\tau^{i,k}\}$ of variables;
   • $K[\![q^{-c_{1}(V)/r}Q]\!]=\left\{\sum_{d\in\operatorname{Eff}(B)}c_{d}q^{-c_{1}(V)\cdot d/r}Q^{d}:c_{d}\in K\right\}$ for a module $K$; and
-  • $\mathfrak{m}$ ((((resp.
-$\mathfrak{m}^{\prime}$)))) is the closed ideal of $\mathbb{C}[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ ((((resp.
-of $\mathbb{C}[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$)))) generated by $q^{-1/r}$ and $q^{-c_{1}(V)\cdot d/r}Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$.
-Moreover, $\sigma_{j}+c_{1}(V)\log\lambda_{j}$ and $\upsilon_{j}$ are homogeneous of degree $2$ and $-(r-1)$ respectively with respect to the usual grading on $H^{*}(B)$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$ and $\deg z=2$.
+  • $\mathfrak{m}$ ((resp.
+$\mathfrak{m}^{\prime}$)) is the closed ideal of $\mathbb{C}[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ ((resp.
+of $\mathbb{C}[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$)) generated by $q^{-1/r}$ and $q^{-c_{1}(V)\cdot d/r}Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$.
+Moreover, $\sigma_{j}$ and $\upsilon_{j}$ are homogeneous of degree $2$ and (r1)-(r-1) respectively with respect to the usual grading on $H^{*}(B)$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$ and $\deg z=2$.
 Proof.
@@
 where we set $u=s/\sqrt{\lambda_{j}}$ as before and $Q\to\lambda^{c_{1}(V)}Q$ means $Q^{d}\to\lambda^{c_{1}(V)\cdot d}Q^{d}$.
-The quantity in the big parenthesis []delimited-[][$\cdots$][ ⋯ ] belongs to $H^{*}(B)[z,z^{-1},s][\![\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and defines a $\mathbb{C}[\![s,\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$-valued point of $\mathcal{L}_{B}$: the Dilaton and String Equations show that $\mathcal{L}_{B}$ is invariant under $e^{-(\frac{r}{2}-1)u}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}$ and the Divisor Equation shows that $\mathcal{L}_{B}$ is invariant under $f(Q)\to e^{-uc_{1}(V)/z}f(Qe^{-uc_{1}(V)})$.
+The quantity in the big parenthesis $\cdots$] belongs to $H^{*}(B)[z,z^{-1},s][\![\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and defines a $\mathbb{C}[\![s,\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$-valued point of $\mathcal{L}_{B}$: the Dilaton and String Equations show that $\mathcal{L}_{B}$ is invariant under $e^{-(\frac{r}{2}-1)u}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}$ and the Divisor Equation shows that $\mathcal{L}_{B}$ is invariant under $f(Q)\to e^{-uc_{1}(V)/z}f(Qe^{-uc_{1}(V)})$.
 By Lemma 2.7, the action of $e^{z(\partial_{s})^{2}/(2r)}$ preserves points on the cone [^9] Strictly speaking, the operator $(z\partial_{s})^{2}/(2r)$ does not vanish at the origin of the formal parameters as required in Lemma 2.7, but the conclusion holds since the action of $e^{z(\partial_{s})^{2}/(2r)}$ is well-defined..
@@
 It follows from the homogeneity of $e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))$ that $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ are homogeneous of degree zero after the change $Q\to\lambda_{j}^{-c_{1}(V)}Q$ of variables.
-By the Divisor Equation again, we obtain $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$ with
-$\sigma_{j}=-c_{1}(V)\log\lambda_{j}+\tilde{\sigma}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q},\quad\upsilon_{j}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\tilde{\upsilon}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q}.$
+By the Divisor Equation again, we obtain $q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$ with
+$\sigma_{j}=-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\tilde{\sigma}_{j}\bigr|_{Q\to\lambda_{j}^{-c_{1}(V)}Q},\quad\upsilon_{j}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\tilde{\upsilon}_{j}\bigr|_{Q\to\lambda_{j}^{-c_{1}(V)}Q}.$
 The conclusion follows. ∎
-Remark 5.7.
-The leading term $-c_{1}(V)\log\lambda_{j}$ of $\sigma_{j}$ accounts for difference of the degree of $Q^{d}$ between $B$ and $V$.
 Example 5.7.
-Computing the formal stationary phase asymptotics of (5.4) at $Q=0$ over the Artinian local ring $H^{*}(B)$, we see that $r\lambda_{j}+\sigma_{j}|_{Q=0}$ is given by the critical value of (cf. (3.7))
+Computing the formal stationary phase asymptotics of (5.6) at $Q=0$ over the Artinian local ring $H^{*}(B)$, we see that $r\lambda_{j}+\sigma_{j}|_{Q=0}-\frac{1}{r}c_{1}(V)\log q$ is given by the critical value of (cf.
+(3.4))
 $f(\lambda)=\tau(\lambda)+\lambda\log q-\sum_{\delta:\text{Chern roots of $V$}}((\lambda+\delta)\log(\lambda+\delta)-(\lambda+\delta))$
 at a unique $H^{*}(B)$-valued critical point $\hat{\lambda}_{j}\in\lambda_{j}(1+\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1},\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!])$, which is a solution $\lambda$ to the equation $q=e_{\lambda}(V)e^{-\tau^{\prime}(\lambda)}$.
-The critical point $\hat{\lambda}_{j}$ and the critical value $f(\hat{\lambda}_{j})$ are the eigenvalues of the quantum products $(p\star)|_{Q=0}$ and $(E_{\mathbb{P}(V)}\star)|_{Q=0}$ in the $H^{*}(B)$-algebra $QH^{*}(\mathbb{P}(V))|_{Q=0}$ respectively.
+Here $\hat{\lambda}_{j}$ is an eigenvalue of the quantum product $(p\star)|_{Q=0}$ in the $H^{*}(B)$-algebra $QH^{*}(\mathbb{P}(V))|_{Q=0}$.
 (It is easy to see that $H^{*}(B)$ acts by cup product on $QH^{*}(\mathbb{P}(V))|_{Q=0}$.)
@@
 Let $\sigma_{j}$ be as in Proposition 5.6.
-We regard it as a multi-valued map from $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ to $H^{*}(B)$, where $r^{\prime}$ is as in (5.1), and consider the pull-back $\sigma_{j}^{*}\operatorname{QDM}(B)$ of the quantum $D$-module of $B$ by $\sigma_{j}$.
-We define $\sigma_{j}^{*}\operatorname{QDM}(B)$ to be the module
-$\sigma_{j}^{*}\operatorname{QDM}(B)=H^{*}(B)[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
-equipped with the pull-back $\sigma_{j}^{*}\nabla$ of the quantum connection of $B$:
+We regard it as a map from $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ to $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,\sigma]\!]$, where $r^{\prime}$ is as in (5.1), and consider the pull-back of the extended quantum $D$-module $\operatorname{QDM}(B)_{\rm ext}$ (5.3) of $B$ by $\sigma_{j}$.
+The pull-back is defined to be the module
+$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}:=H^{*}(B)[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
+equipped with the pull-back $\sigma_{j}^{*}\nabla$ of the connection for $\operatorname{QDM}(B)_{\rm ext}$:
 $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ $=\partial_{\tau^{i,k}}+z^{-1}(\partial_{\tau^{i,k}}\sigma_{j})\star_{\sigma_{j}}$
-$(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(q\partial_{q}\sigma_{j})\star_{\sigma_{j}}$
+$(\sigma_{j}^{*}\nabla)_{q\partial_{q}}$ $=q\partial_{q}+z^{-1}(-(c_{1}(V)/r)\star_{\sigma_{j}}+(q\partial_{q}\sigma_{j})\star_{\sigma_{j}})$
 $(\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\sigma_{j}}+(\xi Q\partial_{Q}\sigma_{j})\star_{\sigma_{j}})\qquad(\xi\in H^{2}(B))$
-$(\sigma_{j}^{*}\nabla)_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}E_{B}\star_{\sigma_{j}}+\mu_{B}$
-and the pairing $P_{B}$ as in (2.8). The Divisor Equation show that these pull-backs are well-defined and define maps
-$q^{-k/r}(\sigma_{j}^{*}\nabla)_{\tau^{i,k}},\ (\sigma_{j}^{*}\nabla)_{q\partial_{q}},\ (\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}},\ (\sigma_{j}^{*}\nabla)_{z\partial_{z}}\colon\sigma_{j}^{*}\operatorname{QDM}(B)\to z^{-1}\sigma_{j}^{*}\operatorname{QDM}(B).$
-The operator $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ has poles along the infinity disivor $(q^{-1/r^{\prime}}=0)$, but this is an apparent singularity: $\partial_{\tau^{i,k}}$ viewed as a vector field on $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ has poles of the same order along $(q^{-1/r^{\prime}}=0)$.
-We introduce a localized version along the infinity disivor $(q^{-1/r^{\prime}}=0)$ as follows:
-$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ $:=\sigma_{j}^{*}\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$
+$(\sigma_{j}^{*}\nabla)_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\sigma_{j}})+\mu_{B}$
+and the pairing $P_{B}$ as in (2.8). The Divisor Equation shows that these pull-backs are well-defined and define maps
+$q^{-k/r}(\sigma_{j}^{*}\nabla)_{\tau^{i,k}},\ (\sigma_{j}^{*}\nabla)_{q\partial_{q}},\ (\sigma_{j}^{*}\nabla)_{\xi Q\partial_{Q}},\ (\sigma_{j}^{*}\nabla)_{z\partial_{z}}\colon\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}\to z^{-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}.$
+The operator $(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}$ has poles along the infinity divisor $(q^{-1/r^{\prime}}=0)$, but this is an apparent singularity: $\partial_{\tau^{i,k}}$ viewed as a vector field on $\operatorname{Spf}\mathbb{C}[\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ has poles of the same order along $(q^{-1/r^{\prime}}=0)$.
+Similarly to (5.3), we denote by $\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ the restriction of $\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}$ to the localized base $\operatorname{Spf}\mathbb{C}(\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$.
+$\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ $:=\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext}\otimes_{\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!]$
 $=H^{*}(B)[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\boldsymbol{\tau}]\!].$
 Proposition 5.8.
-The Fourier transformation $\mathscr{F}_{j}$ induces a map $\Pi_{j}\colon\operatorname{QDM}_{S^{1}}(V)\to\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules such that
-(5.5) $\mathscr{F}_{j}(M_{V}(\tau)s)=M_{B}(\sigma_{j})\Pi_{j}(s)\qquad\text{for all $s\in\operatorname{QDM}_{S^{1}}(V)$.}$
-The map $\Pi_{j}$ satisfies the following:
+The Fourier transformation $\mathscr{F}_{j}$ induces a map $\Pi_{j}\colon\operatorname{QDM}_{S^{1}}(V)\to\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ of $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-modules such that
+(5.7) $\mathscr{F}_{j}(M_{V}(\tau)s)=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\Pi_{j}(s)$
+for all $s\in\operatorname{QDM}_{S^{1}}(V)$. The map $\Pi_{j}$ satisfies the following:
   • (1)
@@
   • (6)
-    $\Pi_{j}$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(B)$ and the degree of the variables as in Proposition 5.6;
+    $\Pi_{j}$ is homogeneous of degree (r1)-(r-1) with respect to the usual grading on $H^{*}_{S^{1}}(V)$, $H^{*}(B)$ and the degree of the variables as in Proposition 5.6;
   • (7)
@@
 Proof.
-By differentiating the identity $\mathscr{F}_{j}(M_{V}(\tau)1)=M_{B}(\sigma_{j})\upsilon_{j}$ from Proposition 5.6 by $\tau^{i,k}$ and using the differential equations (2.7) for $M_{V}$ and $M_{B}$, we obtain
-$\mathscr{F}_{j}(M_{V}(\tau)\phi_{i}\lambda^{k})=M_{B}(\sigma_{j})z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}.$
+By differentiating the identity $\mathscr{F}_{j}(M_{V}(\tau)1)=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$ from Proposition 5.6 by $\tau^{i,k}$ and using the differential equations (2.7) for $M_{V}$ and $M_{B}$, we obtain
+$\mathscr{F}_{j}(M_{V}(\tau)\phi_{i}\lambda^{k})=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}.$
 Therefore the $\mathbb{C}[z][\![Q,\boldsymbol{\tau}]\!]$-module map $\Pi_{j}$ sending $\phi_{i}\lambda^{k}$ to $z(\sigma_{j}^{*}\nabla)_{\tau^{i,k}}\upsilon_{j}$ satisfies (5.7).
 Parts (1)-(5) follow immediately from (5.7), Proposition 5.5 and the intertwining properties (2.7), (2.21) of $M_{V},M_{B}$.
-Part (6) follows from the homogeneity of $\sigma_{j}+c_{1}(V)\log\lambda_{j}$, $\upsilon_{j}$ in Proposition 5.6.
+Part (6) follows from the homogeneity of $\sigma_{j}$, $\upsilon_{j}$ in Proposition 5.6.
 Part (7) follows by a direct computation of
@@
 We combine the isomorphism $\operatorname{FT}\colon\operatorname{QDM}_{S^{1}}(V)\cong\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ in Theorem 4.2 with the maps $\Pi_{j}$ in the above proposition to obtain a map
-$\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}.$
-By Theorem 4.2 and Proposition 5.9, this is a homomorphism of $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$-modules homogeneous of degree $-(r-1)$, and intertwines the connection $\hat{\tau}^{*}\nabla$ with
+$\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\sigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}.$
+By Theorem 4.2 and Proposition 5.8, this is a homomorphism of $\mathbb{C}[z,q][\![Q,\boldsymbol{\tau}]\!]$-modules homogeneous of degree (r1)-(r-1), and intertwines the connection $\hat{\tau}^{*}\nabla$ with
 $\bigoplus_{j=0}^{r-1}\left(\sigma_{j}^{*}\nabla+\frac{\lambda_{j}}{z}\frac{dq}{q}-r\lambda_{j}\frac{dz}{z^{2}}\right).$
-We note that the pull-back $(\sigma_{j}+r\lambda_{j})^{*}\nabla$ of the quantum connection of $B$ by $\sigma_{j}+r\lambda_{j}$ is well-defined [^8] Although $M_{B}(\sigma_{j}+r\lambda_{j})$ is ill-defined in the $q^{-1/r}$-adic topology.
+We note that the pull-back $(\sigma_{j}+r\lambda_{j})^{*}\nabla$ of the quantum connection of $B$ by $\sigma_{j}+r\lambda_{j}$ is well-defined [^10] Despite the fact that the pull-back $M_{B}(\sigma_{j}+r\lambda_{j})$ of the fundamental solution is ill-defined in the $q^{-1/r^{\prime}}$-adic topology.
 and that
@@
 Hence we obtain a map
-(5.6) $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm loc}$
-that is fully compatible with the connection.
-We now restrict the parameter $\tau$ to lie in $\mathsf{H}:=\bigoplus_{k=0}^{r-1}H^{*}(B)\lambda^{k}\subset H^{*}_{S^{1}}(V)$: the mirror map $\tau\mapsto\hat{\tau}(\tau)$ in Theorem 4.2 restricted to $\mathsf{H}$ gives an isomorphism between $\mathsf{H}$ and $H^{*}(\mathbb{P}(V))$ in the formal neighbourhood of $q=Q=0$ since $\hat{\tau}(\tau)|_{q=Q=0}=\kappa(\tau)$.
-We compose the map $\sigma_{j}$ with the inverse $\hat{\tau}\mapsto\tau=\tau(\hat{\tau})$ of this isomorphism and define
+(5.8) $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}\colon\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
+[…]
 $\varsigma_{j}(\hat{\tau}):=\sigma_{j}(\tau(\hat{\tau}))+r\lambda_{j}.$
-Because of the homogeneity of $\hat{\tau}(\tau)$ from Theorem 4.2(6), the inverse mirror map $\tau(\hat{\tau})$ belongs to $H^{*}(B)[q][\![Q,\hat{\tau}]\!]$.
-Therefore the composition is well-defined; $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ belongs to $H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$ and is homogeneous of degree $2$.
-Moreover we have
-(5.7) $\varsigma_{j}(\hat{\tau})\Bigr{|}_{Q=0}\in r\lambda_{j}-c_{1}(V)\log\lambda_{j}+\alpha+\sum_{k=0}^{r-1}\sum_{i=0}^{s}\hat{\tau}^{i,k}\phi_{i}\lambda_{j}^{k}(1+\beta_{k})+(\hat{\tau})^{2}H^{*}(B)$
-for some $\alpha,\beta_{k}\in\lambda_{j}^{-1}H^{*}(B)[\![\lambda_{j}^{-1}]\!]$, where $(\hat{\tau})\subset\mathbb{C}(\!(q^{-1/r})\!)[\![\hat{\tau}]\!]$ denote the ideal generated by $\hat{\tau}^{i,k}$ with $0\leq i\leq s$, $10\leq k\leq r-$ 1.
-This follows by Theorem 4.2(7) and Proposition 5.6.
+[…]
+(5.9) $\varsigma_{j}(\hat{\tau})\Bigr|_{Q=0}\in r\lambda_{j}-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\alpha_{j}+\sum_{k=0}^{r-1}\sum_{i=0}^{s}\hat{\tau}^{i,k}(\phi_{i}+\beta_{j,i,k})\lambda_{j}^{k}+(\hat{\tau})^{2}H^{*}(B)$
+for some $\alpha_{j},\beta_{j,i,k}\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$, where $(\hat{\tau})\subset\mathbb{C}(\!(q^{-1/r})\!)[\![\hat{\tau}$]\!] denote the ideal generated by $\hat{\tau}^{i,k}$ with $0\leq i\leq s$, $0\leq k\leq r-1$.
+Parts (4) and (5) of Theorem 5.1 follow from (5.9).
 The above map induces a $\mathbb{C}[z,q][\![Q,\hat{\tau}$]\!]-module map
-$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
-compatible with the connection and homogeneous of degree $-(r-1)$. Combining Theorem 4.2(8) and Proposition 5.9(7), we have
+$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))\longrightarrow\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
+compatible with the connection and homogeneous of degree (r1)-(r-1). Combining Theorem 4.2(8) and Proposition 5.8(7), we have
 $\Phi_{j}(\phi_{i}p^{k})\Bigr|_{Q=\hat{\tau}=0}\in\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}\left(\phi_{i}+q^{-1/r}H^{*}(B)[z][\![q^{-1/r}]\!]\right)$
-for $10\leq k\leq r-$ 1 and therefore $\Phi$ induces an isomorphism over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$. Inverting $\Phi|_{Q=\hat{\tau}=0}$, we obtain
+for $0\leq k\leq r-1$ and therefore $\Phi$ induces an isomorphism over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!].
+This is part (6) of Theorem 5.1.
+Inverting $\Phi|_{Q=\hat{\tau}=0}$, we obtain
 (5.10) $\Phi^{-1}(\phi_{i}e_{k})\Bigr|_{Q=\hat{\tau}=0}\in\frac{1}{\sqrt{r}}\sum_{j=0}^{r-1}\lambda_{k}^{(r-1)/2-j}p^{j}\left(\phi_{i}+q^{-1/r}H^{*}(B)[z][\![q^{-1/r}]\!]\right)$
@@
 By (5.9) we have
-$(\varsigma_{j}^{*}\nabla)_{q\partial_{q}}\Bigr{|}_{Q=\hat{\tau}=0}=q\partial_{q}+\frac{1}{z}\left(\lambda_{j}+\frac{c_{1}(V)}{r}+\alpha^{\prime}\right)$
-where $c_{1}(V)$ and $\alpha^{\prime}=q\partial_{q}\alpha\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$ acts by cup product.
+$(\varsigma_{j}^{*}\nabla)_{q\partial_{q}}\Bigr|_{Q=\hat{\tau}=0}=q\partial_{q}+\frac{1}{z}\left(\lambda_{j}-\frac{c_{1}(V)}{r}+q\partial_{q}\alpha_{j}\right)$
+where $c_{1}(V)$ and $q\partial_{q}\alpha_{j}\in q^{-1/r}H^{*}(B)[\![q^{-1/r}]\!]$ acts by cup product.
 Writing $x=q^{-1/r}$, we find that $v(x):=P_{2}^{-1}\Phi^{-T}P_{1}\Phi^{-1}|_{Q=\hat{\tau}=0}=v_{0}+v_{1}x+v_{2}x^{2}+\cdots$ satisfies a differential equation of the form
 $(x\partial_{x}+x^{-1}S+(N_{0}+N_{1}x+\cdots))v(x)=0$
-where $S,N_{i}$ are mutually commuting endomorphisms on $\operatorname{End}(\bigoplus_{j=0}^{r-1}H^{*}(B)))$ such that $S$ is semisimple and $N_{0}$ is nilpotent. We obtain a sequence of equations
+where $S,N_{i}$ are mutually commuting endomorphisms on $\operatorname{End}(\bigoplus_{j=0}^{r-1}H^{*}(B))$ such that $S$ is semisimple and $N_{0}$ is nilpotent. We obtain a sequence of equations
 $Sv_{n+1}+(n+N_{0})v_{n}+N_{1}v_{n-1}+N_{2}v_{n-2}+\cdots=0.$
-From $Sv_{0}=N_{i}v_{0}=0$ and (*$*_{0}$) we have $Sv_{1}=0$.
-Then (*$*_{1}$) implies $S^{2}v_{2}=0$.
-Since $S$ is semisimple, $Sv_{2}=0$ and (*$*_{1}$) implies $(1+N_{0})v_{1}=0$; the nilpotence of $N_{0}$ implies $v_{1}=0$.
+From $Sv_{0}=N_{i}v_{0}=0$ and ($_{0}$) we have $Sv_{1}=0$.
+Then ($_{1}$) implies $S^{2}v_{2}=0$.
+Since $S$ is semisimple, $Sv_{2}=0$ and ($_{1}$) implies $(1+N_{0})v_{1}=0$; the nilpotence of $N_{0}$ implies $v_{1}=0$.
 In this way we can show inductively that $v_{1}=v_{2}=\cdots=0$.
@@
 $P_{1}^{-1}\Phi^{T}P_{2}\Phi\Bigr|_{Q=\hat{\tau}=0}=I.$
-We now use the fact that $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ is flat with respect to the connection on the endomorphism bundle of $\operatorname{QDM}(\mathbb{P}(V))$.
-It satisfies a differential equation in $(Q,\hat{\tau})$ with at most logarithmic singularities at $Q=0$ and nilpotent residues.
-Expanding $P_{1}^{-1}\Phi^{T}P_{2}\Phi$ in power series in $Q$ and $\hat{\tau}$ and using a similar argument as before, we can easily show that $P_{1}^{-1}\Phi^{T}P_{2}\Phi=I$ identically.
-Thus $\Phi$ preserves the pairing.
+[…]
 Remark 5.9.
-The map (5.6) glues the connections $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$; $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ is defined around $q=Q=\boldsymbol{\tau}=0$ and $(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)$ is defined around $q^{-1/r^{\prime}}=q^{-c_{1}(V)/r}Q=q^{\bullet/r}\boldsymbol{\tau}=0$.
+The map (5.8) glues the connections $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ and $\bigoplus_{j=0}^{r-1}(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext}$; $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))$ is defined around $q=Q=\boldsymbol{\tau}=0$ and $(\sigma_{j}+r\lambda_{j})^{*}\operatorname{QDM}(B)_{\rm ext}$ is defined around $q^{-1/r^{\prime}}=q^{-c_{1}(V)/r}Q=q^{\bullet/r}\boldsymbol{\tau}=0$.
 They glue to a connection over a global “Kähler moduli space” obtained by adding the infinity divisor $(q=\infty)$ to $\operatorname{Spf}(\mathbb{C}[q][\![Q,\boldsymbol{\tau}]\!])$.
-See Figure 1 for the global Kähler moduli space.
+See Figure 1.
 5.7. Semisimplicity
@@
 $(\phi_{i}p^{k})\star_{\hat{\tau}}=\Phi^{-1}\left(\textstyle\bigoplus_{j=0}^{r-1}(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})\star_{\varsigma_{j}(\hat{\tau})}\right)\Phi+z\Phi^{-1}\partial_{\hat{\tau}^{i,k}}\Phi.$
-Setting $z=0$, this shows that the map sending $\phi_{i}p^{k}$ to $\bigoplus_{j=0}^{r-1}\partial_{\hat{\tau}^{i,k}}\varsigma_{j}$ gives a ring isomorphism $(H^{*}(\mathbb{P}(V)),\star_{\hat{\tau}})\cong\bigoplus_{j=0}^{r-1}(H^{*}(B),\star_{\varsigma_{j}(\hat{\tau})})$ over $\mathbb{C}(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$.
-This proves Corollary 1.8.
+[…]
 Corollary 1.8 can be applied to show the semisimplicity of big quantum cohomology of certain partial flag varieties.
@@
 It is expected that rational homogeneous varieties $G/P$ have semisimple big quantum cohomology. We refer the reader to [49] and references therein for recent developments on this question.
-$Q=\hat{\tau}=0$, $q>0$small deformation of $(Q,\hat{\tau})$
+[…]
 Figure 2. Decomposition of eigenvalues of the Euler multiplication of $\mathbb{P}(V)$ into $r$ groups ($r=6$ in this picture).
@@
 See Figure 2 for a typical decomposition of eigenvalues.
-Similar pictures have appeared in [28, Figure 16], [34, Figure 1].
+Similar pictures have appeared in [28, Figure 16], [38, Figure 1].
+5.8. Reconstruction method of Hinault-Yu-Zhang-Zhang
+[…]
+We describe the initial conditions for $\Phi$ and $\varsigma_{j}$ along $Q=\hat{\tau}=0$.
+Recall that the isomorphism $\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}$ arises from the restriction of the isomorphism $\bigoplus_{j=0}^{r-1}\Pi_{j}\circ\operatorname{FT}^{-1}$ in (5.8) to the subspace $\mathsf{H}=\bigoplus_{k=0}^{r-1}H^{*}(B)\lambda^{k}$.
+Since $\hat{\tau}(\tau)|_{Q=\tau=0}=0$ by Theorem 4.2(7), we have $\Phi_{j}|_{Q=\hat{\tau}=0}=\Pi_{j}\circ\operatorname{FT}^{-1}|_{Q=\tau=0}$.
+We know by Theorem 4.2(8) that $\operatorname{FT}^{-1}(\phi_{i}p^{k})|_{Q=\tau=0}=\phi_{i}\lambda^{k}$.
+Therefore we have
+$\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\Pi_{j}(\phi_{i}\lambda^{k})|_{Q=\tau=0}=e^{-\sigma_{j}|_{Q=\tau=0}/z}q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(\phi_{i}\lambda^{k})$
+by Proposition 5.8.
+Consider the case $\phi_{i}p^{k}=1$.
+Since $q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(1)$ is of the form $\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}e^{2\pi\sqrt{-1}jc_{1}(V)/(rz)}(1+O(q^{-1/r}))$, its logarithm with respect to the cup product structure on $H^{*}(B)$ is well-defined.
+Because $\Pi_{j}(1)$ contains no negative powers of $z$, the initial value $\varsigma_{j}^{\circ}:=\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}+\sigma_{j}|_{Q=\tau=0}$ for $\varsigma_{j}$ can be calculated as
+(5.11) $\varsigma_{j}^{\circ}=r\lambda_{j}+[z^{-1}]\log\left(q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(1)\right)$
+where $z^{-1}](\cdots)$ denotes the coefficient of $z^{-1}$. The initial value $\Phi_{j}^{\circ}:=\Phi_{j}|_{Q=\hat{\tau}=0}$ for $\Phi_{j}$ is then given by
+(5.12) $\Phi_{j}^{\circ}(\phi_{i}p^{k})=e^{-(\varsigma_{j}^{\circ}-r\lambda_{j})/z}q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(\phi_{i}\lambda^{k})$
+for $0\leq i\leq s$ and $0\leq k\leq r-1$. Note that $\mathscr{F}_{j}(\phi_{i}\lambda_{k})$ can be calculated explicitly in terms of the Chern classes of $V$ via the procedure in §5.3.
+We briefly describe the reconstruction of $\Phi$ and $\varsigma_{j}$ from their initial conditions (5.11), (5.12) by means of Birkhoff factorization.
+We leave the detailed verification to the reader, as the argument is completely parallel to the case of blowups treated in [29, §5.8].
+We set:
+(5.13) $s_{j}(\hat{\tau})=\varsigma_{j}(\hat{\tau})-\varsigma_{j}^{\circ}.$
+[…]
+$M=\bigoplus_{j=0}^{r-1}e^{-\varsigma_{j}^{\circ}/z}M_{B}(\varsigma_{j}^{\circ}+s_{j};Qq^{-c_{1}(V)/r}).$
+This satisfies $M|_{Q=s_{0}=\dots=s_{r-1}=0}=\operatorname{id}$ and $M=\operatorname{id}+O(z^{-1})$. Then the composition $(\Phi^{\circ})^{-1}\circ M$ admits a unique Birkhoff facorization of the form
+$(\Phi^{\circ})^{-1}\circ M=M^{\prime}\circ\Phi^{-1}$
+where $M^{\prime}=\operatorname{id}+O(z^{-1})$ is a power series in $z^{-1}$ and $\Phi^{-1}$ is a power series in $z$ such that
+$M^{\prime}$ $\in\operatorname{End}(H^{*}(\mathbb{P}(V)))\otimes\mathbb{C}[z^{-1}](\!(q^{-1/r^{\prime}})\!)[\![Q,s_{0},\dots,s_{r-1}]\!],$
+$\Phi^{-1}$ $\in\operatorname{Hom}(H^{*}(B)^{\oplus r},H^{*}(\mathbb{P}(V)))\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,s_{0},\dots,s_{r-1}]\!],$
+[…]